### v2

# Near-Field Localization and Phase Shift Optimization for RIS-Assisted Non-Ideal OFDM Systems

Abstract
By incorporating reconfigurable intelligent surface (RIS) into communication-assisted localization systems, the issue of signal blockage caused by obstacles can be addressed, and passive beamforming can be employed to enhance localization accuracy. However, existing works mainly consider ideal channels and do not account for the effects of realistic impairments like carrier frequency offset (CFO) and phase noise (PN) on localization. This paper proposes an iterative joint estimation algorithm for CFO, PN, and user position based on maximum a posteriori (MAP) criterion and gradient descent (GD) algorithm. Closed-form expressions for CFO and PN updates are provided. The hybrid Cramér-Rao lower bound (HCRLB) for the estimation parameters is derived, and the ambiguity in CFO and PN estimation is analyzed. To minimize the HCRLB, a non-convex RIS shift optimization problem is formulated and is transformed into a convex semidefinite programming (SDP) problem using the technique of semidefinite relaxation (SDR) and Schur complement. After optimizing the RIS phase shift, the theoretical positioning accuracy within the area of interest (AOI) can be improved by two orders of magnitude, with a maximum positioning root mean square error (RMSE) lower than $\rm 10^{-2}m$.
Index Terms:
Near-field localization, reconfigurable intelligent surface, passive beamforming, carrier frequency offset, phase noise.
I Introduction
With the widespread application of fifth-generation (5G) networks worldwide, research focus has shifted towards future six-th generation (6G) mobile communication technologies. For 5G mobile networks, the demand for positioning accuracy has already reached sub-meter levels due to the need to support various location-based services. In 6G networks, it is expected that high-precision positioning at the centimeter level will play an even more important role [1, 2, 3]. While global navigation satellite systems (GNSSs) have been widely used in various positioning scenarios [25], they suffer from issues such as insufficient positioning accuracy and signal instability due to obstacles [26, 27]. In recent years, research has been conducted on using communication signals to assist in positioning, aiming to achieve more accurate and continuous positioning alongside GNSS signals. For example, Shahmansoori et al. [10] proposed a three-stage positioning algorithm that improves the simultaneous orthogonal matching pursuit (SOMP) algorithm and the space alternating generalized expectation-maximization (SAGE) algorithm, and exploited the sparsity of millimeter-wave multiple-input multiple-output (MIMO) channels to lower the complexity of the channel model. Hu et al. [11] presented a two-dimensional (2D) positioning algorithm based on the estimating signal parameters via rotational invariance technique (ESPRIT), which estimates user positions by constructing signal subspaces. The estimator has closed-form expressions, which reduces the computational complexity significantly compared to traditional estimation methods. However, in these methods that utilize communication signals for localization, the problem of signal blockage by obstacles still remains unavoidable.
In recent years, reconfigurable intelligent surface (RIS) technology has emerged as a key enabling technology for 6G networks [4, 5, 23]. This technology allows for reconfiguring wireless channel environments, enabling communication even when the line-of-sight (LoS) link is obstructed. Introducing the RIS to assist in communication-based positioning has become a research focus [32, 31]. From the perspective of mobility, the research can be divided into two categories: static positioning [17, 18, 19] and dynamic positioning [15, 13]. The former considers scenarios where the user undergoes minimal movement within a coherent time, and the latter involves scenarios with the opposite characteristic. From the perspective of positioning distance, the research can also be divided into two categories: far-field positioning [19, 13, 20] and near-field positioning [18, 19, 13, 24]. In the far-field case, the distance between the user and the RIS exceeds Fraunhofer distance [8], and the signal can be approximated as a planar wave. Conversely, in the near-field case, the signal cannot be treated as parallel propagation and should be considered a spherical wave.
For example, Luan et al. [12] focused on the near-field static 2D positioning scenario. The authors proposed a robust optimization method to optimize the RIS phase shift for improving localization accuracy. The paper also investigated the power optimization problem of the localization system. Elzanaty et al. [13] studied RIS-assisted localization for multi-antenna mobile users. They proposed a three-dimensional (3D) model that can be applied in both far-field and near-field scenarios and analyzed the position and orientation error bounds. Lin et al. [14] introduced a novel twin-RIS configuration comprising two RISs with relative spatial rotation. To extract the parameters of the cascaded channel, the authors employed array signal processing and atomic norm denoising techniques. They established nonlinear equations based on the characteristics of the twin-RIS structure and accurately recovered the multipath parameters. User localization for static users in the far-field scenario is then performed based on the estimated channel parameters. Teng et al. [15] introduced a new message-passing algorithm called BULT to effectively solve the localization problem for mobile users in the far-field scenario. They also derived the Bayesian Cramér-Rao lower bound for the proposed algorithm.
Although there has been extensive research on user localization with RIS in various scenarios such as static, dynamic, far-field, and near-field, most of the existing works assumed ideal channel conditions without considering realistic impairments like carrier frequency offset (CFO) and phase noise (PN) that can introduce uncertainties and affect system performance. Only a few studies have considered modeling certain interferences. For example, [13, 20, 16] took into account clock offset, and [16] considered phase offset. However, to the best of our knowledge, there is currently no research that specifically analyzed CFO and PN in RIS-assisted orthogonal frequency division multiplexing (OFDM) localization systems.
In an OFDM system, CFO and PN can convolve with the data symbols, causing signal constellation rotation and introducing inter-carrier interference (ICI). This interference significantly degrades the overall system performance. Moreover, most aforementioned works have assumed the subcarriers’ signals are independent, which cannot hold in the presence of CFO and PN due to the absence of ICI. Therefore, localization algorithms need to leverage the received time-domain signals to account for these effects and accurately estimate the user’s position in this case.
In this paper, we consider the problem of joint estimation of CFO, PN, and user position in a RIS-assisted OFDM system. The contributions of this paper are outlined as follows.
  • We investigate the problem of static near-field localization in an OFDM system assisted by RIS, considering the effects of CFO and PN. Moreover, we establish a time-domain model and introduce the frequency-domain channel with time-delay parameters into the model through time-frequency transformation to meet the requirements of position estimation.
  • We propose an iterative estimation approach based on the maximum a posteriori (MAP) criterion and gradient descent (GD) algorithm to jointly estimate the CFO, the PN, and the user position. The algorithm significantly reduces complexity by deriving closed-form expressions for CFO and PN estimation.
  • We derive the hybrid Cramér-Rao lower bound (HCRLB) for the joint estimation problem and analyze the phase ambiguity issue in CFO and PN estimation. A new metric is introduced to evaluate the accuracy of joint CFO and PN estimation.
  • We implement phase shift optimization for RIS to enhance the theoretical positioning accuracy. By applying the semidefinite relaxation (SDR) method and utilizing the Schur complement, we convert the initially non-convex optimization problem into a convex semidefinite programming (SDP) problem. It is shown that the theoretical positioning accuracy has been improved by two orders of magnitude.
Figure 1: RIS-aided 3D localization scenario, where the LoS link is blocked.
The paper is structured as follows. Section II introduces the system model, while Section III presents the joint estimation algorithm. Section IV is dedicated to the derivation of the HCRLB for the proposed joint estimation problem. The proposed RIS phase shift optimization method is discussed in Section V. Section VI showcases comprehensive simulation results, and the paper concludes in Section VII.
Notations: $\mathbb{E}(\cdot)$, $p(\cdot)$ denote the expectation and the probability density function (PDF) of the argument, respectively. ${\rm tr}(\mathbf{M})$, $\det\mathbf{M}$, $\mathbf{M}^{\top}$, $\mathbf{M}^{*}$, $\mathbf{M}^{\rm H}$, $\mathbf{M}^{\rm-1}$ denote the trace, determinant, transpose, conjugate, hermitian, and inverse of matrix $\mathbf{M}$, respectively. ${\rm Diag}(\bm{v})$ indicates a diagonal matrix created by assigning the elements of vector $\bm{v}$ along its main diagonal. $\bm{v}[n]$ denotes the $n$-th entry of vector $\bm{v}$. Unless otherwise specified, all indices in the article start from 0. ${\rm Blkdiag}(\mathbf{M}_{0},\mathbf{M}_{1},\cdots,\mathbf{M}_{N-1})$ denotes a block diagonal matrix with $\mathbf{M}_{0},\mathbf{M}_{1},\cdots,\mathbf{M}_{N-1}$ being its diagonal matrices. $\left[\mathbf{M}\right]_{i:j,:}$ refers to a submatrix encompassing rows $i$ to $j$ and all columns of matrix $\mathbf{M}$, while $\left[\mathbf{M}\right]_{:,i:j}$ denotes a submatrix encompassing columns $i$ to $j$ and all rows of matrix $\mathbf{M}$. $\bm{0}_{M\times N}$, and $\mathbf{I}_{M}$ denote $M\times N$ zeros matrix, and $M\times M$ identity matrix, respectively. ${}\Re\left\{\cdot\right\}$ and ${}\Im\left\{\cdot\right\}$ represent the real and imaginary components of the complex value, respectively. $\mathbb{C}^{m\times n}$ and $\mathbb{R}^{m\times n}$ denote $m\times n$ matrices with complex and real entries, respectively. Finally, $\mathcal{CN}(\bm{\mu},\mathbf{\Sigma})$ and $\mathcal{N}(\bm{\mu},\mathbf{\Sigma})$ represent complex and real Gaussian distributions, respectively, characterized by a mean vector $\bm{\mu}$ and covariance matrix $\mathbf{\Sigma}$.
II System Model
Fig. 1 illustrates the RIS-assisted OFDM downlink wireless communication system under consideration, comprising a single anchor node (AN) and a single user equipment (UE). In this scenario, we make the assumption that there are obstacles blocking the LoS path between AN and UE, and the channel state information (CSI) remains unchanged during the localization process, i.e., within the duration of one OFDM symbol, similar to [19, 12, 16]. The positions of the AN and each element of the RIS are assumed to be known, which can be denoted as $\bm{p}_{a}$, and $\bm{p}_{r}=[x_{r},y_{r},z_{r}]^{\top}$, respectively, where $r$ represents the index of the RIS element among a total of $N_{R}$ elements. The position of UE, which is the primary parameter to be estimated, is denoted as $\bm{p}_{u}=[x_{u},y_{u},z_{u}]^{\top}$.
The RIS elements are organized in the form of a uniform planar array (UPA). We take the bottom-left corner of the RIS as the reference point and place it at the coordinate origin. The RIS is positioned on the $yz$ plane, and both edges on the coordinate axes are located in the positive half-axis. The coordinates of the $r$-th RIS element can be represented as $\bm{p}_{r}=[0,d\times((r-1)\mod N_{R}),d\times\lceil\frac{r}{\sqrt{N_{R}}}-1\rceil]^{\top}$, where $d$ stands for the element spacing of RIS, whose value is typically set as half of the wavelength. To satisfy the conditions of the near-field setup, both the distance between AN and RIS, as well as the distance between RIS and UE, should lie within the Fresnel region [8]. Specifically, the distances should lie within the interval $[0.62\sqrt{D^{3}/\lambda},2D^{2}/\lambda]$, where $D=\sqrt{2}d(\sqrt{N_{R}}-1)$ represents the aperture of RIS and $\lambda$ is the operating wavelength. Additionally, we adopt a polar coordinate system for estimating the user’s position, i.e., using $d_{Ru}$, $\phi_{az}$, $\phi_{el}$ to express the user’s position, which represents the distance, azimuth angle, and elevation angle of UE relative to the coordinate origin, respectively. The transformation relationships between these three parameters and $x_{u}$, $y_{u}$, $z_{u}$ are as follows
$d_{Ru}$ $=\sqrt{x_{u}^{2}+y_{u}^{2}+z_{u}^{2}},$ (1a)
$\phi_{az}$ $=\arctan\frac{y_{u}}{x_{u}},$ (1b)
$\phi_{el}$ $=\arccos\frac{z_{u}}{\sqrt{x_{u}^{2}+y_{u}^{2}+z_{u}^{2}}},$ (1c)
and in turn, can be expressed as
$x_{u}$ $=d_{Ru}\sin\phi_{el}\cos\phi_{az},$ (2a)
$y_{u}$ $=d_{Ru}\sin\phi_{el}\sin\phi_{az},$ (2b)
$z_{u}$ $=d_{Ru}\cos\phi_{el}.$ (2c)
The localization process in the paper is performed using one OFDM symbol, with $N$ subcarriers, a center carrier frequency of $f_{c}$, corresponding to a wavelength of $\lambda$, and a signal bandwidth of $B$. Upon sampling at a rate of $T_{s}=\frac{1}{NB}$ and removing the cyclic prefix (CP) at UE, the received time-domain signal can be represented as
$\bm{y}=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\tilde{\bm{h}}+\tilde{\bm{v}},$ (3)
where
  • ${\bm{y}\triangleq\left[\bm{y}[0],\bm{y}[1],\cdots,\bm{y}[N-1]\right]}^{\top}$,
  • $P$ is the transmit power,
  • $\bm{\Lambda}_{\phi}$ is the impact caused by CFO on the signal. It is a diagonal matrix, and each item on the diagonal can be expressed as $\left[\bm{\Lambda}_{\phi}\right]_{n,n}=e^{j2\pi\frac{n\Delta f}{NB}}=e^{j2\pi\frac{n}{N}\phi}$, where $\phi=\frac{\Delta f}{B}$ is called the normalized CFO,
  • $\bm{\Lambda}_{\bm{\theta}}\triangleq{\rm Diag}\left[e^{j\bm{\theta}[0]},e^{j\bm{\theta}[1]},\cdots,e^{j\bm{\theta}[N-1]}\right]$ is the impact caused by PN on the signal, further details about which will be discussed in the later part of this section,
  • $\mathbf{S}\triangleq\sqrt{N}\mathbf{F}^{\rm H}{\rm Diag}(\bm{s})$, where $\mathbf{F}^{\rm H}$ is the inverse fast Fourier transform (IFFT) matrix and $\bm{s}$ is the frequency domain transmitted symbol sequence at the AN,
  • $\tilde{\bm{h}}\triangleq\bm{h}_{LoS}+\bm{h}_{NLoS}$ is the frequency domain channel, with $\bm{h}_{LoS}$ and $\bm{h}_{NLoS}$ representing the LoS and the non-line-of-sight (NLoS) part of the channel, respectively. We assume $\bm{h}_{NLoS}$ follows a known distribution $\bm{h}_{NLoS}\sim\mathcal{CN}(0,\sigma_{NLoS}^{2}\mathbf{I}_{N})$. Hence, The impact caused by $\bm{h}_{NLoS}$ on the signal can be combined with the additive Gaussian noise $\tilde{\bm{v}}$. Consequently, we can rewrite (3) as
    $\bm{y}=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\bm{h}+\bm{v},$ (4)
    where $\bm{h}$ and $\bm{v}$ represent the LoS part of the channel and the combined noise, respectively. $\bm{h}$ contains the known part of the channel from AN to RIS and the unknown part of the channel from RIS to UE. Each item of it can be expressed by $\bm{h}[n]=\sum_{r=1}^{N_{R}}\sqrt{\rho_{aR,n,r}}\sqrt{\rho_{Ru,n,r}}\bm{w}[r]e^{-j2\pi\frac{nB}{N}(\tau_{ar}+\tau_{ru})}$, where $\sqrt{\rho_{aR,n,r}}=\frac{\sqrt{G_{a}}\lambda_{n}}{4\pi d_{ar}}$ and $\sqrt{\rho_{Ru,n,r}}=\frac{\sqrt{G_{t}}\lambda_{n}}{4\pi d_{ru}}$ are the free-space path loss from AN to RIS, and from RIS to UE, respectively. $\bm{w}[r]$ is the phase shift of the $r$-th element of RIS, $\tau_{ar}=\frac{d_{ar}}{c}=\frac{\|\bm{p}_{r}-\bm{p}_{a}\|}{c}$ represents the delay from AN to the $r$-th element of RIS, while $\tau_{ru}=\frac{d_{ru}}{c}=\frac{\|\bm{p}_{u}-\bm{p}_{r}\|}{c}$ denotes the delay from the $r$-th element of RIS to UE, with $c$ representing the light speed,
  • $\bm{v}\triangleq\left[\bm{v}[0],\bm{v}[1],\cdots,\bm{v}[N-1]\right]$ is the combined additive noise at UE, and follows the distribution $\bm{v}\sim\mathcal{CN}(\bm{0}_{N\times 1},\sigma^{2}\mathbf{I}_{N})$.
Drawing upon the attributes of PN observed in practical oscillators, it can be effectively modeled as a Wiener process [22], i.e.,
$\bm{\theta}[n]=\bm{\theta}[n-1]+\bm{\Delta}[n],$ (5)
where $\bm{\Delta}[n]$ is a real Gaussian variable with a distribution of $\bm{\Delta}[n]\sim\mathcal{N}(0,\sigma_{\Delta}^{2})$. As in [21], it is assumed that $\bm{\theta}[-1]=0$, and then we can obtain that the PN vector $\bm{\theta}\triangleq\left[\bm{\theta}[0],\bm{\theta}[1],\cdots,\bm{\theta}[N-1]\right]^{\top}$ follows a distribution of $\bm{\theta}\sim\mathcal{N}(\bm{0}_{N\times 1},\bm{\Psi})$, where $\mathbf{\Psi}$ can be expressed by
$\mathbf{\Psi}=\sigma_{\Delta}^{2}\begin{bmatrix}1&1&1&\cdots&1&1\\ 1&2&2&\cdots&2&2\\ 1&2&3&3&\cdots&3\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 1&2&3&\cdots&N-1&N\end{bmatrix}.$ (6)
III Proposed Joint Position, CFO and PN Estimation
In previous works [12, 16], localization was based on frequency-domain signals. However, in our problem, the introduction of CFO and PN would result in ICI if the signal is transformed into the frequency domain, which will increase the difficulty of our estimation. Therefore, in this paper, we conduct the estimation in the time domain. In this section, we propose an iterative algorithm based on MAP criterion and GD algorithm, which jointly estimates the normalized CFO, PN, and position of UE.
By defining parameter vector $\bm{\lambda}\triangleq\left[\phi,\bm{\theta}^{\top},d_{Ru},\phi_{az},\phi_{el}\right]^{\top}$, whereas $\bm{\theta}$ is a random variable, and the others are deterministic variables, the posterior distribution of $\bm{\lambda}$ given the received signal $\bm{y}$ can be written as
$p(\bm{\lambda}\mid\bm{y})=\frac{p(\bm{y}\mid\bm{\lambda})\times p(\bm{\theta})}{p(\bm{y})}.$ (7)
After removing the constant terms, the object function can be expressed as $\mathcal{L}(\bm{\lambda})=-\log p(\bm{y}\mid\bm{\lambda})-\log p(\bm{\theta})$, and our goal is to achieve the joint estimation of $\bm{\lambda}$ by optimizing the following unconstrained function.
$\hat{\bm{\lambda}}\propto\arg\min_{\bm{\lambda}}\mathcal{L}(\bm{\lambda}).$ (8)
III-A Phase Noise Estimation
III-A1 Approximation of PN vector
Figure 2: Eigenvalues of $\bm{\Psi}$ at $N=32$ and $\sigma_{\Delta}^{2}=10^{-3}$.
From (4), it can be observed that there are multiple parameters that need to be jointly estimated. In order to alleviate the computational complexity at UE, we take advantage of the properties of the PN vector to approximate it. This approach results in a reduction of unknown parameters that require estimation. As depicted in Fig. 2, the eigenvalues of $\bm{\Psi}$ mostly cluster around zero. This observation allows us to approximate the PN vector as follows
$\bm{\theta}\approx\bm{\Pi}\bm{\eta},$ (9)
where $\bm{\eta}\sim\mathcal{N}(\bm{0}_{L\times 1},\mathbf{I}_{L})$ contains $L\leq N$ PN parameters, and $\bm{\Pi}\in\mathbb{R}^{N\times L}$ is the transformation matrix. To derive $\bm{\Pi}$, we first perform singular value decomposition (SVD) on $\bm{\Psi}$, which is given by $\bm{\Psi}=\mathbf{U}\mathbf{D}\mathbf{U}^{\top}$, where $\mathbf{D}={\rm Diag}(\bm{e})$, and $\bm{e}$ represents the eigenvalues of $\bm{\Psi}$ arranged in descending order. Subsequently, $\bm{\Pi}$ can be calculated by $\bm{\Pi}=[\mathbf{U}]_{:,0:L-1}\times\sqrt{\bm{e}}$. In the subsequent analysis, $\bm{\eta}$ is estimated instead of $\bm{\theta}$.
III-A2 Estimation of approximate PN vector
After approximating the PN parameter $\bm{\theta}$, the objective function can be rewritten as
$\mathcal{L}(\bm{\lambda})$ $=-\log p(\bm{y}\mid\bm{\lambda})-\log p(\bm{\eta})$
$\propto\frac{1}{\sigma^{2}}\|\bm{y}-\bm{\mu}\|^{2}+\frac{1}{2}\bm{\eta}^{\top}\bm{\eta},$ (10)
where $\bm{\mu}\triangleq\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\bm{h}$ is the received signal at UE without additive noise, and the noise variance $\sigma^{2}$ is assumed to be known.
Due to the typically small PN variance in practical oscillators, the elements of $\bm{\Lambda}_{\bm{\theta}}$ can be approximated by employing a first-order Taylor series expansion, resulting in $e^{j\bm{\theta}[n]}\approx 1+j\bm{\theta}[n]$. Consequently, we can approximate $\bm{\Lambda}_{\bm{\theta}}$ as $\bm{\Lambda}_{\bm{\theta}}\approx\mathbf{I}_{N}+j{\rm Diag}(\bm{\Pi\eta})$. With this approximation, the objective function in (III-A2) can be reformulated as:
$\mathcal{L}_{\bm{\eta}}^{[q+1]}$ $\approx\frac{1}{\sigma^{2}}(\bar{\bm{y}}-\mathbf{Q}\bm{\eta})^{\rm H}(\bar{\bm{y}}-\mathbf{Q}\bm{\eta})+\frac{1}{2}\bm{\eta}^{\top}\bm{\eta}$
$=\frac{1}{\sigma^{2}}\left(\bar{\bm{y}}^{\rm H}\bar{\bm{y}}-2\Re\left\{\bar{\bm{y}}^{\rm H}\mathbf{Q}\right\}\bm{\eta}+\bm{\eta}^{\top}\Re\left\{\mathbf{Q}^{\rm H}\mathbf{Q}\right\}\bm{\eta}\right)$
$\quad+\frac{1}{2}\bm{\eta}^{\top}\bm{\eta},$ (11)
where $\bar{\bm{y}}\triangleq\bm{y}-\sqrt{P}\hat{\bm{\Lambda}}_{\phi}^{[q]}\mathbf{S}^{\top}\hat{\bm{h}}^{[q]}$, and $\mathbf{Q}\triangleq j\sqrt{P}{\rm Diag}(\hat{\bm{\Lambda}}_{\phi}^{[q]}\mathbf{S}^{\top}\hat{\bm{h}}^{[q]})\bm{\Pi}$. The superscript $q$ represents the round number of iterations.
Subsequently, by setting the gradient of (III-A2) equal to zero, i.e.,
$\frac{\partial\mathcal{L}^{[q+1]}_{\bm{\eta}}}{\partial\bm{\eta}}=-\frac{2}{\sigma^{2}}\Re\left\{\mathbf{Q}^{\rm H}\bar{\bm{y}}\right\}+\frac{2}{\sigma^{2}}\Re\left\{\mathbf{Q}^{\rm H}\mathbf{Q}\right\}\bm{\eta}+\bm{\eta}=\bm{0}_{L\times 1},$ (12)
we obtain
$\hat{\bm{\eta}}^{[q+1]}=\left[\Re\left\{\mathbf{Q}^{\rm H}\mathbf{Q}\right\}+\frac{1}{2}\sigma^{2}\mathbf{I}_{L}\right]^{-1}\Re\left\{\mathbf{Q}^{\rm H}\bar{\bm{y}}\right\}.$ (13)
III-B CFO Estimation
To estimate the CFO, we solve the following unconstrained problem
$\hat{\phi}\propto\arg\min_{\phi}\mathcal{L}_{\phi}\propto\arg\min_{\phi}\|\bm{y}-\bm{\mu}\|^{2}.$ (14)
In order to enhance the tractability of the problem presented in (14) and obtain a closed-form solution, we employ a Taylor series approximation once more. The $n$-th term on the diagonal of $\bm{\Lambda}_{\phi}$ can be approximated as
$\left[\bm{\Lambda}_{\phi}\right]_{n,n}=e^{j2\pi\frac{n}{N}\phi}\approx e^{j2\pi\frac{n}{N}\hat{\phi}}+\left(\phi-\hat{\phi}\right)\frac{j2\pi}{N}e^{j2\pi\frac{n}{N}\hat{\phi}}.$ (15)
Hence, $\hat{\bm{\Lambda}}_{\phi}^{[q+1]}$ can be approximated as
$\hat{\bm{\Lambda}}_{\phi}^{[q+1]}\approx\hat{\bm{\Lambda}}_{\phi}^{[q]}+\left(\phi-\hat{\phi}^{[q]}\right)\tilde{\bm{\Lambda}}_{\phi}^{[q]},$ (16)
where $\left[\tilde{\bm{\Lambda}}_{\phi}^{[q]}\right]_{n,n}=\frac{j2\pi n}{N}e^{\frac{j2\pi n\hat{\phi}^{[q]}}{N}}$.
With (16), $\mathcal{L}_{\phi}$ in (14) can be rewritten as
$\mathcal{L}_{\phi}^{[q+1]}\approx\frac{1}{\sigma^{2}}\|{\bm{y}-\left[\hat{\bm{\Lambda}}_{\phi}^{[q]}+\left(\phi-\hat{\phi}^{[q]}\right)\tilde{\bm{\Lambda}}_{\phi}^{[q]}\right]\bm{d}}\|^{2},$ (17)
where $\bm{d}\triangleq\sqrt{P}\hat{\bm{\Lambda}}_{\bm{\theta}}^{[q+1]}\mathbf{S}^{\top}\hat{\bm{h}}^{[q]}$. Upon setting the gradient of (17) equal to zero, and solving for $\phi$, we can obtain the closed-form expression of $\hat{\phi}$ at the $(q+1)$-th iteration as
$\hat{\phi}^{[q+1]}=\hat{\phi}^{[q]}+\frac{\Re\left\{\left(\bm{y}-\hat{\bm{\Lambda}}_{\phi}^{[q]}\bm{d}\right)^{\rm H}\tilde{\bm{\Lambda}}_{\phi}^{[q]}\bm{d}\right\}}{\bm{d}^{\rm H}\tilde{\bm{\Lambda}}_{\phi}^{[q]{\rm H}}\tilde{\bm{\Lambda}}_{\phi}^{[q]}\bm{d}}.$ (18)
III-C Position Estimation
For position estimation, we employ the GD algorithm. Due to the similarity in the derivative expressions of the objective function with respect to three positional parameters $d_{Ru}$, $\phi_{az}$, and $\phi_{el}$, we use the symbol $\xi$ to represent any one of them. The expression for the partial derivative of the objective function with respect to $\xi$ is as follows
$\frac{\partial\mathcal{L}_{\xi}}{\partial\xi}=\frac{2}{\sigma^{2}}\Re\left\{\left(P\bm{h}^{\rm H}\mathbf{S}^{*}\mathbf{S}^{\top}-\sqrt{P}\bm{y}^{\rm H}\bm{\Lambda}_{\bm{\theta}}\bm{\Lambda}_{\phi}\mathbf{S}^{\top}\right)\frac{\partial\bm{h}}{\partial\xi}\right\},$ (19)
where each term of $\frac{\partial\bm{h}}{\partial\xi}$ can be calculated as follows
$\frac{\partial\bm{h}}{\partial\xi}[n]$ $=\frac{\partial\bm{h}[n]}{\partial\xi}=-\sum_{r=1}^{N_{R}}\sqrt{\rho_{aR,n,r}}\sqrt{\rho_{Ru,n,r}}\bm{w}[r]$
×$\times e^{-j2\pi\frac{nB}{N}(\tau_{ar}+\tau_{ru})}\left(\frac{1}{d_{ru}}+j2\pi\frac{nB}{Nc}\right)\frac{\partial d_{ru}}{\partial\xi},$ (20)
with $\frac{\partial d_{ru}}{\partial\xi}$, $\xi=d_{Ru},\phi_{az},\phi_{el}$, being
$\frac{\partial d_{ru}}{\partial d_{Ru}}$ $=\frac{1}{d_{Ru}}\left[(d_{Ru}\sin\phi_{el}\cos\phi_{az}-x_{r})\sin\phi_{el}\cos\phi_{az}\right.$
$\left.\quad+(d_{Ru}\sin\phi_{el}\sin\phi_{az}-y_{r})\sin\phi_{el}\sin\phi_{az}\right.$
$\left.\quad+(d_{Ru}\cos\phi_{el}-z_{r})\cos\phi_{el}\right],$ (21)
$\frac{\partial d_{ru}}{\partial\phi_{az}}$ $=\frac{1}{d_{Ru}}\left[(d_{Ru}\sin\phi_{el}\cos\phi_{az}-x_{r})(-\sin\phi_{az}d_{Ru}\sin\phi_{el})\right.$
$\left.\quad+(d_{Ru}\sin\phi_{el}\sin\phi_{az}-y_{r})d_{Ru}\sin\phi_{el}\cos\phi_{az}\right],$ (22)
and
$\frac{\partial d_{ru}}{\partial\phi_{el}}$ $=\frac{1}{d_{Ru}}\left[(d_{Ru}\sin\phi_{el}\cos\phi_{az}-x_{r})d_{Ru}\cos\phi_{az}\cos\phi_{el}\right.$
$\left.\quad+(d_{Ru}\sin\phi_{el}\sin\phi_{az}-y_{r})d_{Ru}\sin\phi_{az}\cos\phi_{el}\right.$
$\left.\quad+(d_{Ru}cos\phi_{el}-z_{r})(-d_{Ru}\sin\phi_{el})\right].$ (23)
The general iterative joint estimation algorithm is outlined in Algorithm 1.
Algorithm 1 Iterative joint estimation algorithm.
1:  Optimize RIS phase shift $\bm{w}$ using the method to be introduced in Section V.
2:  Initialize estimation of normalized CFO $\hat{\phi}$ and UE position $\hat{\bm{p}}_{u}$, step length $k_{p}$, convergence accuracy for outer loop $\epsilon_{o}$, convergence accuracy for inner loop $\epsilon_{i}$, and calculate corresponding polar coordinate of $\hat{\bm{p}}_{u}$ using (1).
3:  repeat
4:     Update $\hat{\bm{\eta}}$ with $\hat{\phi}$, $\hat{\bm{p}}_{u}$ being fixed using (13);
5:     Update $\hat{\bm{\theta}}$ using (9);
6:     Update $\hat{\phi}$ with $\hat{\bm{\theta}}$, $\hat{\bm{p}}_{u}$ being fixed using (18);
7:     repeat
8:        $\hat{d}_{Ru}\leftarrow\hat{d}_{Ru}-\frac{\partial\mathcal{L}_{\xi}}{\partial d_{Ru}}k_{p}$;
9:        $\hat{\phi}_{az}\leftarrow\hat{\phi}_{az}-\frac{\partial\mathcal{L}_{\xi}}{\partial\phi_{az}}k_{p}$;
10:        $\hat{\phi}_{el}\leftarrow\hat{\phi}_{el}-\frac{\partial\mathcal{L}_{\xi}}{\partial\phi_{el}}k_{p}$;
11:     until $\mathcal{L}^{[p]}-\mathcal{L}^{[p+1]}\leq\epsilon_{i}$ after the $p$-th iteration.
12:  until $\mathcal{L}^{[q]}-\mathcal{L}^{[q+1]}\leq\epsilon_{o}$ after the $q$-th iteration. $\epsilon_{i}$ and $\epsilon_{o}$ are both pre-set convergence accuracy and should vary with changes in transmit power.
13:  Calculate Cartesian coordinate of $\hat{\bm{p}}_{u}$ using (2).
III-D Estimation Performance Evaluation
For position estimation, we use root mean square error (RMSE) as the performance metric. For any testing UE position $\bm{p}_{u}$, ${\rm RMSE}_{u}$ can be expressed as
${\rm RMSE}_{u}=\sqrt{\frac{1}{M_{e}}\sum_{m=1}^{M_{e}}\|\hat{\bm{p}}_{u,m}-\bm{p}_{u}\|^{2}},$ (24)
where $M_{e}$ is the Monte Carlo experiment amount, and $\hat{\bm{p}}_{u,m}$ represents the estimation of UE position in the $m$-th trail.
For CFO and PN estimation, it can be straightforwardly observed from (4) that, the received signal $\bm{y}$ is not altered under a common phase rotation $\epsilon$, i.e.,
$\hat{\phi}\rightarrow\phi-\epsilon,\hat{\bm{\theta}}\rightarrow\bm{\theta}+\bm{\epsilon},$ (25)
where $\bm{\epsilon}[n]=\frac{2\pi n\epsilon}{N}$. Equation (25) indicates that there exists ambiguity in estimating CFO and PN. The ambiguity leads to difficulties in achieving high estimation accuracy for individual parameters of CFO and PN. However, in this paper, we are not focusing on the separate estimation accuracy of CFO and PN but rather on the estimation accuracy of UE position. Therefore, it is acceptable as long as the overall estimation accuracy of CFO and PN is reduced. In this paper, to evaluate the overall estimation accuracy of CFO and PN, we utilize a novel joint mean square error (MSE) measure, which can be calculated as
${\rm MSE}_{\phi,\bm{\theta}}=\mathbb{E}_{\phi,\bm{\theta}}\left(\|\bm{\underline{\gamma}}-\hat{\bm{\underline{\gamma}}}\|^{2}\right),$ (26)
where $\bm{\underline{\gamma}}=\bm{\gamma}-\bm{\gamma}[0]\bm{1}$, $\hat{\bm{\underline{\gamma}}}=\hat{\bm{\gamma}}-\hat{\bm{\gamma}}[0]\bm{1}$, $\bm{\gamma}=[\bm{\gamma}[0],\bm{\gamma}[1],\cdots,\bm{\gamma}[N-1]]^{\top}$, $\bm{\gamma}[n]=\bm{\theta}[n]+\frac{2\pi n\phi}{N}$, $\hat{\bm{\gamma}}=[\hat{\bm{\gamma}}[0],\hat{\bm{\gamma}}[1],\cdots,\hat{\bm{\gamma}}[N-1]]^{\top}$, $\hat{\bm{\gamma}}[n]=\hat{\bm{\theta}}[n]+\frac{2\pi n\hat{\phi}}{N}$.
IV Hybrid Cramér-Rao Lower Bound Analysis
In this section, we derive HCRLB to evaluate the joint estimation, in the presence of a random parameter $\bm{\theta}$ in the set of parameters to be estimated. The HCRLB for $\bm{\lambda}$, the parameters to be estimated, is given by [7]
$\mathbb{E}_{\bm{y},\bm{\theta}\mid\phi,d_{Ru},\phi_{az},\phi_{el}}\left[(\bm{\lambda}-\hat{\bm{\lambda}})(\bm{\lambda}-\hat{\bm{\lambda}})^{\top}\right]\succeq\mathbf{B}^{\rm-1},$ (27)
where $\mathbf{B}$ is the Bayesian information matrix (BIM), which can be calculated as
$\mathbf{B}=\mathbb{E}_{\bm{\theta}}\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]+\mathbb{E}_{\bm{\theta}}\left[-\frac{\partial}{\partial\bm{\lambda}}\left(\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}\right)^{\top}\right].$ (28)
In order to evaluate the positioning accuracy and the CFO-PN joint MSE, we define a new parameter vector $\bm{\alpha}=\bm{g}(\bm{\lambda})\triangleq\left[\bm{\underline{\gamma}},x_{u},y_{u},z_{u}\right]^{\top}$. In the following subsections, we first derive the expression for matrix $\mathbf{B}$. Then, we derive the transition matrix to transform $\mathbf{B}$ into the HCRLB relative to the parameter vector $\bm{\alpha}$.
IV-A Derivation of $\mathbb{E}_{\bm{\theta}}\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]$
To derive $\mathbb{E}_{\bm{\theta}}\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]$, we begin by calculating the Fisher information matrix (FIM) for $\bm{\lambda}$, as expressed by
$\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]_{i,j}=\frac{2}{\sigma^{2}}\Re\left\{\frac{\partial\bm{\mu}{\rm{}^{\rm H}}}{\partial\lambda_{i}}\frac{\partial\bm{\mu}}{\partial\lambda_{j}}\right\}.$ (29)
To calculate (29), the following gives the derivative of $\bm{\mu}$ with respect to each element in vector $\bm{\lambda}$.
$\frac{\partial\bm{\mu}}{\partial\phi}$ $=\sqrt{P}\bm{\Lambda}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\bm{h},$ (30a)
$\frac{\partial\bm{\mu}}{\partial\bm{\theta}[n]}$ $=\sqrt{P}{\rm Diag}(\bm{a}[n])\bm{\Lambda}_{\phi}\mathbf{S}^{\top}\bm{h},$ (30b)
$\frac{\partial\bm{\mu}}{\partial d_{Ru}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial d_{Ru}},$ (30c)
$\frac{\partial\bm{\mu}}{\partial\phi_{az}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial\phi_{az}},$ (30d)
$\frac{\partial\bm{\mu}}{\partial\phi_{el}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial\phi_{el}},$ (30e)
where $\bm{\Lambda}$ is a diagnal matrix, $\left[\bm{\Lambda}\right]_{n,n}=\frac{j2\pi n}{N}$, $\bm{a}[n]=[\bm{0}_{1\times n},je^{j\bm{\theta}[n]},\bm{0}_{1\times(N-n-1)}]^{\top}$.
It is not difficult to observe that in (30), the terms related to $\phi$ and $\bm{\theta}$ all appear in the exponential terms of the diagonal elements of the corresponding matrix. Hence, they would disappear when calculating $\frac{\partial\bm{\mu}{\rm{}^{\rm H}}}{\partial\lambda_{i}}\frac{\partial\bm{\mu}}{\partial\lambda_{j}}$ in (29). That is to say, the calculated FIM is not related to parameter $\bm{\theta}$, and
$\mathbb{E}_{\bm{\theta}}\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]=\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda}).$ (31)
IV-B Derivation of $\mathbb{E}_{\bm{\theta}}\left[-\frac{\partial}{\partial\bm{\lambda}}\left(\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}\right)^{\top}\right]$
Since $\bm{\theta}\sim\mathcal{N}(\bm{0}_{N\times 1},\bm{\Psi})$, we can calculate $\log p(\bm{\theta})$ by $\log p(\bm{\theta})=-\frac{1}{2}\bm{\theta}^{\top}\bm{\Psi}^{-1}\bm{\theta}-\frac{N}{2}\log(2\pi)-\frac{1}{2}\log\det\bm{\Psi}$, and its partial to $\bm{\lambda}$ can be calculated by
$\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}=\left[0,-(\bm{\Psi}^{-1}\bm{\theta})^{\top},0,0,0\right]^{\top}.$ (32)
Subsequently, $\mathbb{E}_{\bm{\theta}}\left[-\frac{\partial}{\partial\bm{\lambda}}\left(\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}\right)^{\top}\right]$ can be expressed by
$\mathbb{E}_{\bm{\theta}}\left[-\frac{\partial}{\partial\bm{\lambda}}\left(\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}\right)^{\top}\right]$ $=-\frac{\partial}{\partial\bm{\lambda}}\left(\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}\right)^{\top}$
$=\mathrm{Blkdiag}(0,\bm{\Psi}^{-1},\bm{0}_{3\times 3}).$ (33)
IV-C Derivation of the Transformed HCRLB
Due to the ambiguity between the parameters to be estimated, the MSE of CFO and PN is jointly calculated using (26), and when computing the position error bound (PEB), the position parameters are required to be in the form of 3D Cartesian coordinates $x_{u}$, $y_{u}$, and $z_{u}$. Therefore, the parameter vector $\bm{\lambda}$ needs to be transformed into $\bm{\alpha}$. The HCRLB for $\bm{\alpha}$ is given by [9]
$\mathbf{HCRLB}(\bm{p}_{u},\bm{w})=\bm{\Xi}(\bm{p}_{u})\mathbf{B}^{-1}(\bm{p}_{u},\bm{w})\bm{\Xi}^{\top}({\bm{p}_{u}}),$ (34)
where for any UE position $\bm{p}_{u}$, $\mathbf{\Xi}(\bm{p}_{u})$ is the transition matrix defined as
$\bm{\Xi}(\bm{p}_{u})=\frac{\partial\bm{g}(\bm{\lambda})}{\partial\bm{\lambda}}={\rm Blkdiag}\left(\bm{\Xi}_{1},\bm{\Xi}_{2}(\bm{p}_{u})\right),$ (35)
and $\bm{\Xi}_{1}$, $\bm{\Xi}_{2}(\bm{p}_{u})$ can be calculated by (36) and (37).
$\bm{\Xi}_{1}=\begin{bmatrix}\frac{\partial\bm{\underline{\gamma}}[0]}{\partial\phi}&\frac{\partial\bm{\underline{\gamma}}[0]}{\partial\bm{\theta}[0]}&\cdots&\frac{\partial\bm{\underline{\gamma}}[0]}{\partial\bm{\theta}[N-1]}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{\partial\bm{\underline{\gamma}}[N-1]}{\partial\phi}&\frac{\partial\bm{\underline{\gamma}}[N-1]}{\partial\bm{\theta}[0]}&\cdots&\frac{\partial\bm{\underline{\gamma}}[N-1]}{\partial\bm{\theta}[N-1]}\end{bmatrix}=\begin{bmatrix}0&0&0&\cdots&0\\ \frac{2\pi}{N}&-1&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \frac{2\pi(N-1)}{N}&-1&0&\cdots&1\\ \end{bmatrix}\in\mathbb{R}^{N\times(N+1)},$ (36)
$\bm{\Xi}_{2}(\bm{p}_{u})=\begin{bmatrix}\frac{\partial x_{u}}{\partial d_{Ru}}&\frac{\partial x_{u}}{\partial\phi_{az}}&\frac{\partial x_{u}}{\partial\phi_{el}}\\ \frac{\partial y_{u}}{\partial d_{Ru}}&\frac{\partial y_{u}}{\partial\phi_{az}}&\frac{\partial y_{u}}{\partial\phi_{el}}\\ \frac{\partial z_{u}}{\partial d_{Ru}}&\frac{\partial z_{u}}{\partial\phi_{az}}&\frac{\partial z_{u}}{\partial\phi_{el}}\\ \end{bmatrix}=\begin{bmatrix}\sin\phi_{el}\cos\phi_{az}&-d_{Ru}\sin\phi_{el}\sin\phi_{az}&d_{Ru}\cos\phi_{az}\cos\phi_{el}\\ \sin\phi_{el}\sin\phi_{az}&d_{Ru}\sin\phi_{el}\cos\phi_{az}&d_{Ru}\sin\phi_{az}\cos\phi_{el}\\ \cos\phi_{el}&0&-d_{Ru}\sin\phi_{el}\end{bmatrix}\in\mathbb{R}^{3\times 3}.$ (37)
After obtaining $\mathbf{HCRLB}(\bm{p}_{u},\bm{w})$, $\rm PEB$ for position estimation and $\rm HCRLB_{CFO-PN}$ for CFO-PN estimation can be calculated as
${\rm PEB}=\sqrt{{\rm tr}([\mathbf{HCRLB}(\bm{p}_{u},\bm{w})]_{N:N+2,N:N+2})},$ (38)
${\rm HCRLB_{CFO-PN}}={\rm tr}([\mathbf{HCRLB}(\bm{p}_{u},\bm{w})]_{0:N-1,0:N-1}).$ (39)
Both ${\rm PEB}$ and ${\rm HCRLB_{CFO-PN}}$ are related to UE’s location $\bm{p}_{u}$ and RIS phase shift $\bm{w}$.
V RIS Phase Shift Optimization
In this section, our goal is to optimize the phase shift of the RIS to enhance theoretical positioning accuracy, i.e., PEB. According to equation (38), minimizing the PEB at a specific location requires knowledge of that location. However, this is not possible in practical positioning scenarios. To address this issue, we select multiple sampling points within the AOI randomly, then calculate and minimize their average PEB. Accordingly, the optimization problem for RIS phase shift can be expressed as
$\min_{\bm{w}}\quad\frac{1}{U}\sum_{u=1}^{U}{\rm PEB}(\bm{p}_{u},\bm{w})$
$s.t.\quad|\bm{w}[r]|=1,r=1,\cdots,N_{R}.$ (40)
Problem (V) is non-convex because the objective function does not exhibit convexity with respect to vector $\bm{w}$. We define $\mathbf{W}\triangleq\bm{\bar{w}}\bm{\bar{w}}^{\rm H}$, where $\bm{\bar{w}}\triangleq\bm{w}^{*}$, which needs to satisfy $\mathbf{W}\succcurlyeq 0$ and ${\rm rank}(\mathbf{W})=1$. Since the rank-1 constraint is non-convex, we employ the SDR method to relax this constraint. Consequently, the problem (V) can be transformed into
$\min_{\mathbf{W}}\quad\frac{1}{U}\sum_{u=1}^{U}{\rm PEB}(\bm{p}_{u},\mathbf{W})$
$\begin{array}[]{ll}s.t.&\left[\mathbf{W}\right]_{r,r}=1,r=1,\cdots,N_{R},\\ &\mathbf{W}\succcurlyeq 0.\end{array}$ (43)
Inspired by [12], we utilize the property of the Schur complement [6], and the problem can be expressed in the form of SDP. For each calculated PEB at a particular position $\bm{p}_{u}$, we define an auxiliary matrix $\mathbf{Z}_{u}$ that satisfies the following expression
$\mathbf{Z}_{u}\succcurlyeq\tilde{\bm{\Xi}}(\bm{p}_{u})\mathbf{B}^{-1}(\bm{p}_{u},\mathbf{W})\tilde{\bm{\Xi}}^{\top}(\bm{p}_{u}),$ (44)
where $\tilde{\bm{\Xi}}(\bm{p}_{u})\triangleq\mathbf{T}\mathbf{\Xi}(\bm{p}_{u})$, and $\mathbf{T}\triangleq[\bm{0}_{N\times N}\quad\mathbf{I}_{3}]\in\mathbb{R}^{(N+3)\times(N+3)}$ can extract the PEB portion from the HCRLB.
The optimization problem in (V) can be converted to
$\min_{\mathbf{W},\mathbf{Z}_{1},\cdots,\mathbf{Z}_{U}}\quad\frac{1}{U}\sum_{u=1}^{U}{\rm tr}(\mathbf{Z}_{u})$
$\begin{array}[]{ll}s.t.&\left[\mathbf{W}\right]_{r,r}=1,r=1,\cdots,N_{R},\\ &\mathbf{W}\succcurlyeq 0,\\ &\displaystyle\frac{1}{U}\sum_{u=1}^{U}\begin{bmatrix}\mathbf{Z}_{u}&\tilde{\bm{\Xi}}(\bm{p}_{u})\\ \tilde{\bm{\Xi}}^{\top}(\bm{p}_{u})&\mathbf{B}(\bm{p}_{u},\mathbf{W})\end{bmatrix}\succcurlyeq 0.\end{array}$ (48)
Problem (V) is a convex SDP problem, and can be optimally solved using established convex optimization solvers, such as CVX [30]. Since there is no rank-1 constraint in the optimization problem, it is highly likely that the obtained matrix $\mathbf{W}$ will have a rank greater than one. Therefore, decomposing it into the desired RIS phase shift vector $\bm{w}$ is a crucial subsequent problem. As in [29], we first perform an eigenvalue decomposition on the optimal $\mathbf{W}$, then take the largest eigenvector and multiply it by the square root of the largest eigenvalue and normalize the vector, obtaining $\bm{\bar{w}}$. Finally, we calculate the RIS phase shift by $\bm{w}=\bm{\bar{w}}^{*}$. The decomposition process can also be accomplished using the Gaussian randomization method as in [28].
In order to meet the requirements of (V), we next derive the expression of $\mathbf{B}$ that is related to $\mathbf{W}$. We first rewrite (4) as
$\bm{y}=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\mathbf{G}^{\top}\bm{w}+\bm{v},$ (49)
which separates $\bm{w}$ from $\bm{h}$, with $\mathbf{G}\in\mathbb{C}^{N_{R}\times N}$ and $\left[\mathbf{G}\right]_{r,n}=\sqrt{\rho_{aR,n,r}}\sqrt{\rho_{Ru,n,r}}e^{-j2\pi\frac{nB}{N}(\tau_{ar}+\tau_{ru})}$.
Correspondingly, $\bm{\mu}$ can be rewritten as $\bm{\mu}=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\mathbf{G}^{\top}\bm{\bar{w}}^{*}$. Furthermore, we can express each term of $\bm{\mu}$ and its conjugate as
$\bm{\mu}[n]=\bm{\mu}[n]^{\top}$ $=\sqrt{P}e^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}^{\top}]_{n,:}\mathbf{G}^{\top}\bm{\bar{w}}^{*}$
$=\sqrt{P}e^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}\bm{\bar{w}}^{\rm H}\mathbf{G}[\mathbf{S}]_{:,n},$ (50)
and
$\bm{\mu}[n]^{*}=\sqrt{P}e^{-j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}]_{:,n}^{\rm H}\mathbf{G}^{\rm H}\bm{\bar{w}}.$ (51)
According to (28), only the $\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})$ part is related to $\bm{w}$, hence we only need to derive the relationship of this part of $\mathbf{B}$ and $\mathbf{W}$. We can rewrite the FIM in a summation form:
$\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]_{i,j}=\frac{2}{\sigma^{2}}\sum_{n=0}^{N-1}\Re\left\{\frac{\partial\bm{\mu}[n]^{*}}{\partial\lambda_{i}}\frac{\partial\bm{\mu}[n]}{\partial\lambda_{j}}\right\}.$ (52)
The partial derivatives of $\bm{\mu}[n]$ and $\bm{\mu}[n]^{*}$ with respect to all parameters to be estimated are derived in Appendix A. It is observed that the expressions for $\frac{\partial\bm{\mu}[n]^{*}}{\partial\lambda_{i}}$ all end with $\bm{\bar{w}}$, while the expressions for $\frac{\partial\bm{\mu}[n]}{\partial\lambda_{j}}$ all start with $\bm{\bar{w}}^{\rm H}$ for the vector parts. Therefore, when calculating the FIM, we can easily obtain the expression with respect to $\mathbf{W}$.
VI Numerical Results
This section presents comprehensive simulations conducted to assess the performance of the proposed estimation and optimization algorithms. We consider a 3D Cartesian coordinate system, where the AN is located at $\bm{p}_{a}=[\rm 2m,-2m,1m]^{\top}$, and the reference point of RIS is located at $\bm{0}_{3\times 1}$. Without loss of generality, RIS is placed on the $yz$ plane. We assume AOI in a cube centered at $[\rm 2m,2m,0m]^{\top}$ with the edge length of $\rm 1m$. We set light speed $c=3\times 10^{8}{\rm m/s}$, noise power $\sigma^{2}=-109{\rm dBm}$. $N=32$ subcarriers are used in one OFDM symbol, and the central frequency and bandwidth of the OFDM symbol are set to $f_{c}=2.8{\rm GHz}$ and $B=100{\rm MHz}$, respectively. The normalized CFO $\phi$ is uniformly drawn from $[-0.15,0.15]$, and the length of shortened PN vector $L$ is set as $L=16$. Additionally, we make the assumption that the antenna gains, $G_{a}$ and $G_{t}$, are equal to $1$.
VI-A Convergence analysis
Figure 3: The convergence of Algorithm 1.
To begin, we analyze the convergence of the proposed joint estimation algorithm. Fig. 3 illustrates the convergence of the algorithm in the case of transmit power of $-10{\rm dBm}$, which corresponds to a receiving signal-to-noise ratio (SNR) of approximately $17.5{\rm dB}$. Specifically, Fig. 3 (a) shows the variation of the objective function during each inner GD iteration in Algorithm 1, while Fig. 3 (b) focuses solely on the variation of the objective function based on the results after every inner iteration. From Fig. 3 (a), we observe that the inner GD iterations corresponding to each outer iteration converge earlier and earlier. This phenomenon occurs because the convergence for outer iteration approaches, the magnitude of the parameter updates decreases, and the variation of the posterior function becomes closer to the predefined stopping threshold. From Fig. 3, it is not difficult to notice that the inner loop typically converges within 2000 iterations, while the outer loop generally converges within 40 iterations.
VI-B Performance of RIS Phase Shift Optimization
Figure 4: Apply random RIS phase shift and calculate the PEB of different positions.
Figure 5: Apply the RIS phase shift optimization algorithm and calculate the PEB of different positions.
In this subsection, our attention is directed toward evaluating the performance of the optimization algorithm. Fig. 4 and Fig. 5 illustrate the PEB near the AOI before and after the optimization of RIS phase shift, respectively. More precisely, we configure the number of RIS elements, denoted as $N_{R}$, to be $81$, and transmit power $P$ to be $-20{\rm dBm}$, which corresponds to a receiving SNR of approximately $7.5{\rm dB}$. Without loss of generality, the selected positions in the figures are located in the $xy$ plane. A comparison between the two figures reveals that the optimized PEB has improved by approximately two orders of magnitude, which fully demonstrates the effectiveness of the employed RIS optimization algorithm. Additionally, we have observed that the optimized RIS phase shift not only improves the theoretical positioning accuracy within the AOI, but also has a significant improvement in the surrounding areas of it.
VI-C Localization Accuracy
Figure 6: The CDF of RMSE and PEB with different $N_{R}$ and transmit powers.
Figure 7: The RMSE of $\bm{p}_{u}$ estimation with random and optimized RIS phase shift.
Figure 8: The RMSE of $\bm{p}_{u}$ estimation at $N_{R}=49$, $N_{R}=81$, and $N_{R}=121$.
Figure 9: ${\rm MSE}_{\phi,\bm{\theta}}$ at $N_{R}=49$, $N_{R}=81$, and $N_{R}=121$.
Figure 10: ${\rm RMSE}_{u}$ at $\sigma_{\Delta}^{2}=10^{-3}{\rm rad^{2}}$ and $\sigma_{\Delta}^{2}=10^{-4}{\rm rad^{2}}$.
Figure 11: ${\rm MSE}_{\phi,\bm{\theta}}$ at $\sigma_{\Delta}^{2}=10^{-3}{\rm rad^{2}}$ and $\sigma_{\Delta}^{2}=10^{-4}{\rm rad^{2}}$.
In this subsection, we first illustrate the overall positioning accuracy. Fig. 6 presents the cumulative distribution functions (CDFs) of RMSE and PEB for 500 random testing positions, under different transmit powers and the number of RIS elements. It can be observed that increasing the RIS size and the transmit power both lead to an improvement in the overall positioning accuracy. Furthermore, we find that when the RIS size is set to $N_{R}=49$, the CDF of RMSE deviates significantly from the CDF of PEB, and doesn’t achieve the theoretical positioning accuracy. However, when the RIS size is increased to $N_{R}=121$, the CDF of the RMSE curve aligns more closely with the CDF of PEB. This phenomenon can also be observed in Fig. 8.
Next, we select a fixed testing position at coordinate $\rm[1.50m,2.15m,0.45m]^{\top}$ (rounded to two decimal places for convenience) to investigate the impact of transmit power $P$, RIS size $N_{R}$, and PN variance $\sigma_{\Delta}^{2}$ on the estimation accuracy of UE position, CFO, and PN.
In Fig. 7, we fix the number of RIS elements at 81 and assess the impact of random RIS phase shift versus optimized RIS configuration on positioning accuracy. It is evident that the actual positioning performance with random RIS phase shift does not experience a significant decline as the transmit power increases. At the other extreme, the optimized RIS phase shift provides an improvement of approximately two orders of magnitude in positioning accuracy. This fully demonstrates the crucial role of optimizing the RIS phase shift in localization problems and the effectiveness of the optimization method we employed.
Fig. 8 and Fig. 9 respectively depict the curves of RMSE for positioning and MSE for estimation of $\underline{\bm{\gamma}}$ under different transmit powers, at the number of RIS sizes of 49, 81, and 121. It can be observed that increasing the transmit power and the number of RIS elements results in gains in all parameters to be estimated. This demonstrates that increasing the signal transmit power or increasing the number of RIS reflection elements is beneficial for parameter estimation.
Fig. 10 and Fig. 11 respectively illustrate the RMSE for positioning and MSE for estimation of $\underline{\bm{\gamma}}$ when $N_{R}$ is set to 81, with PN variances $\sigma_{\Delta}^{2}=10^{-3}{\rm rad^{2}}$ and $\sigma_{\Delta}^{2}=10^{-4}{\rm rad^{2}}$. The results demonstrate that decreasing the PN covariance leads to significant improvements in combined estimation performance.
VII Conclusion
This paper investigates the optimization of RIS phase shift and the joint estimation of CFO, PN, and UE position under imperfect channel conditions for an OFDM system in a near-field static localization scenario. An iterative joint estimation algorithm based on the MAP criterion and GD algorithm is proposed. Extensive simulation results have demonstrated that by optimizing the RIS phase shift, the theoretical positioning accuracy has been improved by two orders of magnitude. Moreover, the maximum positioning accuracy can be lower than $\rm 10^{-2}m$. These findings strongly support the efficacy of the proposed estimation and optimization algorithms in the paper. In order to further improve the estimation performance and support a larger AOI, the optimization algorithm used in this paper requires averaging a large number of PEBs related to random positions and will result in extremely high computation complexity. Hence, it becomes imperative to utilize an optimization algorithm with reduced complexity in this scenario. This would be a fruitful area for further work.
Appendix A Derivation of partial derivatives of $\bm{\mu}[n]$ and $\bm{\mu}[n]^{*}$
The partial derivatives of $\bm{\mu}[n]$ with respect to all parameters to be estimated are derived as follows
$\frac{\partial\bm{\mu}[n]}{\partial\phi}$ $=\sqrt{P}j\frac{2\pi n}{N}e^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}\bm{\bar{w}}^{\rm H}\mathbf{G}[\mathbf{S}]_{:,n},$ (53a)
$\frac{\partial\bm{\mu}[n]}{\partial\bm{\theta}[n]}$ $=\sqrt{P}je^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}\bm{\bar{w}}^{\rm H}\mathbf{G}[\mathbf{S}]_{:,n},$ (53b)
$\frac{\partial\bm{\mu}[n]}{\partial\bm{\theta}[m]}$ $=0,m\neq n,$ (53c)
$\frac{\partial\bm{\mu}[n]}{\partial d_{Ru}}$ $=\sqrt{P}e^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}\bm{\bar{w}}^{\rm H}\frac{\partial\mathbf{G}}{\partial d_{Ru}}[\mathbf{S}]_{:,n},$ (53d)
$\frac{\partial\bm{\mu}[n]}{\partial\phi_{az}}$ $=\sqrt{P}e^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}\bm{\bar{w}}^{\rm H}\frac{\partial\mathbf{G}}{\partial\phi_{az}}[\mathbf{S}]_{:,n},$ (53e)
$\frac{\partial\bm{\mu}[n]}{\partial\phi_{el}}$ $=\sqrt{P}e^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}\bm{\bar{w}}^{\rm H}\frac{\partial\mathbf{G}}{\partial\phi_{el}}[\mathbf{S}]_{:,n},$ (53f)
and the partial derivatives of $\bm{\mu}[n]^{*}$ with respect to all parameters to be estimated are derived as follows
$\frac{\partial\bm{\mu}[n]^{*}}{\partial\phi}$ $=-j\sqrt{P}\frac{2\pi n}{N}e^{-j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}]_{:,n}^{\rm H}\mathbf{G}^{\rm H}\bm{\bar{w}},$ (54a)
$\frac{\partial\bm{\mu}[n]^{*}}{\partial\bm{\theta}[n]}$ $=-j\sqrt{P}e^{-j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}]_{:,n}^{\rm H}\mathbf{G}^{\rm H}\bm{\bar{w}},$ (54b)
$\frac{\partial\bm{\mu}[n]^{*}}{\partial\bm{\theta}[m]}$ $=0,m\neq n,$ (54c)
$\frac{\partial\bm{\mu}[n]^{*}}{\partial d_{Ru}}$ $=\sqrt{P}e^{-j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}]_{:,n}^{\rm H}\left(\frac{\partial\mathbf{G}}{\partial d_{Ru}}\right)^{\rm H}\bm{\bar{w}},$ (54d)
$\frac{\partial\bm{\mu}[n]^{*}}{\partial\phi_{az}}$ $=\sqrt{P}e^{-j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}]_{:,n}^{\rm H}\left(\frac{\partial\mathbf{G}}{\partial\phi_{az}}\right)^{\rm H}\bm{\bar{w}},$ (54e)
$\frac{\partial\bm{\mu}[n]^{*}}{\partial\phi_{el}}$ $=\sqrt{P}e^{-j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}]_{:,n}^{\rm H}\left(\frac{\partial\mathbf{G}}{\partial\phi_{el}}\right)^{\rm H}\bm{\bar{w}}.$ (54f)

### v3

@@
-Near-Field Localization and Phase Shift Optimization for RIS-Assisted Non-Ideal OFDM Systems
+Reconfigurable Intelligent Surface-Assisted Localization in OFDM Systems with Carrier Frequency Offset and Phase Noise
 Abstract
-By incorporating reconfigurable intelligent surface (RIS) into communication-assisted localization systems, the issue of signal blockage caused by obstacles can be addressed, and passive beamforming can be employed to enhance localization accuracy.
-However, existing works mainly consider ideal channels and do not account for the effects of realistic impairments like carrier frequency offset (CFO) and phase noise (PN) on localization.
-This paper proposes an iterative joint estimation algorithm for CFO, PN, and user position based on maximum a posteriori (MAP) criterion and gradient descent (GD) algorithm.
-Closed-form expressions for CFO and PN updates are provided.
-The hybrid Cramér-Rao lower bound (HCRLB) for the estimation parameters is derived, and the ambiguity in CFO and PN estimation is analyzed.
-To minimize the HCRLB, a non-convex RIS shift optimization problem is formulated and is transformed into a convex semidefinite programming (SDP) problem using the technique of semidefinite relaxation (SDR) and Schur complement.
-After optimizing the RIS phase shift, the theoretical positioning accuracy within the area of interest (AOI) can be improved by two orders of magnitude, with a maximum positioning root mean square error (RMSE) lower than $\rm 10^{-2}m$.
+Reconfigurable intelligent surface (RIS)-assisted communication systems have been extensively studied for providing high-precision location services.
+However, most studies have overlooked the impact of carrier frequency offset (CFO) and phase noise (PN) resulting from hardware impairments on localization.
+This paper presents a novel, alternating optimization (AO)-based algorithm to jointly estimate the CFO, PN, and user position, where, provided the user position, closed-form expressions for the CFO and PN are derived per iteration, significantly reducing the complexity and enhancing the stability of the algorithm.
+Another important aspect is a new RIS phase shift optimization algorithm developed to minimize the analytical lower bound of localization accuracy, hence benefiting localization.
+The semidefinite relaxation method and Schur complement are utilized to convexify this challenging non-convex optimization problem to a semidefinite program.
+Simulations demonstrate the effectiveness of the proposed algorithms, with the localization accuracy enhanced by two orders of magnitude.
+The localization accuracy of the proposed algorithm is close to the analytical lower bound, with a root mean square error of lower than $\rm 10^{-2}\>m$.
 Index Terms:
-Near-field localization, reconfigurable intelligent surface, passive beamforming, carrier frequency offset, phase noise.
+Localization, reconfigurable intelligent surface, phase shift configuration, carrier frequency offset, phase noise.
 I Introduction
-With the widespread application of fifth-generation (5G) networks worldwide, research focus has shifted towards future six-th generation (6G) mobile communication technologies.
-For 5G mobile networks, the demand for positioning accuracy has already reached sub-meter levels due to the need to support various location-based services.
-In 6G networks, it is expected that high-precision positioning at the centimeter level will play an even more important role [1, 2, 3].
-While global navigation satellite systems (GNSSs) have been widely used in various positioning scenarios [25], they suffer from issues such as insufficient positioning accuracy and signal instability due to obstacles [26, 27].
-In recent years, research has been conducted on using communication signals to assist in positioning, aiming to achieve more accurate and continuous positioning alongside GNSS signals.
+I-A Motivation and Challenges
+With the development of industrial Internet of Things, there is an increasing demand for centimeter-level localization accuracy in the sixth-generation (6G) mobile networks [1, 2, 3].
+While global navigation satellite systems (GNSSs) have been widely applied to outdoor scenarios [4], they undergo insufficient localization accuracy and signal instability arising from obstacles [5, 6].
+In recent years, communication systems have been considered to assist in localization, e.g., integrated sensing and communication (ISAC), aiming to achieve more accurate and continuous localization alongside GNSS signals.
+Although communication-assisted localization systems have been extensively studied, few studies have considered the adverse effects of hardware impairments, such as carrier frequency offset (CFO) and phase noise (PN), on localization accuracy.
+CFO is a constant parameter due to the frequency offset between the transmitter and the receiver.
+Doppler shift can also lead to CFO.
+PN is a random process resulting from the fluctuation of the oscillators at the transmitters and receivers [7].
+CFO and PN can result in signal constellation rotations in single carrier communication systems, or inter-carrier interference (ICI) in broadband communication systems, for example, orthogonal frequency division multiplexing (OFDM) systems.
+The overall communication performance and the localization accuracy can degrade significantly, due to the CFO and PN.
+Another challenge arises from the increasingly likely blockages of signals, e.g., millimeter wave (mmWave) signals with the quasi-optical property, in many dense urban areas [8].
+Reconfigurable intelligent surface (RIS) [9, 10] has been considered in communication systems to deal with the problem of signal blockage.
+Localization can be implemented with the help of virtual line-of-sight (LoS) paths provided by RISs even when the LoS paths are blocked, and can be improved via passive beamforming.
+In this sense, RIS-assisted user localization is of practical interest, and could be particularly challenging with the further consideration of the CFO and PN.
+However, this has not been studied in the existing literature.
+I-B Related works
+Communication systems have been widely employed to assist in high-precision user localization, typically under a far-field setting; i.e., the user and the anchors are in the far field of each other.
 For example, Shahmansoori et al.
-[10] proposed a three-stage positioning algorithm that improves the simultaneous orthogonal matching pursuit (SOMP) algorithm and the space alternating generalized expectation-maximization (SAGE) algorithm, and exploited the sparsity of millimeter-wave multiple-input multiple-output (MIMO) channels to lower the complexity of the channel model.
+[11] proposed a three-stage localization algorithm that improves the simultaneous orthogonal matching pursuit algorithm and the space alternating generalized expectation-maximization algorithm, and exploited the sparsity of millimeter-wave multiple-input multiple-output (MIMO) channels to simplify the channel model.
 Hu et al.
-[11] presented a two-dimensional (2D) positioning algorithm based on the estimating signal parameters via rotational invariance technique (ESPRIT), which estimates user positions by constructing signal subspaces.
-The estimator has closed-form expressions, which reduces the computational complexity significantly compared to traditional estimation methods.
-However, in these methods that utilize communication signals for localization, the problem of signal blockage by obstacles still remains unavoidable.
-In recent years, reconfigurable intelligent surface (RIS) technology has emerged as a key enabling technology for 6G networks [4, 5, 23].
-This technology allows for reconfiguring wireless channel environments, enabling communication even when the line-of-sight (LoS) link is obstructed.
-Introducing the RIS to assist in communication-based positioning has become a research focus [32, 31].
-From the perspective of mobility, the research can be divided into two categories: static positioning [17, 18, 19] and dynamic positioning [15, 13].
-The former considers scenarios where the user undergoes minimal movement within a coherent time, and the latter involves scenarios with the opposite characteristic.
-From the perspective of positioning distance, the research can also be divided into two categories: far-field positioning [19, 13, 20] and near-field positioning [18, 19, 13, 24].
-In the far-field case, the distance between the user and the RIS exceeds Fraunhofer distance [8], and the signal can be approximated as a planar wave.
-Conversely, in the near-field case, the signal cannot be treated as parallel propagation and should be considered a spherical wave.
+[12] presented a two-dimensional (2D) localization algorithm based on the celebrated estimating signal parameters via rotational invariance technique (ESPRIT), which estimates user positions by constructing signal subspaces.
+Some studies [13, 14] have explored the impact of realistic hardware impairments, including the CFO and PN, with no RIS considered.
+For instance, Meles et al.
+[13] simulated the effects of CFO and PN on angle-of-arrival (AOA) estimation.
+Chen et al.
+[14] considered the impacts of CFO, PN, mutual coupling, and power amplifier nonlinearity on 6G localization, and performed analysis using the derived misspecified Cramér-Rao lower bound (CRLB).
+However, neither of them provided an estimation algorithm to handle the hardware impairments.
+Some other works [15, 16, 17, 18] have delved into localization algorithm design for scenarios with either CFO or PN.
+On the one hand, the CFO was studied in [15, 16].
+Qiu et al.
+[15] designed an maximum likelihood estimation (MLE)-based weighted iterative algorithm for the joint estimation of CFO and AOA, yet the coupling of CFO and position parameters in each iteration incurs high complexity.
+Zakariyya et al.
+[16] proposed a high-resolution multiple signal classification (MUSIC) algorithm for localization, but did not utilize the CFO estimates in localization, leading to a loss in localization accuracy.
+On the other hand, the PN was considered in [17, 18].
+Xiao et al.
+[17] combined the expectation maximization (EM) and MUSIC algorithms for joint estimation of AOA and PN, but only considered a single-carrier system under a 2D setting.
+Zheng et al.
+[18] formulated a maximum a-posteriori (MAP) problem iteratively solved using the block majorization minimization (MM) algorithm after approximation, where the PN was simplified as a Gaussian variable.
+However, the above methods have not considered the use of RISs nor addressed the signal blockage by obstacles.
+RIS has been integrated into localization systems to address the problem of signal blockage, but little consideration has been given to the above-mentioned hardware impairments, i.e., CFO and PN.
+From the perspective of user mobility, existing studies can be divided between static localization [19, 20, 21] and dynamic localization [22, 23].
+The former encompasses scenarios where the users undergo minimal movement within a coherent time.
+The latter involves scenarios when the users move at a relatively high speed, and the Doppler shift cannot be overlooked.
+From the perspective of localization distance, existing research can be divided between far-field localization [21, 23, 24] and near-field localization [20, 21, 23, 25].
+In the far-field case, the user-RIS distance exceeds Fraunhofer distance [26], and the signal can be approximated as a planar wave.
+Conversely, in the near-field case, the signal cannot be treated as parallel propagation and should be considered as a spherical wave.
 For example, Luan et al.
-[12] focused on the near-field static 2D positioning scenario.
-The authors proposed a robust optimization method to optimize the RIS phase shift for improving localization accuracy.
-The paper also investigated the power optimization problem of the localization system.
+[27] focused on the near-field static 2D localization scenario, and proposed a robust optimization algorithm to configure the RIS phase shifts for improving localization accuracy.
+The transmit powers were also optimized in the considered localization system.
+Ozturk et al.
+[28] further introduced a more practical RIS model with phase-dependent amplitude variations in the near-field circumstance.
+The misspecified CRLB was derived and an approximated mismatched maximum likelihood-based algorithm was proposed for joint UE localization and RIS parameter estimation.
+The system model in this paper also belongs to the near-field static category.
+Although we adopt the classical ideal RIS phase shift model with unit amplitude for simplicity, we first take the CFO and PN into consideration.
 Elzanaty et al.
 [23] studied RIS-assisted localization for multi-antenna mobile users.
-They proposed a three-dimensional (3D) model that can be applied in both far-field and near-field scenarios and analyzed the position and orientation error bounds.
+They proposed a three-dimensional (3D) model that applies to both far-field and near-field scenarios, and analyzed the position and orientation error bounds.
 Lin et al.
-[14] introduced a novel twin-RIS configuration comprising two RISs with relative spatial rotation.
+[29] introduced a novel twin-RIS configuration comprising two RISs with relative spatial rotations.
 To extract the parameters of the cascaded channel, the authors employed array signal processing and atomic norm denoising techniques.
-They established nonlinear equations based on the characteristics of the twin-RIS structure and accurately recovered the multipath parameters.
-User localization for static users in the far-field scenario is then performed based on the estimated channel parameters.
+They established nonlinear equations exploiting the characteristics of the twin-RIS structure to recover the multipath parameters.
+User localization for static users in the far-field scenario was then performed based on the estimated channel parameters.
 Teng et al.
-[15] introduced a new message-passing algorithm called BULT to effectively solve the localization problem for mobile users in the far-field scenario.
-They also derived the Bayesian Cramér-Rao lower bound for the proposed algorithm.
-Although there has been extensive research on user localization with RIS in various scenarios such as static, dynamic, far-field, and near-field, most of the existing works assumed ideal channel conditions without considering realistic impairments like carrier frequency offset (CFO) and phase noise (PN) that can introduce uncertainties and affect system performance.
-Only a few studies have considered modeling certain interferences.
-For example, [13, 20, 16] took into account clock offset, and [16] considered phase offset.
-However, to the best of our knowledge, there is currently no research that specifically analyzed CFO and PN in RIS-assisted orthogonal frequency division multiplexing (OFDM) localization systems.
-In an OFDM system, CFO and PN can convolve with the data symbols, causing signal constellation rotation and introducing inter-carrier interference (ICI).
-This interference significantly degrades the overall system performance.
-Moreover, most aforementioned works have assumed the subcarriers’ signals are independent, which cannot hold in the presence of CFO and PN due to the absence of ICI.
-Therefore, localization algorithms need to leverage the received time-domain signals to account for these effects and accurately estimate the user’s position in this case.
-In this paper, we consider the problem of joint estimation of CFO, PN, and user position in a RIS-assisted OFDM system. The contributions of this paper are outlined as follows.
-  • We investigate the problem of static near-field localization in an OFDM system assisted by RIS, considering the effects of CFO and PN.
-Moreover, we establish a time-domain model and introduce the frequency-domain channel with time-delay parameters into the model through time-frequency transformation to meet the requirements of position estimation.
-  • We propose an iterative estimation approach based on the maximum a posteriori (MAP) criterion and gradient descent (GD) algorithm to jointly estimate the CFO, the PN, and the user position.
-The algorithm significantly reduces complexity by deriving closed-form expressions for CFO and PN estimation.
-  • We derive the hybrid Cramér-Rao lower bound (HCRLB) for the joint estimation problem and analyze the phase ambiguity issue in CFO and PN estimation.
-A new metric is introduced to evaluate the accuracy of joint CFO and PN estimation.
-  • We implement phase shift optimization for RIS to enhance the theoretical positioning accuracy.
-By applying the semidefinite relaxation (SDR) method and utilizing the Schur complement, we convert the initially non-convex optimization problem into a convex semidefinite programming (SDP) problem.
-It is shown that the theoretical positioning accuracy has been improved by two orders of magnitude.
-Figure 1: RIS-aided 3D localization scenario, where the LoS link is blocked.
-The paper is structured as follows.
-Section II introduces the system model, while Section III presents the joint estimation algorithm.
-Section IV is dedicated to the derivation of the HCRLB for the proposed joint estimation problem.
-The proposed RIS phase shift optimization method is discussed in Section V.
-Section VI showcases comprehensive simulation results, and the paper concludes in Section VII.
-Notations: $\mathbb{E}(\cdot)$, $p(\cdot)$ denote the expectation and the probability density function (PDF) of the argument, respectively.
-${\rm tr}(\mathbf{M})$, $\det\mathbf{M}$, $\mathbf{M}^{\top}$, $\mathbf{M}^{*}$, $\mathbf{M}^{\rm H}$, $\mathbf{M}^{\rm-1}$ denote the trace, determinant, transpose, conjugate, hermitian, and inverse of matrix $\mathbf{M}$, respectively.
-${\rm Diag}(\bm{v})$ indicates a diagonal matrix created by assigning the elements of vector $\bm{v}$ along its main diagonal.
+[22] introduced a new message-passing algorithm called BULT to locate mobile users in the far-field scenario.
+The Bayesian CRLB was derived for BULT.
+The existing RIS-assisted localization studies have assumed ideal channel conditions without considering realistic hardware impairments.
+Only several studies have considered modeling some specific impairments.
+For example, the studies in [23, 24, 30] took into account clock offset, and the work in [30] considered phase offsets.
+Zhang et al.
+[31] considered the CFO while overlooking the effects of PN.
+They proposed a MUSIC-based algorithm that directly estimates the position parameters together with the CFO.
+However, the algorithm is only suitable for far-field scenarios.
+To the best of our knowledge, no existing research has specifically analyzed the impact of both CFO and PN on RIS-assisted near-field localization systems.
+I-C Contributions
+In this paper, we investigate a RIS-assisted OFDM system, where practical CFO and PN convolve with the data symbols, causing signal constellation rotation and ICI and, consequently, hindering user localization in near-field environments.
+The key contributions of this paper are outlined as follows:
+  • We investigate a problem of static near-field localization in a RIS-aided OFDM system, while taking into account the CFO and PN.
+We propose an alternating optimization (AO)-based iterative estimation approach, where the MAP criterion and gradient descent (GD) algorithm are employed with closed-form expressions derived for the CFO and PN estimation.
+  • We derive the hybrid CRLB (HCRLB) for the considered joint estimation problem.
+The phase ambiguity in CFO and PN estimation is analyzed.
+To evaluate the accuracy of joint CFO and PN estimation, a joint mean square error (MSE) metric is introduced.
+  • We further optimize the phase shifts of the RIS to improve the performance of localization.
+By applying the semi-definite relaxation (SDR) and the Schur complement, we convert the non-convex optimization problem into a convex semi-definite programming (SDP) problem solved efficiently.
+  • We perform comprehensive numerical simulations to assess the performance of the proposed algorithms.
+It is shown that the localization accuracy is improved by two orders of magnitude compared to random RIS phase shift configuration.
+Compared to the state-of-the-art benchmarks, our algorithms can improve the localization accuracy by about one order of magnitude under relatively low transmit power and even by two orders of magnitude under high transmit power.
+Figure 1: The considered RIS-aided 3D near-field localization system, including an AN, a UE, and an $N_{R}$-element RIS.
+The LoS link between the AN and UE is obstructed.
+The UE utilizes the signals transmitted by the AN and reflected by the RIS to localize its position.
+The rest of this paper is organized as follows.
+Section II presents the system model.
+Section III presents the joint estimation algorithm.
+Section IV derives the HCRLB for the joint estimation algorithm.
+The proposed RIS phase shift optimization method is delineated in Section V.
+Section VI showcases comprehensive simulation results, and the paper is concluded in Section VII.
+Notation: $\mathbb{E}(\cdot)$ and $p(\cdot)$ denote expectation and probability density function (PDF), respectively.
+${\rm tr}(\cdot)$, $\det(\cdot)$, $(\cdot)^{\top}$, $(\cdot)^{*}$, $(\cdot)^{\rm H}$, and $(\cdot)^{\rm-1}$ denote trace, determinant, transpose, conjugate, hermitian, and inverse, respectively.
+${\rm Diag}(\bm{v})$ indicates a diagonal matrix where the entries of vector $\bm{v}$ appear on its main diagonal.
 $\bm{v}[n]$ denotes the $n$-th entry of vector $\bm{v}$.
-Unless otherwise specified, all indices in the article start from 0.
-${\rm Blkdiag}(\mathbf{M}_{0},\mathbf{M}_{1},\cdots,\mathbf{M}_{N-1})$ denotes a block diagonal matrix with $\mathbf{M}_{0},\mathbf{M}_{1},\cdots,\mathbf{M}_{N-1}$ being its diagonal matrices.
-$\left[\mathbf{M}\right]_{i:j,:}$ refers to a submatrix encompassing rows $i$ to $j$ and all columns of matrix $\mathbf{M}$, while $\left[\mathbf{M}\right]_{:,i:j}$ denotes a submatrix encompassing columns $i$ to $j$ and all rows of matrix $\mathbf{M}$.
-$\bm{0}_{M\times N}$, and $\mathbf{I}_{M}$ denote $M\times N$ zeros matrix, and $M\times M$ identity matrix, respectively.
-${}\Re\left\{\cdot\right\}$ and ${}\Im\left\{\cdot\right\}$ represent the real and imaginary components of the complex value, respectively.
-$\mathbb{C}^{m\times n}$ and $\mathbb{R}^{m\times n}$ denote $m\times n$ matrices with complex and real entries, respectively.
-Finally, $\mathcal{CN}(\bm{\mu},\mathbf{\Sigma})$ and $\mathcal{N}(\bm{\mu},\mathbf{\Sigma})$ represent complex and real Gaussian distributions, respectively, characterized by a mean vector $\bm{\mu}$ and covariance matrix $\mathbf{\Sigma}$.
+${\rm Blkdiag}(\mathbf{M}_{0},\mathbf{M}_{1},\cdots,\mathbf{M}_{N-1})$ denotes a block diagonal matrix with $\mathbf{M}_{0},\mathbf{M}_{1},\cdots,\mathbf{M}_{N-1}$ being the matrices along its main diagonal.
+$\left[\mathbf{M}\right]_{i:j,:}$ refers to a sub-matrix encompassing rows $i$ to $j$ of matrix $\mathbf{M}$, while $\left[\mathbf{M}\right]_{:,i:j}$ denotes a sub-matrix encompassing columns $i$ to $j$ of matrix $\mathbf{M}$.
+$\bm{0}_{M\times N}$ and $\mathbf{I}_{M}$ denote the $M\times N$ all-zero matrix and the $M\times M$ identity matrix, respectively.
+${}\Re\left\{\cdot\right\}$ and ${}\Im\left\{\cdot\right\}$ represent the real and imaginary components of a complex value, respectively.
+$\mathbb{C}^{m\times n}$ and $\mathbb{R}^{m\times n}$ collect all $m\times n$ matrices with complex and real entries, respectively.
+$\mathcal{CN}(\bm{\mu},\mathbf{\Sigma})$ and $\mathcal{N}(\bm{\mu},\mathbf{\Sigma})$ represent complex and real Gaussian distributions, respectively, with a mean vector $\bm{\mu}$ and covariance matrix $\mathbf{\Sigma}$.
+Unless otherwise specified, all indices in the paper start from 0.
+Moreover, we summarize the notation used in this paper in Table I.
+TABLE I: Notation and definition.
+Notation Definition
+$N_{R}$ Number of RIS elements
+$N$ Number of subcarriers
+$L$ Length of the approximate PN vector
+$d$ Element spacing of the RIS
+$D$ Aperture of the RIS
+$f_{c}$ Carrier frequency of the system
+$B$ Bandwidth of the system
+$\lambda$ Wavelength of the system
+$c$ Speed of light
+$P$ Transmit power at the AN
+$\Delta f$ CFO
+$\phi$ Normalized CFO
+$U$ Position sampling amount
+$\sigma^{2}$ Variance of the noise at the UE
+$\sigma_{\Delta}^{2}$ Innovation variance for PN
+$\bm{p}_{a}$ Coordinate of the AN
+$\bm{p}_{r}$ Coordinate of the reference point of the RIS
+$\bm{p}_{u}$ Coordinate of the UE position
+$\bm{\theta}$ PN vector
+$\bm{\eta}$ Approximate PN vector
+$\bm{y}$ Received signal at the UE
+$\bm{\mu}$ Noise-free received signal at UE
+$\bm{h}$ AN-RIS-UE frequency-domain channel
+$\bm{w}$ RIS phase shift vector
+$\mathbf{s}$ Transmitted symbol sequence at the AN
+$\bm{v}$ Additive noise at UE
+$\bm{\lambda}$ Parameter vector to be estimated
+$\bm{\alpha}$ Transformed parameter vector
+$\mathbf{B}$ Bayesian information matrix
+$\bm{\Lambda}_{\phi}$ CFO matrix
+$\bm{\Lambda}_{\bm{\theta}}$ PN matrix
+$\mathbf{\Pi}$ Transition matrix for the PN vector
+$\mathbf{\Xi}$ Transition matrix for the HCRLB
 II System Model
-Fig.
-1 illustrates the RIS-assisted OFDM downlink wireless communication system under consideration, comprising a single anchor node (AN) and a single user equipment (UE).
-In this scenario, we make the assumption that there are obstacles blocking the LoS path between AN and UE, and the channel state information (CSI) remains unchanged during the localization process, i.e., within the duration of one OFDM symbol, similar to [19, 12, 16].
-The positions of the AN and each element of the RIS are assumed to be known, which can be denoted as $\bm{p}_{a}$, and $\bm{p}_{r}=[x_{r},y_{r},z_{r}]^{\top}$, respectively, where $r$ represents the index of the RIS element among a total of $N_{R}$ elements.
-The position of UE, which is the primary parameter to be estimated, is denoted as $\bm{p}_{u}=[x_{u},y_{u},z_{u}]^{\top}$.
-The RIS elements are organized in the form of a uniform planar array (UPA).
+The considered RIS-assisted OFDM downlink wireless localization system is composed of an anchor node (AN), a user equipment (UE), and an $N_{R}$-element RIS, as illustrated in Fig.
+1.
+We make the assumption that the LoS path between the AN and UE is obstructed, and the channel state information (CSI) remains unchanged during localization, as in [21, 27, 30].
+The positions of the AN and each RIS element, denoted as $\bm{p}_{a}$ and $\bm{p}_{r}=[x_{r},y_{r},z_{r}]^{\top}$, respectively, are known, where $r$ is the index of a RIS element.
+The position of UE, denoted as $\bm{p}_{u}=[x_{u},y_{u},z_{u}]^{\top}$, is to be estimated.
+The RIS elements are organized in a uniform planar array (UPA).
 We take the bottom-left corner of the RIS as the reference point and place it at the coordinate origin.
-The RIS is positioned on the $yz$ plane, and both edges on the coordinate axes are located in the positive half-axis.
-The coordinates of the $r$-th RIS element can be represented as $\bm{p}_{r}=[0,d\times((r-1)\mod N_{R}),d\times\lceil\frac{r}{\sqrt{N_{R}}}-1\rceil]^{\top}$, where $d$ stands for the element spacing of RIS, whose value is typically set as half of the wavelength.
-To satisfy the conditions of the near-field setup, both the distance between AN and RIS, as well as the distance between RIS and UE, should lie within the Fresnel region [8].
-Specifically, the distances should lie within the interval $[0.62\sqrt{D^{3}/\lambda},2D^{2}/\lambda]$, where $D=\sqrt{2}d(\sqrt{N_{R}}-1)$ represents the aperture of RIS and $\lambda$ is the operating wavelength.
-Additionally, we adopt a polar coordinate system for estimating the user’s position, i.e., using $d_{Ru}$, $\phi_{az}$, $\phi_{el}$ to express the user’s position, which represents the distance, azimuth angle, and elevation angle of UE relative to the coordinate origin, respectively.
-The transformation relationships between these three parameters and $x_{u}$, $y_{u}$, $z_{u}$ are as follows
-$d_{Ru}$ $=\sqrt{x_{u}^{2}+y_{u}^{2}+z_{u}^{2}},$ (1a)
-$\phi_{az}$ $=\arctan\frac{y_{u}}{x_{u}},$ (1b)
+The RIS is placed on the $yz$-plane, and both edges on the coordinate axes are located in the positive half-axis.
+Hence, the coordinates of the $r$-th RIS element can be represented as $\bm{p}_{r}=[0,d\times((r-1)\mod N_{R}),d\times\lceil\frac{r}{\sqrt{N_{R}}}-1\rceil]^{\top}$, where $d$ stands for the element spacing of the RIS, and is typically half of a wavelength.
+To satisfy the conditions of the near-field setup, both the AN-RIS distance, as well as the RIS-UE distance, should lie within the Fresnel region [26], i.e., the interval $[0.62\sqrt{D^{3}/\lambda},2D^{2}/\lambda]$, where $D=\sqrt{2}d(\sqrt{N_{R}}-1)$ represents the aperture of the RIS and $\lambda$ denotes the wavelength.
+Additionally, we adopt a polar coordinate system for estimating the user’s position, i.e., using $d_{Ru}$, $\phi_{az}$ and $\phi_{el}$, to express the user’s position, indicating the distance, azimuth angle, and elevation angle of UE relative to the coordinate origin, respectively.
+The relationships between these three parameters and $x_{u}$, $y_{u}$ and $z_{u}$ are given by
+$d_{Ru}$ $=\sqrt{x_{u}^{2}+y_{u}^{2}+z_{u}^{2}};$ (1a)
+$\phi_{az}$ $=\arctan\frac{y_{u}}{x_{u}};$ (1b)
 $\phi_{el}$ $=\arccos\frac{z_{u}}{\sqrt{x_{u}^{2}+y_{u}^{2}+z_{u}^{2}}},$ (1c)
-and in turn, can be expressed as
-$x_{u}$ $=d_{Ru}\sin\phi_{el}\cos\phi_{az},$ (2a)
-$y_{u}$ $=d_{Ru}\sin\phi_{el}\sin\phi_{az},$ (2b)
+and, in turn, $x_{u}$, $y_{u}$ and $z_{u}$ can be written as
+$x_{u}$ $=d_{Ru}\sin\phi_{el}\cos\phi_{az};$ (2a)
+$y_{u}$ $=d_{Ru}\sin\phi_{el}\sin\phi_{az};$ (2b)
 $z_{u}$ $=d_{Ru}\cos\phi_{el}.$ (2c)
-The localization process in the paper is performed using one OFDM symbol, with $N$ subcarriers, a center carrier frequency of $f_{c}$, corresponding to a wavelength of $\lambda$, and a signal bandwidth of $B$.
-Upon sampling at a rate of $T_{s}=\frac{1}{NB}$ and removing the cyclic prefix (CP) at UE, the received time-domain signal can be represented as
-$\bm{y}=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\tilde{\bm{h}}+\tilde{\bm{v}},$ (3)
+An OFDM symbol with a center carrier frequency of $f_{c}$, a signal bandwidth of $B$ and $N$ subcarriers is considered in this paper.
+Upon sampling at a rate of $T_{s}=\frac{1}{NB}$ and removing the cyclic prefix (CP) at the UE, the received time-domain signal $\bm{y}$ can be written as
+$\bm{y}=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\bm{h}_{LoS}+\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\bm{h}_{NLoS}+\tilde{\bm{v}},$ (3)
 where
-  • ${\bm{y}\triangleq\left[\bm{y}[0],\bm{y}[1],\cdots,\bm{y}[N-1]\right]}^{\top}$,
-  • $P$ is the transmit power,
-  • $\bm{\Lambda}_{\phi}$ is the impact caused by CFO on the signal.
-It is a diagonal matrix, and each item on the diagonal can be expressed as $\left[\bm{\Lambda}_{\phi}\right]_{n,n}=e^{j2\pi\frac{n\Delta f}{NB}}=e^{j2\pi\frac{n}{N}\phi}$, where $\phi=\frac{\Delta f}{B}$ is called the normalized CFO,
-  • $\bm{\Lambda}_{\bm{\theta}}\triangleq{\rm Diag}\left[e^{j\bm{\theta}[0]},e^{j\bm{\theta}[1]},\cdots,e^{j\bm{\theta}[N-1]}\right]$ is the impact caused by PN on the signal, further details about which will be discussed in the later part of this section,
-  • $\mathbf{S}\triangleq\sqrt{N}\mathbf{F}^{\rm H}{\rm Diag}(\bm{s})$, where $\mathbf{F}^{\rm H}$ is the inverse fast Fourier transform (IFFT) matrix and $\bm{s}$ is the frequency domain transmitted symbol sequence at the AN,
-  • $\tilde{\bm{h}}\triangleq\bm{h}_{LoS}+\bm{h}_{NLoS}$ is the frequency domain channel, with $\bm{h}_{LoS}$ and $\bm{h}_{NLoS}$ representing the LoS and the non-line-of-sight (NLoS) part of the channel, respectively.
-We assume $\bm{h}_{NLoS}$ follows a known distribution $\bm{h}_{NLoS}\sim\mathcal{CN}(0,\sigma_{NLoS}^{2}\mathbf{I}_{N})$.
-Hence, The impact caused by $\bm{h}_{NLoS}$ on the signal can be combined with the additive Gaussian noise $\tilde{\bm{v}}$.
-Consequently, we can rewrite (3) as
+  • $P$ is the transmit power;
+  • $\bm{\Lambda}_{\phi}$ is the CFO matrix resulting from a frequency mismatch between the transmitter and receiver.
+$\bm{\Lambda}_{\phi}$ is a diagonal matrix with $\left[\bm{\Lambda}_{\phi}\right]_{n,n}=e^{j2\pi\frac{n\Delta f}{NB}}=e^{j2\pi\frac{n}{N}\phi}$ being the $(n,n)$-th element along its main diagonal, where $\Delta f$ is the CFO, and $\phi=\frac{\Delta f}{B}$ is called the normalized CFO;
+  • $\bm{\Lambda}_{\bm{\theta}}\triangleq{\rm Diag}\left[e^{j\bm{\theta}[0]},e^{j\bm{\theta}[1]},\cdots,e^{j\bm{\theta}[N-1]}\right]$ is the PN matrix.
+Drawing upon the attributes of the PN observed in practical oscillators, the PN vector $\bm{\theta}$ can be modeled as a Wiener process [32], i.e.,
+    $\bm{\theta}[n]=\bm{\theta}[n-1]+\bm{\Delta}[n],$ (4)
+    where $\bm{\Delta}[n]$ is a Gaussian variable, i.e., $\bm{\Delta}[n]\sim\mathcal{N}(0,\sigma_{\Delta}^{2})$.
+As in [33], we assume that $\bm{\theta}[-1]=0$.
+Then, the PN vector $\bm{\theta}\triangleq\left[\bm{\theta}[0],\bm{\theta}[1],\cdots,\bm{\theta}[N-1]\right]^{\top}$ yields $\bm{\theta}\sim\mathcal{N}(\bm{0}_{N\times 1},\bm{\Psi})$, where $\mathbf{\Psi}$ is given by
+    $\mathbf{\Psi}=\sigma_{\Delta}^{2}\begin{bmatrix}1&1&1&\cdots&1&1\\ 1&2&2&\cdots&2&2\\ 1&2&3&3&\cdots&3\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 1&2&3&\cdots&N-1&N\end{bmatrix};$ (5)
+  • $\mathbf{S}\triangleq\sqrt{N}\mathbf{F}^{\rm H}{\rm Diag}(\bm{s})$, where $\mathbf{F}^{\rm H}$ is the inverse fast Fourier transform (FFT) matrix and $\bm{s}$ is the frequency-domain transmitted symbol sequence at the AN;
+  • $\bm{h}_{LoS}$ and $\bm{h}_{NLoS}$ represent the LoS and non-LoS parts of the frequency-domain channel, respectively;
+  • $\tilde{\bm{v}}$ is the additive Gaussian noise at the UE, and follows the distribution $\tilde{\bm{v}}\sim\mathcal{CN}(\bm{0}_{N\times 1},\sigma_{AWGN}^{2}\mathbf{I}_{N})$.
+Assume that $\bm{h}_{NLoS}$ follows a known Gaussian distribution $\bm{h}_{NLoS}\sim\mathcal{CN}(0,\sigma_{NLoS}^{2}\mathbf{I}_{N})$.
+Thus, $\bm{\mu}_{NLoS}\triangleq\mathbf{X}\bm{h}_{NLoS}$ also follows a known Gaussian distribution $\bm{\mu}_{NLoS}\sim\mathcal{CN}(0,\sigma_{NLoS}^{2}\mathbf{X}\mathbf{X}^{\rm H})$ with $\mathbf{X}\triangleq\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}$.
+Combine $\bm{\mu}_{NLoS}$ with the additive Gaussian noise $\tilde{\bm{v}}$.
+For conciseness, denote $\bm{h}_{LoS}$ with $\bm{h}$.
+The received signal is rewritten as
 $\bm{y}=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\bm{h}+\bm{v},$ (6)
-    where $\bm{h}$ and $\bm{v}$ represent the LoS part of the channel and the combined noise, respectively.
-$\bm{h}$ contains the known part of the channel from AN to RIS and the unknown part of the channel from RIS to UE.
-Each item of it can be expressed by $\bm{h}[n]=\sum_{r=1}^{N_{R}}\sqrt{\rho_{aR,n,r}}\sqrt{\rho_{Ru,n,r}}\bm{w}[r]e^{-j2\pi\frac{nB}{N}(\tau_{ar}+\tau_{ru})}$, where $\sqrt{\rho_{aR,n,r}}=\frac{\sqrt{G_{a}}\lambda_{n}}{4\pi d_{ar}}$ and $\sqrt{\rho_{Ru,n,r}}=\frac{\sqrt{G_{t}}\lambda_{n}}{4\pi d_{ru}}$ are the free-space path loss from AN to RIS, and from RIS to UE, respectively.
-$\bm{w}[r]$ is the phase shift of the $r$-th element of RIS, $\tau_{ar}=\frac{d_{ar}}{c}=\frac{\|\bm{p}_{r}-\bm{p}_{a}\|}{c}$ represents the delay from AN to the $r$-th element of RIS, while $\tau_{ru}=\frac{d_{ru}}{c}=\frac{\|\bm{p}_{u}-\bm{p}_{r}\|}{c}$ denotes the delay from the $r$-th element of RIS to UE, with $c$ representing the light speed,
-  • $\bm{v}\triangleq\left[\bm{v}[0],\bm{v}[1],\cdots,\bm{v}[N-1]\right]$ is the combined additive noise at UE, and follows the distribution $\bm{v}\sim\mathcal{CN}(\bm{0}_{N\times 1},\sigma^{2}\mathbf{I}_{N})$.
-Drawing upon the attributes of PN observed in practical oscillators, it can be effectively modeled as a Wiener process [22], i.e.,
-$\bm{\theta}[n]=\bm{\theta}[n-1]+\bm{\Delta}[n],$ (5)
-where $\bm{\Delta}[n]$ is a real Gaussian variable with a distribution of $\bm{\Delta}[n]\sim\mathcal{N}(0,\sigma_{\Delta}^{2})$.
-As in [21], it is assumed that $\bm{\theta}[-1]=0$, and then we can obtain that the PN vector $\bm{\theta}\triangleq\left[\bm{\theta}[0],\bm{\theta}[1],\cdots,\bm{\theta}[N-1]\right]^{\top}$ follows a distribution of $\bm{\theta}\sim\mathcal{N}(\bm{0}_{N\times 1},\bm{\Psi})$, where $\mathbf{\Psi}$ can be expressed by
-$\mathbf{\Psi}=\sigma_{\Delta}^{2}\begin{bmatrix}1&1&1&\cdots&1&1\\ 1&2&2&\cdots&2&2\\ 1&2&3&3&\cdots&3\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 1&2&3&\cdots&N-1&N\end{bmatrix}.$ (6)
-III Proposed Joint Position, CFO and PN Estimation
-In previous works [12, 16], localization was based on frequency-domain signals.
-However, in our problem, the introduction of CFO and PN would result in ICI if the signal is transformed into the frequency domain, which will increase the difficulty of our estimation.
-Therefore, in this paper, we conduct the estimation in the time domain.
-In this section, we propose an iterative algorithm based on MAP criterion and GD algorithm, which jointly estimates the normalized CFO, PN, and position of UE.
-By defining parameter vector $\bm{\lambda}\triangleq\left[\phi,\bm{\theta}^{\top},d_{Ru},\phi_{az},\phi_{el}\right]^{\top}$, whereas $\bm{\theta}$ is a random variable, and the others are deterministic variables, the posterior distribution of $\bm{\lambda}$ given the received signal $\bm{y}$ can be written as
+where $\bm{v}\triangleq\bm{\mu}_{NLoS}+\tilde{\bm{v}}$ is Gaussian distributed, i.e, $\bm{v}\sim\mathcal{CN}(0,\sigma^{2}\mathbf{I}_{N})$, with $\sigma^{2}=\sigma_{AWGN}^{2}+\sigma_{NLoS}^{2}$.
+Each element of $\bm{h}$ can be written as $\bm{h}[n]=\sum_{r=1}^{N_{R}}\sqrt{\rho_{aR,n,r}}\sqrt{\rho_{Ru,n,r}}\bm{w}[r]e^{-j2\pi\frac{nB}{N}(\tau_{ar}+\tau_{ru})}$, where $\sqrt{\rho_{aR,n,r}}=\frac{\sqrt{G_{a}}\lambda_{n}}{4\pi d_{ar}}$ and $\sqrt{\rho_{Ru,n,r}}=\frac{\sqrt{G_{t}}\lambda_{n}}{4\pi d_{ru}}$ are the free-space path losses from the AN to the $r$-th RIS element, and from the $r$-th RIS element to the UE, respectively.
+$\bm{w}[r]$ is the phase shift of the $r$-th RIS element, $\tau_{ar}=\frac{d_{ar}}{c}=\frac{\|\bm{p}_{r}-\bm{p}_{a}\|}{c}$ is the delay from the AN to the $r$-th RIS element, while $\tau_{ru}=\frac{d_{ru}}{c}=\frac{\|\bm{p}_{u}-\bm{p}_{r}\|}{c}$ denotes the delay from the $r$-th RIS element to the UE, with the light speed $c$.
+III Proposed Joint CFO, PN and UE Position Estimation Algorithm
+In previous works [27, 30], localization was based on frequency-domain signals, and different subcarriers’ signals were often assumed to be independent.
+However, the CFO and PN would result in ICI, which increases the difficulty in the estimation of the CFO, PN and the position of the UE.
+The estimation can only be conducted in the time domain.
+In this section, we propose an AO-based iterative algorithm, which employs the MAP criterion and GD algorithm to jointly estimate the normalized CFO, PN, and the UE position.
+Define the parameter vector $\bm{\lambda}\triangleq\left[\phi,\bm{\theta}^{\top},d_{Ru},\phi_{az},\phi_{el}\right]^{\top}$, wherein $\bm{\theta}$ is a random variable and the others are deterministic.
+The posterior distribution of $\bm{\lambda}$, given the received signal $\bm{y}$, can be expressed as
 $p(\bm{\lambda}\mid\bm{y})=\frac{p(\bm{y}\mid\bm{\lambda})\times p(\bm{\theta})}{p(\bm{y})}.$ (7)
-After removing the constant terms, the object function can be expressed as $\mathcal{L}(\bm{\lambda})=-\log p(\bm{y}\mid\bm{\lambda})-\log p(\bm{\theta})$, and our goal is to achieve the joint estimation of $\bm{\lambda}$ by optimizing the following unconstrained function.
-$\hat{\bm{\lambda}}\propto\arg\min_{\bm{\lambda}}\mathcal{L}(\bm{\lambda}).$ (8)
+According to the MAP criterion, $\bm{\lambda}$ can be estimated by optimizing the following unconstrained function:
+$\hat{\bm{\lambda}}\propto\arg\min_{\bm{\lambda}}\mathcal{L}(\bm{\lambda}),$ (8)
+where $\mathcal{L}(\bm{\lambda})=-\log p(\bm{y}\mid\bm{\lambda})-\log p(\bm{\theta})$, which is obtained by calculating the negative logarithmic function of (7) and then suppressing constants.
 III-A Phase Noise Estimation
 III-A1 Approximation of PN vector
-Figure 2: Eigenvalues of $\bm{\Psi}$ at $N=32$ and $\sigma_{\Delta}^{2}=10^{-3}$.
-From (4), it can be observed that there are multiple parameters that need to be jointly estimated.
-In order to alleviate the computational complexity at UE, we take advantage of the properties of the PN vector to approximate it.
-This approach results in a reduction of unknown parameters that require estimation.
-As depicted in Fig.
-2, the eigenvalues of $\bm{\Psi}$ mostly cluster around zero.
-This observation allows us to approximate the PN vector as follows
+Figure 2: Eigenvalues of $\bm{\Psi}$ at the number of subcarriers $N=32$, and the innovation variance for PN $\sigma_{\Delta}^{2}=10^{-3}$.
+From the definition of $\bm{\lambda}$, many variables need to be jointly estimated at the UE.
+To lower the computational complexity of the estimation, we exploit the redundancy of the covariance matrix of the PN vector to approximate it and reduce variables requiring estimation.
+As illustrated in Fig. 2, the eigenvalues of $\bm{\Psi}$ are mostly close to zero, indicating significant redundancy in $\bm{\Psi}$. With this property, we can approximate the PN vector as
 $\bm{\theta}\approx\bm{\Pi}\bm{\eta},$ (9)
-where $\bm{\eta}\sim\mathcal{N}(\bm{0}_{L\times 1},\mathbf{I}_{L})$ contains $L\leq N$ PN parameters, and $\bm{\Pi}\in\mathbb{R}^{N\times L}$ is the transformation matrix.
-To derive $\bm{\Pi}$, we first perform singular value decomposition (SVD) on $\bm{\Psi}$, which is given by $\bm{\Psi}=\mathbf{U}\mathbf{D}\mathbf{U}^{\top}$, where $\mathbf{D}={\rm Diag}(\bm{e})$, and $\bm{e}$ represents the eigenvalues of $\bm{\Psi}$ arranged in descending order.
+where $\bm{\eta}\sim\mathcal{N}(\bm{0}_{L\times 1},\mathbf{I}_{L})$ contains $L\leq N$ PN variables, and $\bm{\Pi}\in\mathbb{R}^{N\times L}$ is the transition matrix.
+We set $L=\frac{N}{2}$ in this paper.
+To derive $\bm{\Pi}$, we first perform eigenvalue decomposition (EVD) on $\bm{\Psi}$, i.e., $\bm{\Psi}=\mathbf{U}\mathbf{D}\mathbf{U}^{\top}$, where $\mathbf{D}={\rm Diag}(\bm{e})$ and $\bm{e}$ collects the eigenvalues of $\bm{\Psi}$ arranged in descending order.
 Subsequently, $\bm{\Pi}$ can be calculated by $\bm{\Pi}=[\mathbf{U}]_{:,0:L-1}\times\sqrt{\bm{e}}$.
-In the subsequent analysis, $\bm{\eta}$ is estimated instead of $\bm{\theta}$.
+In the following analysis, $\bm{\eta}$ is estimated, instead of $\bm{\theta}$.
 III-A2 Estimation of approximate PN vector
-After approximating the PN parameter $\bm{\theta}$, the objective function can be rewritten as
+By approximating the PN vector $\bm{\theta}$, the objective function on the right-hand side (RHS) of (8) can be rewritten as
 $\mathcal{L}(\bm{\lambda})$ $=-\log p(\bm{y}\mid\bm{\lambda})-\log p(\bm{\eta})$
 $\propto\frac{1}{\sigma^{2}}\|\bm{y}-\bm{\mu}\|^{2}+\frac{1}{2}\bm{\eta}^{\top}\bm{\eta},$ (10)
-where $\bm{\mu}\triangleq\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\bm{h}$ is the received signal at UE without additive noise, and the noise variance $\sigma^{2}$ is assumed to be known.
-Due to the typically small PN variance in practical oscillators, the elements of $\bm{\Lambda}_{\bm{\theta}}$ can be approximated by employing a first-order Taylor series expansion, resulting in $e^{j\bm{\theta}[n]}\approx 1+j\bm{\theta}[n]$.
+where $\bm{\mu}\triangleq\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\bm{h}$ stands for the noise-free received signal at the UE.
+Owing to the typically small PN variance of practical oscillators, the elements of the PN matrix $\bm{\Lambda}_{\bm{\theta}}$ can be approximated by employing its first-order Taylor expansion, resulting in $e^{j\bm{\theta}[n]}\approx 1+j\bm{\theta}[n]$.
 Consequently, we can approximate $\bm{\Lambda}_{\bm{\theta}}$ as $\bm{\Lambda}_{\bm{\theta}}\approx\mathbf{I}_{N}+j{\rm Diag}(\bm{\Pi\eta})$.
-With this approximation, the objective function in (III-A2) can be reformulated as:
+With this approximation, the objective function in (III-A2) can be rewritten as
 $\mathcal{L}_{\bm{\eta}}^{[q+1]}$ $\approx\frac{1}{\sigma^{2}}(\bar{\bm{y}}-\mathbf{Q}\bm{\eta})^{\rm H}(\bar{\bm{y}}-\mathbf{Q}\bm{\eta})+\frac{1}{2}\bm{\eta}^{\top}\bm{\eta}$
@@
 where $\bar{\bm{y}}\triangleq\bm{y}-\sqrt{P}\hat{\bm{\Lambda}}_{\phi}^{[q]}\mathbf{S}^{\top}\hat{\bm{h}}^{[q]}$, and $\mathbf{Q}\triangleq j\sqrt{P}{\rm Diag}(\hat{\bm{\Lambda}}_{\phi}^{[q]}\mathbf{S}^{\top}\hat{\bm{h}}^{[q]})\bm{\Pi}$.
-The superscript $q$ represents the round number of iterations.
-Subsequently, by setting the gradient of (III-A2) equal to zero, i.e.,
+The superscript ``$^{q}"$ indicates the $q$-th iteration.
+Set the partial derivative of (III-A2) regarding $\bm{\eta}$ to zero, i.e.,
 $\frac{\partial\mathcal{L}^{[q+1]}_{\bm{\eta}}}{\partial\bm{\eta}}=-\frac{2}{\sigma^{2}}\Re\left\{\mathbf{Q}^{\rm H}\bar{\bm{y}}\right\}+\frac{2}{\sigma^{2}}\Re\left\{\mathbf{Q}^{\rm H}\mathbf{Q}\right\}\bm{\eta}+\bm{\eta}=\bm{0}_{L\times 1},$ (12)
-we obtain
+we can solve
 $\hat{\bm{\eta}}^{[q+1]}=\left[\Re\left\{\mathbf{Q}^{\rm H}\mathbf{Q}\right\}+\frac{1}{2}\sigma^{2}\mathbf{I}_{L}\right]^{-1}\Re\left\{\mathbf{Q}^{\rm H}\bar{\bm{y}}\right\}.$ (13)
 III-B CFO Estimation
-To estimate the CFO, we solve the following unconstrained problem
+To estimate the CFO, we solve the following unconstrained problem:
 $\hat{\phi}\propto\arg\min_{\phi}\mathcal{L}_{\phi}\propto\arg\min_{\phi}\|\bm{y}-\bm{\mu}\|^{2}.$ (14)
-In order to enhance the tractability of the problem presented in (14) and obtain a closed-form solution, we employ a Taylor series approximation once more.
-The $n$-th term on the diagonal of $\bm{\Lambda}_{\phi}$ can be approximated as
+To enhance the tractability of (14) and derive a closed-form solution, the first-order Taylor approximation is employed. The $n$-th term on the diagonal of $\bm{\Lambda}_{\phi}$ can be approximated as
 $\left[\bm{\Lambda}_{\phi}\right]_{n,n}=e^{j2\pi\frac{n}{N}\phi}\approx e^{j2\pi\frac{n}{N}\hat{\phi}}+\left(\phi-\hat{\phi}\right)\frac{j2\pi}{N}e^{j2\pi\frac{n}{N}\hat{\phi}}.$ (15)
@@
 where $\bm{d}\triangleq\sqrt{P}\hat{\bm{\Lambda}}_{\bm{\theta}}^{[q+1]}\mathbf{S}^{\top}\hat{\bm{h}}^{[q]}$.
-Upon setting the gradient of (17) equal to zero, and solving for $\phi$, we can obtain the closed-form expression of $\hat{\phi}$ at the $(q+1)$-th iteration as
+Upon setting the partial derivative of (17) with regard to $\phi$ to zero and solving it for $\phi$, we acquire the closed-form solution of $\hat{\phi}$ at the $(q+1)$-th iteration as
 $\hat{\phi}^{[q+1]}=\hat{\phi}^{[q]}+\frac{\Re\left\{\left(\bm{y}-\hat{\bm{\Lambda}}_{\phi}^{[q]}\bm{d}\right)^{\rm H}\tilde{\bm{\Lambda}}_{\phi}^{[q]}\bm{d}\right\}}{\bm{d}^{\rm H}\tilde{\bm{\Lambda}}_{\phi}^{[q]{\rm H}}\tilde{\bm{\Lambda}}_{\phi}^{[q]}\bm{d}}.$ (18)
@@
 For position estimation, we employ the GD algorithm.
-Due to the similarity in the derivative expressions of the objective function with respect to three positional parameters $d_{Ru}$, $\phi_{az}$, and $\phi_{el}$, we use the symbol $\xi$ to represent any one of them.
-The expression for the partial derivative of the objective function with respect to $\xi$ is as follows
+Due to the same structures in the respective derivatives of the objective function with regard to the three position parameters $d_{Ru}$, $\phi_{az}$, and $\phi_{el}$, we use the symbol $\xi$ to represent any one of them.
+The partial derivative of the objective function with regard to $\xi$ is given by
 $\frac{\partial\mathcal{L}_{\xi}}{\partial\xi}=\frac{2}{\sigma^{2}}\Re\left\{\left(P\bm{h}^{\rm H}\mathbf{S}^{*}\mathbf{S}^{\top}-\sqrt{P}\bm{y}^{\rm H}\bm{\Lambda}_{\bm{\theta}}\bm{\Lambda}_{\phi}\mathbf{S}^{\top}\right)\frac{\partial\bm{h}}{\partial\xi}\right\},$ (19)
-where each term of $\frac{\partial\bm{h}}{\partial\xi}$ can be calculated as follows
+where $\frac{\partial\bm{h}}{\partial\xi}$ is calculated by
 $\frac{\partial\bm{h}}{\partial\xi}[n]$ $=\frac{\partial\bm{h}[n]}{\partial\xi}=-\sum_{r=1}^{N_{R}}\sqrt{\rho_{aR,n,r}}\sqrt{\rho_{Ru,n,r}}\bm{w}[r]$
@@
 $\left.\quad+(d_{Ru}\sin\phi_{el}\sin\phi_{az}-y_{r})d_{Ru}\sin\phi_{az}\cos\phi_{el}\right.$
-$\left.\quad+(d_{Ru}cos\phi_{el}-z_{r})(-d_{Ru}\sin\phi_{el})\right].$ (23)
-The general iterative joint estimation algorithm is outlined in Algorithm 1.
-Algorithm 1 Iterative joint estimation algorithm.
-1:  Optimize RIS phase shift $\bm{w}$ using the method to be introduced in Section V.
-2:  Initialize estimation of normalized CFO $\hat{\phi}$ and UE position $\hat{\bm{p}}_{u}$, step length $k_{p}$, convergence accuracy for outer loop $\epsilon_{o}$, convergence accuracy for inner loop $\epsilon_{i}$, and calculate corresponding polar coordinate of $\hat{\bm{p}}_{u}$ using (1).
-3:  repeat
-4:     Update $\hat{\bm{\eta}}$ with $\hat{\phi}$, $\hat{\bm{p}}_{u}$ being fixed using (13);
-5:     Update $\hat{\bm{\theta}}$ using (9);
-6:     Update $\hat{\phi}$ with $\hat{\bm{\theta}}$, $\hat{\bm{p}}_{u}$ being fixed using (18);
-7:     repeat
-8:        $\hat{d}_{Ru}\leftarrow\hat{d}_{Ru}-\frac{\partial\mathcal{L}_{\xi}}{\partial d_{Ru}}k_{p}$;
-9:        $\hat{\phi}_{az}\leftarrow\hat{\phi}_{az}-\frac{\partial\mathcal{L}_{\xi}}{\partial\phi_{az}}k_{p}$;
-10:        $\hat{\phi}_{el}\leftarrow\hat{\phi}_{el}-\frac{\partial\mathcal{L}_{\xi}}{\partial\phi_{el}}k_{p}$;
-11:     until $\mathcal{L}^{[p]}-\mathcal{L}^{[p+1]}\leq\epsilon_{i}$ after the $p$-th iteration.
-12:  until $\mathcal{L}^{[q]}-\mathcal{L}^{[q+1]}\leq\epsilon_{o}$ after the $q$-th iteration.
-$\epsilon_{i}$ and $\epsilon_{o}$ are both pre-set convergence accuracy and should vary with changes in transmit power.
-13:  Calculate Cartesian coordinate of $\hat{\bm{p}}_{u}$ using (2).
-III-D Estimation Performance Evaluation
-For position estimation, we use root mean square error (RMSE) as the performance metric. For any testing UE position $\bm{p}_{u}$, ${\rm RMSE}_{u}$ can be expressed as
+$\left.\quad+(d_{Ru}\cos\phi_{el}-z_{r})(-d_{Ru}\sin\phi_{el})\right].$ (23)
+III-D Overall Algorithm
+The proposed joint estimation of the CFO, PN and UE position is summarized in Algorithm 1.
+Specifically, we initialize the parameters, including the initial estimation of the normalized CFO $\hat{\phi}$ and UE position $\hat{\bm{p}}_{u}$, the step size, $k_{p}$, the convergence accuracy for outer and inner loops, $\epsilon_{o}$ and $\epsilon_{i}$, and the maximum iteration number for the outer and inner loops, $N_{o}$ and $N_{i}$.
+After transforming the coordinates into the polar coordinates, we estimate the PN, CFO, and UE position alternately.
+Closed-form expressions are utilized to update the PN and CFO estimates in each iteration, while the GD algorithm is employed to estimate the UE position.
+This repeats until either the objective function $\mathcal{L}$ converges or the maximum number of iterations is reached.
+It is noteworthy that this AO-based algorithm takes the RIS phase shifts, i.e., $\bm{w}$, as input, which can affect the algorithm performance significantly.
+The configuration of the RIS phase shifts, $\bm{w}$, will be discussed in Section V.
+In Algorithm 1, the objective function is expected to be non-increasing throughout iterations when updating the CFO and PN parameters because the approximate functions in (III-A2) and (17) show convexity with respect to $\bm{\eta}$ and $\phi$, respectively.
+In Steps 9, 10 and 11, with the UE position estimated in the previous iteration as input, the GD algorithm can ensure a lower objective function.
+Since the posterior probability is no larger than $1$, the negative logarithmic function $\mathcal{L}(\bm{\lambda})$ is non-negative and therefore lower bounded.
+Algorithm 1 converges.
+The computational complexity for updating $\bm{\eta}$, $\bm{\theta}$, $\phi$, and position parameters $d_{Rt}$, $\phi_{az}$, and $\phi_{el}$ in Algorithm 1 are $\mathcal{O}(L^{3}+N^{2}+NL^{2})$, $\mathcal{O}(NL)$, $\mathcal{O}(N^{2})$, and $\mathcal{O}(\max\left\{N^{3},N_{R}\right\})$, respectively.
+Suppose the inner GD algorithm for updating the UE position converges after $K_{i}$ iterations.
+Consequently, the computational complexity of Algorithm 1 is $\mathcal{O}(\max\left\{L^{3}+N^{2}+NL^{2},K_{i}N^{3},K_{i}N_{R}\right\})$ per iteration.
+Algorithm 1 Proposed iterative joint estimation algorithm.
+1:  Input: $\bm{w}$;
+2:  Initialization: $\hat{\phi}$, $\hat{\bm{p}}_{u}$, $k_{p}$, $\epsilon_{o}$, $\epsilon_{i}$, $N_{o}$, $N_{i}$;
+3:  Calculate polar coordinate of UE using (1);
+4:  repeat
+5:     Given fixed $\hat{\bm{p}}_{u}$, update $\hat{\bm{\eta}}$ with $\hat{\phi}$ using (13);
+6:     Update $\hat{\bm{\theta}}$ using (9);
+7:     Given fixed $\hat{\bm{p}}_{u}$, update $\hat{\phi}$ with $\hat{\bm{\theta}}$ using (18);
+8:     repeat
+9:        $\hat{d}_{Ru}\leftarrow\hat{d}_{Ru}-\frac{\partial\mathcal{L}_{\xi}}{\partial d_{Ru}}k_{p}$;
+10:        $\hat{\phi}_{az}\leftarrow\hat{\phi}_{az}-\frac{\partial\mathcal{L}_{\xi}}{\partial\phi_{az}}k_{p}$;
+11:        $\hat{\phi}_{el}\leftarrow\hat{\phi}_{el}-\frac{\partial\mathcal{L}_{\xi}}{\partial\phi_{el}}k_{p}$;
+12:     until $\mathcal{L}^{[p]}-\mathcal{L}^{[p+1]}\leq\epsilon_{i}$ after the $p$-th iteration or $p\textgreater N_{i}$.
+13:  until $\mathcal{L}^{[q]}-\mathcal{L}^{[q+1]}\leq\epsilon_{o}$ after the $q$-th iteration or $q\textgreater N_{o}$.
+14:  Calculate Cartesian coordinate of UE using (2);
+15:  Output: $\hat{\phi}$, $\hat{\bm{\theta}}$, $\hat{\bm{p}}_{u}$.
+III-E Performance Metrics
+We specify the performance metrics for all estimation parameters in $\bm{\lambda}$ for the sake of evaluating Algorithm 1.
+For the position estimation, we take root mean square error (RMSE) as the performance index.
+For any testing UE position $\bm{p}_{u}$, ${\rm RMSE}_{u}$ can be expressed as
 ${\rm RMSE}_{u}=\sqrt{\frac{1}{M_{e}}\sum_{m=1}^{M_{e}}\|\hat{\bm{p}}_{u,m}-\bm{p}_{u}\|^{2}},$ (24)
-where $M_{e}$ is the Monte Carlo experiment amount, and $\hat{\bm{p}}_{u,m}$ represents the estimation of UE position in the $m$-th trail.
-For CFO and PN estimation, it can be straightforwardly observed from (4) that, the received signal $\bm{y}$ is not altered under a common phase rotation $\epsilon$, i.e.,
-$\hat{\phi}\rightarrow\phi-\epsilon,\hat{\bm{\theta}}\rightarrow\bm{\theta}+\bm{\epsilon},$ (25)
+where $M_{e}$ is the number of Monte Carlo tests, and $\hat{\bm{p}}_{u,m}$ denotes the estimation of the UE position in the $m$-th trial.
+For the CFO and PN estimation, it can be straightforwardly observed from (6) that the received signal $\bm{y}$ does not change under a common phase rotation $\epsilon$, i.e.,
+$\hat{\phi}\rightarrow\phi-\epsilon\quad\text{and }\quad\hat{\bm{\theta}}\rightarrow\bm{\theta}+\bm{\epsilon},$ (25)
 where $\bm{\epsilon}[n]=\frac{2\pi n\epsilon}{N}$.
-Equation (25) indicates that there exists ambiguity in estimating CFO and PN.
-The ambiguity leads to difficulties in achieving high estimation accuracy for individual parameters of CFO and PN.
-However, in this paper, we are not focusing on the separate estimation accuracy of CFO and PN but rather on the estimation accuracy of UE position.
-Therefore, it is acceptable as long as the overall estimation accuracy of CFO and PN is reduced.
-In this paper, to evaluate the overall estimation accuracy of CFO and PN, we utilize a novel joint mean square error (MSE) measure, which can be calculated as
+Here, (25) indicates that there exists ambiguity in estimating the CFO and PN.
+The ambiguity leads to difficulties in achieving high estimation accuracy for individual parameters of the CFO and PN.
+Nevertheless, a high-precision joint estimation of CFO and PN is achievable, which is beneficial for estimating the UE position and will be illustrated in section VI.
+To evaluate the joint estimation performance, we utilize a joint MSE measure, which can be calculated as
 ${\rm MSE}_{\phi,\bm{\theta}}=\mathbb{E}_{\phi,\bm{\theta}}\left(\|\bm{\underline{\gamma}}-\hat{\bm{\underline{\gamma}}}\|^{2}\right),$ (26)
-where $\bm{\underline{\gamma}}=\bm{\gamma}-\bm{\gamma}[0]\bm{1}$, $\hat{\bm{\underline{\gamma}}}=\hat{\bm{\gamma}}-\hat{\bm{\gamma}}[0]\bm{1}$, $\bm{\gamma}=[\bm{\gamma}[0],\bm{\gamma}[1],\cdots,\bm{\gamma}[N-1]]^{\top}$, $\bm{\gamma}[n]=\bm{\theta}[n]+\frac{2\pi n\phi}{N}$, $\hat{\bm{\gamma}}=[\hat{\bm{\gamma}}[0],\hat{\bm{\gamma}}[1],\cdots,\hat{\bm{\gamma}}[N-1]]^{\top}$, $\hat{\bm{\gamma}}[n]=\hat{\bm{\theta}}[n]+\frac{2\pi n\hat{\phi}}{N}$.
+where $\bm{\underline{\gamma}}=\bm{\gamma}-\bm{\gamma}[0]\bm{1}$, $\hat{\bm{\underline{\gamma}}}=\hat{\bm{\gamma}}-\hat{\bm{\gamma}}[0]\bm{1}$, $\bm{\gamma}=[\bm{\gamma}[0],\bm{\gamma}[1],\cdots,\bm{\gamma}[N-1]]^{\top}$, $\bm{\gamma}[n]=\bm{\theta}[n]+\frac{2\pi n\phi}{N}$, $\hat{\bm{\gamma}}=[\hat{\bm{\gamma}}[0],\hat{\bm{\gamma}}[1],\cdots,\hat{\bm{\gamma}}[N-1]]^{\top}$, and $\hat{\bm{\gamma}}[n]=\hat{\bm{\theta}}[n]+\frac{2\pi n\hat{\phi}}{N}$.
+Using this measure, the performance of the joint CFO-PN estimation rather than each of the individual variables alone is considered.
 IV Hybrid Cramér-Rao Lower Bound Analysis
-In this section, we derive HCRLB to evaluate the joint estimation, in the presence of a random parameter $\bm{\theta}$ in the set of parameters to be estimated.
-The HCRLB for $\bm{\lambda}$, the parameters to be estimated, is given by [7]
+The HCRLB for the joint estimation of the CFO, PN, and UE position is derived in this section.
+This helps quantify the performance lower bound of the proposed AO-based estimation algorithm, and also serves as the objective function for optimizing the RIS phase shifts, $\bm{w}$, in the next section.
+The HCRLB for the estimation of $\bm{\lambda}$ is given by [34, Eq.
+(4.610)]
 $\mathbb{E}_{\bm{y},\bm{\theta}\mid\phi,d_{Ru},\phi_{az},\phi_{el}}\left[(\bm{\lambda}-\hat{\bm{\lambda}})(\bm{\lambda}-\hat{\bm{\lambda}})^{\top}\right]\succeq\mathbf{B}^{\rm-1},$ (27)
-where $\mathbf{B}$ is the Bayesian information matrix (BIM), which can be calculated as
-$\mathbf{B}=\mathbb{E}_{\bm{\theta}}\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]+\mathbb{E}_{\bm{\theta}}\left[-\frac{\partial}{\partial\bm{\lambda}}\left(\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}\right)^{\top}\right].$ (28)
-In order to evaluate the positioning accuracy and the CFO-PN joint MSE, we define a new parameter vector $\bm{\alpha}=\bm{g}(\bm{\lambda})\triangleq\left[\bm{\underline{\gamma}},x_{u},y_{u},z_{u}\right]^{\top}$.
-In the following subsections, we first derive the expression for matrix $\mathbf{B}$.
-Then, we derive the transition matrix to transform $\mathbf{B}$ into the HCRLB relative to the parameter vector $\bm{\alpha}$.
+where $\mathbf{B}$ is the Bayesian information matrix (BIM), which can be given by [34, Eq. (4.611)]
+$\mathbf{B}=\mathbb{E}_{\bm{\theta}}\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]+\mathbb{E}_{\bm{\theta}}\left[-\frac{\partial}{\partial\bm{\lambda}}\left(\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}\right)^{\top}\right],$ (28)
+where $\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})$ stands for the Fisher’s information matrix (FIM) for $\bm{\lambda}$ under the observation of received signal $\bm{y}$ at the UE, and $\mathbb{E}_{\bm{\theta}}\left[-\frac{\partial}{\partial\bm{\lambda}}\left(\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}\right)^{\top}\right]$ is the prior information matrix.
+Interested readers may refer to [34] for more details.
 IV-A Derivation of $\mathbb{E}_{\bm{\theta}}\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]$
-To derive $\mathbb{E}_{\bm{\theta}}\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]$, we begin by calculating the Fisher information matrix (FIM) for $\bm{\lambda}$, as expressed by
+To derive $\mathbb{E}_{\bm{\theta}}\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]$ in (28), we first analyze the FIM for $\bm{\lambda}$, as given by
 $\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]_{i,j}=\frac{2}{\sigma^{2}}\Re\left\{\frac{\partial\bm{\mu}{\rm{}^{\rm H}}}{\partial\lambda_{i}}\frac{\partial\bm{\mu}}{\partial\lambda_{j}}\right\}.$ (29)
-To calculate (29), the following gives the derivative of $\bm{\mu}$ with respect to each element in vector $\bm{\lambda}$.
-$\frac{\partial\bm{\mu}}{\partial\phi}$ $=\sqrt{P}\bm{\Lambda}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\bm{h},$ (30a)
-$\frac{\partial\bm{\mu}}{\partial\bm{\theta}[n]}$ $=\sqrt{P}{\rm Diag}(\bm{a}[n])\bm{\Lambda}_{\phi}\mathbf{S}^{\top}\bm{h},$ (30b)
-$\frac{\partial\bm{\mu}}{\partial d_{Ru}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial d_{Ru}},$ (30c)
-$\frac{\partial\bm{\mu}}{\partial\phi_{az}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial\phi_{az}},$ (30d)
+To calculate (29), the first-order derivative of $\bm{\mu}$ with respect to each element in vector $\bm{\lambda}$ is derived as follows.
+$\frac{\partial\bm{\mu}}{\partial\phi}$ $=\sqrt{P}\bm{\Lambda}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\bm{h};$ (30a)
+$\frac{\partial\bm{\mu}}{\partial\bm{\theta}[n]}$ $=\sqrt{P}{\rm Diag}(\bm{a}[n])\bm{\Lambda}_{\phi}\mathbf{S}^{\top}\bm{h};$ (30b)
+$\frac{\partial\bm{\mu}}{\partial d_{Ru}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial d_{Ru}};$ (30c)
+$\frac{\partial\bm{\mu}}{\partial\phi_{az}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial\phi_{az}};$ (30d)
 $\frac{\partial\bm{\mu}}{\partial\phi_{el}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial\phi_{el}},$ (30e)
-where $\bm{\Lambda}$ is a diagnal matrix, $\left[\bm{\Lambda}\right]_{n,n}=\frac{j2\pi n}{N}$, $\bm{a}[n]=[\bm{0}_{1\times n},je^{j\bm{\theta}[n]},\bm{0}_{1\times(N-n-1)}]^{\top}$.
-It is not difficult to observe that in (30), the terms related to $\phi$ and $\bm{\theta}$ all appear in the exponential terms of the diagonal elements of the corresponding matrix.
-Hence, they would disappear when calculating $\frac{\partial\bm{\mu}{\rm{}^{\rm H}}}{\partial\lambda_{i}}\frac{\partial\bm{\mu}}{\partial\lambda_{j}}$ in (29).
-That is to say, the calculated FIM is not related to parameter $\bm{\theta}$, and
+where $\bm{\Lambda}$ is a diagnal matrix, $\left[\bm{\Lambda}\right]_{n,n}=\frac{j2\pi n}{N}$, and $\bm{a}[n]=[\bm{0}_{1\times n},je^{j\bm{\theta}[n]},\bm{0}_{1\times(N-n-1)}]^{\top}$.
+It is not difficult to see that in (30), only $\bm{\Lambda}_{\phi}$, $\bm{\Lambda}_{\bm{\theta}}$ and ${\rm Diag}(\bm{a}[n])$ are related to $\phi$ and $\bm{\theta}$, which are all diagonal matrices.
+Additionally, both $\phi$ and $\bm{\theta}$ appear in the exponent of diagonal elements in these matrices.
+Hence, they can be suppressed when calculating $\frac{\partial\bm{\mu}{\rm{}^{\rm H}}}{\partial\lambda_{i}}\frac{\partial\bm{\mu}}{\partial\lambda_{j}}$ in (29).
+In other words, the calculated FIM does not rely on the parameter $\bm{\theta}$, i.e.,
 $\mathbb{E}_{\bm{\theta}}\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]=\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda}).$ (31)
+Take $\frac{\partial\bm{\mu}^{\rm H}}{\partial d_{Ru}}\frac{\partial\bm{\mu}}{\partial d_{Ru}}$ as an example. We have
+$\frac{\partial\bm{\mu}^{\rm H}}{\partial d_{Ru}}\frac{\partial\bm{\mu}}{\partial d_{Ru}}$ $=\sqrt{P}\frac{\partial\bm{h}^{\rm H}}{\partial d_{Ru}}\mathbf{S}^{*}\bm{\Lambda}_{\bm{\theta}}^{\rm{H}}\bm{\Lambda}_{\phi}^{\rm H}\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial d_{Ru}}$ (32)
+$=P\frac{\partial\bm{h}^{\rm H}}{\partial d_{Ru}}\mathbf{S}^{*}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial d_{Ru}}$
+which is unrelated to $\phi$ or $\bm{\theta}$, since $\bm{\Lambda}_{\phi}^{\rm H}\bm{\Lambda}_{\phi}=\bm{\Lambda}_{\bm{\theta}}^{\rm{H}}\bm{\Lambda}_{\bm{\theta}}=\mathbf{I}_{N}$.
+This can be extended to $\frac{\partial\bm{\mu}{\rm{}^{\rm H}}}{\partial\lambda_{i}}\frac{\partial\bm{\mu}}{\partial\lambda_{j}},\forall i,j$, and is omitted here.
 IV-B Derivation of $\mathbb{E}_{\bm{\theta}}\left[-\frac{\partial}{\partial\bm{\lambda}}\left(\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}\right)^{\top}\right]$
-Since $\bm{\theta}\sim\mathcal{N}(\bm{0}_{N\times 1},\bm{\Psi})$, we can calculate $\log p(\bm{\theta})$ by $\log p(\bm{\theta})=-\frac{1}{2}\bm{\theta}^{\top}\bm{\Psi}^{-1}\bm{\theta}-\frac{N}{2}\log(2\pi)-\frac{1}{2}\log\det\bm{\Psi}$, and its partial to $\bm{\lambda}$ can be calculated by
+Since $\bm{\theta}\sim\mathcal{N}(\bm{0}_{N\times 1},\bm{\Psi})$, we can calculate $\log p(\bm{\theta})$ as $\log p(\bm{\theta})=-\frac{1}{2}\bm{\theta}^{\top}\bm{\Psi}^{-1}\bm{\theta}-\frac{N}{2}\log(2\pi)-\frac{1}{2}\log\det\bm{\Psi}$, and its partial derivative with respect to $\bm{\lambda}$ can be calculated by
 $\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}=\left[0,-(\bm{\Psi}^{-1}\bm{\theta})^{\top},0,0,0\right]^{\top}.$ (33)
@@
 IV-C Derivation of the Transformed HCRLB
-Due to the ambiguity between the parameters to be estimated, the MSE of CFO and PN is jointly calculated using (26), and when computing the position error bound (PEB), the position parameters are required to be in the form of 3D Cartesian coordinates $x_{u}$, $y_{u}$, and $z_{u}$.
-Therefore, the parameter vector $\bm{\lambda}$ needs to be transformed into $\bm{\alpha}$.
-The HCRLB for $\bm{\alpha}$ is given by [9]
+So far, we have derived the expression for $\mathbf{B}$, which is an HCRLB matrix that gives analytical estimation accuracy lower bounds for the parameters in $\bm{\lambda}$.
+However, when assessing the proposed estimation algorithm, we adopt the RMSE and CFO-PN joint MSE metrics introduced in Section III-E.
+Hence, we define a new parameter vector $\bm{\alpha}=\bm{g}(\bm{\lambda})\triangleq\left[\bm{\underline{\gamma}},x_{u},y_{u},z_{u}\right]^{\top}$, corresponding to the two metrics.
+Then, $\mathbf{B}$ can be transformed into the $\bm{\alpha}$-related HCRLB matrix by [35]
 $\mathbf{HCRLB}(\bm{p}_{u},\bm{w})=\bm{\Xi}(\bm{p}_{u})\mathbf{B}^{-1}(\bm{p}_{u},\bm{w})\bm{\Xi}^{\top}({\bm{p}_{u}}),$ (35)
-where for any UE position $\bm{p}_{u}$, $\mathbf{\Xi}(\bm{p}_{u})$ is the transition matrix defined as
+where, for any UE position $\bm{p}_{u}$, $\mathbf{\Xi}(\bm{p}_{u})$ is the transition matrix defined as
 $\bm{\Xi}(\bm{p}_{u})=\frac{\partial\bm{g}(\bm{\lambda})}{\partial\bm{\lambda}}={\rm Blkdiag}\left(\bm{\Xi}_{1},\bm{\Xi}_{2}(\bm{p}_{u})\right),$ (36)
-and $\bm{\Xi}_{1}$, $\bm{\Xi}_{2}(\bm{p}_{u})$ can be calculated by (36) and (37).
+with $\bm{\Xi}_{1}$ and $\bm{\Xi}_{2}(\bm{p}_{u})$ given by (37) and (38), respectively.
 $\bm{\Xi}_{1}=\begin{bmatrix}\frac{\partial\bm{\underline{\gamma}}[0]}{\partial\phi}&\frac{\partial\bm{\underline{\gamma}}[0]}{\partial\bm{\theta}[0]}&\cdots&\frac{\partial\bm{\underline{\gamma}}[0]}{\partial\bm{\theta}[N-1]}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{\partial\bm{\underline{\gamma}}[N-1]}{\partial\phi}&\frac{\partial\bm{\underline{\gamma}}[N-1]}{\partial\bm{\theta}[0]}&\cdots&\frac{\partial\bm{\underline{\gamma}}[N-1]}{\partial\bm{\theta}[N-1]}\end{bmatrix}=\begin{bmatrix}0&0&0&\cdots&0\\ \frac{2\pi}{N}&-1&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \frac{2\pi(N-1)}{N}&-1&0&\cdots&1\\ \end{bmatrix}\in\mathbb{R}^{N\times(N+1)},$ (37)
 $\bm{\Xi}_{2}(\bm{p}_{u})=\begin{bmatrix}\frac{\partial x_{u}}{\partial d_{Ru}}&\frac{\partial x_{u}}{\partial\phi_{az}}&\frac{\partial x_{u}}{\partial\phi_{el}}\\ \frac{\partial y_{u}}{\partial d_{Ru}}&\frac{\partial y_{u}}{\partial\phi_{az}}&\frac{\partial y_{u}}{\partial\phi_{el}}\\ \frac{\partial z_{u}}{\partial d_{Ru}}&\frac{\partial z_{u}}{\partial\phi_{az}}&\frac{\partial z_{u}}{\partial\phi_{el}}\\ \end{bmatrix}=\begin{bmatrix}\sin\phi_{el}\cos\phi_{az}&-d_{Ru}\sin\phi_{el}\sin\phi_{az}&d_{Ru}\cos\phi_{az}\cos\phi_{el}\\ \sin\phi_{el}\sin\phi_{az}&d_{Ru}\sin\phi_{el}\cos\phi_{az}&d_{Ru}\sin\phi_{az}\cos\phi_{el}\\ \cos\phi_{el}&0&-d_{Ru}\sin\phi_{el}\end{bmatrix}\in\mathbb{R}^{3\times 3}.$ (38)
-After obtaining $\mathbf{HCRLB}(\bm{p}_{u},\bm{w})$, $\rm PEB$ for position estimation and $\rm HCRLB_{CFO-PN}$ for CFO-PN estimation can be calculated as
-${\rm PEB}=\sqrt{{\rm tr}([\mathbf{HCRLB}(\bm{p}_{u},\bm{w})]_{N:N+2,N:N+2})},$ (38)
+After obtaining $\mathbf{HCRLB}(\bm{p}_{u},\bm{w})$, the position error bound (${\rm PEB}$) for evaluating the UE position estimation and $\rm HCRLB_{CFO-PN}$ for evaluating the CFO-PN joint estimation can be given by
+${\rm PEB}=\sqrt{{\rm tr}([\mathbf{HCRLB}(\bm{p}_{u},\bm{w})]_{N:N+2,N:N+2})};$ (39)
 ${\rm HCRLB_{CFO-PN}}={\rm tr}([\mathbf{HCRLB}(\bm{p}_{u},\bm{w})]_{0:N-1,0:N-1}).$ (40)
-Both ${\rm PEB}$ and ${\rm HCRLB_{CFO-PN}}$ are related to UE’s location $\bm{p}_{u}$ and RIS phase shift $\bm{w}$.
-V RIS Phase Shift Optimization
-In this section, our goal is to optimize the phase shift of the RIS to enhance theoretical positioning accuracy, i.e., PEB.
-According to equation (38), minimizing the PEB at a specific location requires knowledge of that location.
-However, this is not possible in practical positioning scenarios.
-To address this issue, we select multiple sampling points within the AOI randomly, then calculate and minimize their average PEB.
-Accordingly, the optimization problem for RIS phase shift can be expressed as
+Both ${\rm PEB}$ and ${\rm HCRLB_{CFO-PN}}$ depend on the UE’s position, $\bm{p}_{u}$, and the RIS phase shifts, $\bm{w}$.
+V RIS Configuration
+In this section, we further configure the phase shifts $\bm{w}$ of the RIS to minimize the analytical lower bound of the localization accuracy, i.e., PEB.
+According to (39), the PEB depends on the UE position, which indicates that minimizing the PEB at a specific UE position requires knowledge of that position.
+However, this is not possible in practice.
+To address this issue and enhance the algorithm’s robustness, we take the average PEB within the area of interest (AOI) as the objective function.
+Specifically, we randomly sample $U$ positions within the AOI and calculate their average PEB.
+Accordingly, the problem of interest is formulated as
 $\min_{\bm{w}}\quad\frac{1}{U}\sum_{u=1}^{U}{\rm PEB}(\bm{p}_{u},\bm{w})$
 $s.t.\quad|\bm{w}[r]|=1,r=1,\cdots,N_{R}.$ (41)
-Problem (V) is non-convex because the objective function does not exhibit convexity with respect to vector $\bm{w}$.
-We define $\mathbf{W}\triangleq\bm{\bar{w}}\bm{\bar{w}}^{\rm H}$, where $\bm{\bar{w}}\triangleq\bm{w}^{*}$, which needs to satisfy $\mathbf{W}\succcurlyeq 0$ and ${\rm rank}(\mathbf{W})=1$.
+Problem (V) is not convex because of the non-convexity of the objective function with respect to vector $\bm{w}$, as well as the unit-module constraint.
+To convexify (V), we first define $\mathbf{W}\triangleq\bm{\bar{w}}\bm{\bar{w}}^{\rm H}$, where $\bm{\bar{w}}\triangleq\bm{w}^{*}$, which satisfies $\mathbf{W}\succcurlyeq 0$ and ${\rm rank}(\mathbf{W})=1$.
 Since the rank-1 constraint is non-convex, we employ the SDR method to relax this constraint.
-Consequently, the problem (V) can be transformed into
+Consequently, the problem in (V) can be transformed into
 $\min_{\mathbf{W}}\quad\frac{1}{U}\sum_{u=1}^{U}{\rm PEB}(\bm{p}_{u},\mathbf{W})$
 $\begin{array}[]{ll}s.t.&\left[\mathbf{W}\right]_{r,r}=1,r=1,\cdots,N_{R},\\ &\mathbf{W}\succcurlyeq 0.\end{array}$ (44)
-Inspired by [12], we utilize the property of the Schur complement [6], and the problem can be expressed in the form of SDP.
-For each calculated PEB at a particular position $\bm{p}_{u}$, we define an auxiliary matrix $\mathbf{Z}_{u}$ that satisfies the following expression
+According to (35) and (39), the optimization problem in (V) remains non-convex because of the complex objective function.
+To transform (V) into the convex SDP form, for each calculated PEB at a particular position $\bm{p}_{u}$, we define an auxiliary matrix $\mathbf{Z}_{u}$ that satisfies
 $\mathbf{Z}_{u}\succcurlyeq\tilde{\bm{\Xi}}(\bm{p}_{u})\mathbf{B}^{-1}(\bm{p}_{u},\mathbf{W})\tilde{\bm{\Xi}}^{\top}(\bm{p}_{u}),$ (45)
-where $\tilde{\bm{\Xi}}(\bm{p}_{u})\triangleq\mathbf{T}\mathbf{\Xi}(\bm{p}_{u})$, and $\mathbf{T}\triangleq[\bm{0}_{N\times N}\quad\mathbf{I}_{3}]\in\mathbb{R}^{(N+3)\times(N+3)}$ can extract the PEB portion from the HCRLB.
-The optimization problem in (V) can be converted to
+where $\tilde{\bm{\Xi}}(\bm{p}_{u})\triangleq\mathbf{T}\mathbf{\Xi}(\bm{p}_{u})$, and $\mathbf{T}\triangleq[\bm{0}_{N\times N}\quad\mathbf{I}_{3}]\in\mathbb{R}^{(N+3)\times(N+3)}$ can extract the PEB from the HCRLB.
+According to the property of the Schur complement and the positive semidefiniteness of $\mathbf{B}(\bm{p}_{u},\mathbf{W})$, (45) can be transformed to
+$\begin{bmatrix}\mathbf{Z}_{u}&\tilde{\bm{\Xi}}(\bm{p}_{u})\\ \tilde{\bm{\Xi}}^{\top}(\bm{p}_{u})&\mathbf{B}(\bm{p}_{u},\mathbf{W})\end{bmatrix}\succcurlyeq 0.$ (46)
+Thus, the problem in (V) can be converted to
 $\min_{\mathbf{W},\mathbf{Z}_{1},\cdots,\mathbf{Z}_{U}}\quad\frac{1}{U}\sum_{u=1}^{U}{\rm tr}(\mathbf{Z}_{u})$
-$\begin{array}[]{ll}s.t.&\left[\mathbf{W}\right]_{r,r}=1,r=1,\cdots,N_{R},\\ &\mathbf{W}\succcurlyeq 0,\\ &\displaystyle\frac{1}{U}\sum_{u=1}^{U}\begin{bmatrix}\mathbf{Z}_{u}&\tilde{\bm{\Xi}}(\bm{p}_{u})\\ \tilde{\bm{\Xi}}^{\top}(\bm{p}_{u})&\mathbf{B}(\bm{p}_{u},\mathbf{W})\end{bmatrix}\succcurlyeq 0.\end{array}$ (48)
-Problem (V) is a convex SDP problem, and can be optimally solved using established convex optimization solvers, such as CVX [30].
-Since there is no rank-1 constraint in the optimization problem, it is highly likely that the obtained matrix $\mathbf{W}$ will have a rank greater than one.
-Therefore, decomposing it into the desired RIS phase shift vector $\bm{w}$ is a crucial subsequent problem.
-As in [29], we first perform an eigenvalue decomposition on the optimal $\mathbf{W}$, then take the largest eigenvector and multiply it by the square root of the largest eigenvalue and normalize the vector, obtaining $\bm{\bar{w}}$.
-Finally, we calculate the RIS phase shift by $\bm{w}=\bm{\bar{w}}^{*}$.
-The decomposition process can also be accomplished using the Gaussian randomization method as in [28].
-In order to meet the requirements of (V), we next derive the expression of $\mathbf{B}$ that is related to $\mathbf{W}$. We first rewrite (4) as
+$\begin{array}[]{ll}s.t.&\left[\mathbf{W}\right]_{r,r}=1,r=1,\cdots,N_{R},\\ &\mathbf{W}\succcurlyeq 0,\\ &\displaystyle\frac{1}{U}\sum_{u=1}^{U}\begin{bmatrix}\mathbf{Z}_{u}&\tilde{\bm{\Xi}}(\bm{p}_{u})\\ \tilde{\bm{\Xi}}^{\top}(\bm{p}_{u})&\mathbf{B}(\bm{p}_{u},\mathbf{W})\end{bmatrix}\succcurlyeq 0,\end{array}$ (50)
+where, according to the relationship in (45) and the calculation of PEB in (39), minimizing the objective function results in minimizing the average square of PEB, which also minimizes the initial objective function, the average PEB.
+Proposition 1. $\mathbf{B}(\bm{p}_{u},\mathbf{W})$ is linear to $\mathbf{W}$.
+Proof.
+According to (28), only $\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})$ depends on $\bm{w}$ in $\mathbf{B}$.
+Hence, we only need to prove the linearity of $\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})$ to $\mathbf{W}$, which can be rewritten as
+$\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]_{i,j}=\frac{2}{\sigma^{2}}\sum_{n=0}^{N-1}\Re\left\{\frac{\partial\bm{\mu}[n]^{*}}{\partial\lambda_{i}}\frac{\partial\bm{\mu}[n]}{\partial\lambda_{j}}\right\}.$ (51)
+To derive the explicit expressions of $\frac{\partial\bm{\mu}[n]^{*}}{\partial\lambda_{i}},\forall i$ and $\frac{\partial\bm{\mu}[n]}{\partial\lambda_{j}},\forall j$ with respect to $\bar{\bm{w}}$, we rewrite (6) as
 $\bm{y}=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\mathbf{G}^{\top}\bm{w}+\bm{v},$ (52)
-which separates $\bm{w}$ from $\bm{h}$, with $\mathbf{G}\in\mathbb{C}^{N_{R}\times N}$ and $\left[\mathbf{G}\right]_{r,n}=\sqrt{\rho_{aR,n,r}}\sqrt{\rho_{Ru,n,r}}e^{-j2\pi\frac{nB}{N}(\tau_{ar}+\tau_{ru})}$.
+with $\mathbf{G}\in\mathbb{C}^{N_{R}\times N}$ and $\left[\mathbf{G}\right]_{r,n}=\sqrt{\rho_{aR,n,r}}\sqrt{\rho_{Ru,n,r}}e^{-j2\pi\frac{nB}{N}(\tau_{ar}+\tau_{ru})}$.
 Correspondingly, $\bm{\mu}$ can be rewritten as $\bm{\mu}=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\mathbf{G}^{\top}\bm{\bar{w}}^{*}$.
@@
 $\bm{\mu}[n]^{*}=\sqrt{P}e^{-j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}]_{:,n}^{\rm H}\mathbf{G}^{\rm H}\bm{\bar{w}}.$ (54)
-According to (28), only the $\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})$ part is related to $\bm{w}$, hence we only need to derive the relationship of this part of $\mathbf{B}$ and $\mathbf{W}$.
-We can rewrite the FIM in a summation form:
-$\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]_{i,j}=\frac{2}{\sigma^{2}}\sum_{n=0}^{N-1}\Re\left\{\frac{\partial\bm{\mu}[n]^{*}}{\partial\lambda_{i}}\frac{\partial\bm{\mu}[n]}{\partial\lambda_{j}}\right\}.$ (52)
-The partial derivatives of $\bm{\mu}[n]$ and $\bm{\mu}[n]^{*}$ with respect to all parameters to be estimated are derived in Appendix A.
-It is observed that the expressions for $\frac{\partial\bm{\mu}[n]^{*}}{\partial\lambda_{i}}$ all end with $\bm{\bar{w}}$, while the expressions for $\frac{\partial\bm{\mu}[n]}{\partial\lambda_{j}}$ all start with $\bm{\bar{w}}^{\rm H}$ for the vector parts.
-Therefore, when calculating the FIM, we can easily obtain the expression with respect to $\mathbf{W}$.
-VI Numerical Results
-This section presents comprehensive simulations conducted to assess the performance of the proposed estimation and optimization algorithms.
-We consider a 3D Cartesian coordinate system, where the AN is located at $\bm{p}_{a}=[\rm 2m,-2m,1m]^{\top}$, and the reference point of RIS is located at $\bm{0}_{3\times 1}$.
-Without loss of generality, RIS is placed on the $yz$ plane.
-We assume AOI in a cube centered at $[\rm 2m,2m,0m]^{\top}$ with the edge length of $\rm 1m$.
-We set light speed $c=3\times 10^{8}{\rm m/s}$, noise power $\sigma^{2}=-109{\rm dBm}$.
-$N=32$ subcarriers are used in one OFDM symbol, and the central frequency and bandwidth of the OFDM symbol are set to $f_{c}=2.8{\rm GHz}$ and $B=100{\rm MHz}$, respectively.
-The normalized CFO $\phi$ is uniformly drawn from $[-0.15,0.15]$, and the length of shortened PN vector $L$ is set as $L=16$.
-Additionally, we make the assumption that the antenna gains, $G_{a}$ and $G_{t}$, are equal to $1$.
-VI-A Convergence analysis
-Figure 3: The convergence of Algorithm 1.
-To begin, we analyze the convergence of the proposed joint estimation algorithm.
+The partial derivatives of $\bm{\mu}[n]$ and $\bm{\mu}[n]^{*}$ with regard to all parameters to be estimated are derived in Appendix A.
+It can be observed that $\bm{\bar{w}}$ is only in the last term of $\frac{\partial\bm{\mu}[n]^{*}}{\partial\lambda_{i}},\forall i$, while $\bm{\bar{w}}^{\rm H}$ appears as the first term of $\frac{\partial\bm{\mu}[n]}{\partial\lambda_{j}},\forall j$, for the vector parts.
+Hence, the product of $\frac{\partial\bm{\mu}[n]^{*}}{\partial\lambda_{i}},\forall i$ and $\frac{\partial\bm{\mu}[n]}{\partial\lambda_{j}},\forall j$ would lead to $\mathbf{W}$.
+As a result, as can be seen from (51), every term of $\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})$ is linear to $\mathbf{W}$ because the operations of taking the real parts, summation, and scalar multiplication preserve linearity.
+Hence, $\mathbf{B}(\bm{p}_{u},\mathbf{W})$ is linear to $\mathbf{W}$.
+∎
+According to Proposition 1, (V) is a convex SDP problem, and can be optimally solved using established convex optimization solvers, such as off-the-shelf CVX toolboxes [36].
+Since the rank-1 constraint is relaxed via the SDR process, the obtained matrix $\mathbf{W}$ is likely to have a rank greater than 1.
+Therefore, decomposing it into the desired RIS phase shift vector $\bm{w}$ is a crucial subsequent problem.
+As done in [37], we first perform EVD on the optimal $\mathbf{W}$, then take the largest eigenvector and multiply it with the square root of the largest eigenvalue and normalize the vector, obtaining $\bm{\bar{w}}$.
+Finally, we calculate the RIS phase shifts by $\bm{w}=\bm{\bar{w}}^{*}$.
+The decomposition process can also be accomplished using the Gaussian randomization method, as done in [38].
+Being an SDP, the problem in (V) can be solved using the interior point method in the CVX toolboxes.
+Thus, the convergence of the proposed RIS configuration is guaranteed.
+The computational complexity of the proposed RIS configuration is primarily influenced by the interior point method employed to solve (V).
+The worst-case complexity of the interior point method is $\mathcal{O}(\max\left\{N_{con},N_{var}\right\}^{4}\sqrt{N_{v}})$, where $N_{con}$ and $N_{var}$ are the constraint and variable numbers, respectively [39].
+For the problem in (V), ${N_{con}}=N_{R}+2$ and ${N_{var}}=N_{R}^{2}+U(N+3)^{2}$.
+Suppose that the convergence accuracy is $\epsilon$.
+The complexity of solving (V) is $\mathcal{O}((N_{R}^{2}+UN^{2})^{4.5}\log\frac{1}{\epsilon})$.
+VI Numerical Evaluation
+This section elaborates on the comprehensive simulation results conducted to assess the proposed CFO, PN and position estimation and the RIS configuration algorithms.
+We adopt a 3D Cartesian coordinate system, where the AN is located at $\bm{p}_{a}=[\rm 2\>m,-2\>m,1\>m]^{\top}$, and the reference point of the RIS is located at $\bm{0}_{3\times 1}{\rm\>m}$.
+The RIS is placed on the $yz$ plane, without loss of generality.
+The AOI is set to be a cube centered at $[\rm 2\>m,2\>m,0\>m]^{\top}$ with the side length of $\rm 1\>m$.
+The light speed is $c=3\times 10^{8}\>{\rm m/s}$, and the noise power is $\sigma^{2}=-109\;{\rm dBm}$.
+$N=32$ subcarriers are used in an OFDM symbol, the central frequency is $f_{c}=2.8{\rm\>GHz}$, and the bandwidth is $B=100{\rm\>MHz}$.
+The antenna gains $G_{a}$ and $G_{t}$ are both set to $1$.
+Unless otherwise specified, $N_{R}=81$ and $\sigma_{\Delta}^{2}=10^{-3}\;{\rm rad^{2}}$ are used in the simulations.
+The length of the approximate PN vector $\bm{\eta}$ is $L=16$.
+The normalized CFO $\phi$ is uniformly drawn from $[-0.15,0.15]$, corresponding to a CFO of up to $15\;{\rm MHz}$.
+The CFO can be caused by frequency mismatches (resulting from severe hardware impairments) between the transmitter (the ANs) and receiver (the UE) when the UE is stationary.
+[^1] The CFO can also be a result of the Doppler shift, in which case the normalized CFO is expected to be within a much smaller range, e.g., $[-9.96\times 10^{-5},9.96\times 10^{-5}]$ when the UE moves at the speed of $\rm 120\;km/h$.
+The proposed localization algorithm can directly apply to this case due to the direct mapping between the Doppler shift and the normalized CFO.
+We gauge the performance of the proposed algorithms in comparison with the state-of-the-art approaches developed in [27], where a 2D near-field localization problem was solved with the assistance of a RIS.
+An MLE-based algorithm was designed to estimate the UE position, and an SDR-based optimization algorithm was proposed to minimize the CRLB.
+Note that the CFO and PN were not considered in [27].
+The RMSE and the CFO-PN joint MSE are utilized to assess the performance of the localization and joint CFO and PN estimation; see (24) and (26).
+VI-A Convergence of Algorithm 1
+(a) Convergence of the GD algorithm for estimating UE position.
+(b) Convergence of the overall algorithm.
+Figure 3: The convergence of the proposed Algorithm 1, where the transmit power is $P=-10{\rm\;dBm}$.
+First, we illustrate the convergence of Algorithm 1.
 Fig.
-3 illustrates the convergence of the algorithm in the case of transmit power of $-10{\rm dBm}$, which corresponds to a receiving signal-to-noise ratio (SNR) of approximately $17.5{\rm dB}$.
+3 illustrates the case when the transmit power is $-10{\rm\>dBm}$, which corresponds to a receive signal-to-noise ratio (SNR) of approximately $17.5{\rm\>dB}$.
 Specifically, Fig.
-3 (a) shows the variation of the objective function during each inner GD iteration in Algorithm 1, while Fig.
-3 (b) focuses solely on the variation of the objective function based on the results after every inner iteration.
+3(a) depicts the change of the objective function when operating the GD algorithm to estimate the UE position, and different curves correspond to the GD algorithm in different rounds of iteration.
+Fig.
+3(b) focuses on the convergence of the overall algorithm, and different points on the curves represent the objective function’s values after different rounds of iteration.
+In general, the GD algorithm typically converges within 2,000 iterations in this setup, which converges increasingly earlier within an iteration as the overall algorithm runs.
+The overall algorithm generally converges within 55 iterations.
+TABLE II: Average number of iterations for Algorithm 1
+Transmit Power (dBm) $-20$ $-10$ 00 $10$
+Avg. No. of Iterations 19.45 50.59 84.31 134.95
+Further, Table II shows the average number of iterations under different transmit powers for Algorithm 1.
+In this simulation, 100 tests are conducted, and $N_{o}$ and $N_{i}$ are set to be sufficiently large.
+The results indicate that the required iteration number increases as the transmit power increases.
+VI-B Performance of RIS Configuration
+(a) Random RIS phase shift
+(b) Approach developed in [27]
+(c) Proposed optimization algorithm
+Figure 4: The PEB within AOI for various RIS configuration strategies.
+Next, our attention is directed towards evaluating the performance of the proposed RIS configuration algorithm.
+Fig.
+4 illustrates the PEB near the AOI across various RIS phase shift design schemes, where the black squares represent the boundary of AOI.
+More precisely, we set the transmit power $P$ to $-20{\rm\>dBm}$, corresponding to a receive SNR of approximately $7.5{\rm\>dB}$.
+Without loss of generality, the selected positions in the figures are all located in the $xy$-plane.
+Fig.
+4 indicates that the proposed optimization algorithm improves the PEB by approximately two orders of magnitude by comparison with the random RIS phase shifts, and by approximately one order of magnitude compared to the situation where there are no hardware impairments [27].
+This fully demonstrates the significance of considering the CFO and PN in the localization of the UE, as well as the effectiveness, progressiveness, and superiority of the proposed RIS configuration algorithm.
+Additionally, observed from Fig.
+4, the optimized RIS phase shifts not only improve the localization accuracy within the AOI but also significantly improve localization performance in its surrounding areas.
+VI-C Estimation Accuracy
+Figure 5: The CDF of RMSE and PEB with different $N_{R}$ and transmit powers.
+Then, we concentrate on the CFO, PN, and UE position estimation accuracy.
+We first illustrate the overall localization accuracy.
+Fig.
+5 illustrates the cumulative distribution functions (CDFs) of RMSE and PEB for 500 random testing positions, under different transmit powers and different numbers of RIS elements.
+It is observed that increasing the RIS size and the transmit power results in significant enhancement of the overall localization accuracy.
+Furthermore, when the RIS size is $N_{R}=49$, the CDF of the RMSE deviates significantly from the CDF of the PEB, and does not achieve the theoretical localization accuracy.
+However, when the RIS size is increased to $N_{R}=121$, the CDF of the RMSE curve aligns more closely with the CDF of the PEB.
+This phenomenon can also be observed in Fig.
+8.
+Next, we select a fixed testing position with the coordinates $\rm[1.50\>m,2.15\>m,0.45\>m]^{\top}$ to investigate the impact of the transmit power $P$, the RIS size $N_{R}$, and the PN variance $\sigma_{\Delta}^{2}$ on the CFO, PN, and UE position estimation accuracy.
+Figure 6: The RMSE of $\bm{p}_{u}$ estimation under random and optimized RIS phase shifts.
+Figure 7: ${\rm MSE}_{\phi,\bm{\theta}}$ under random and optimized RIS phase shifts.
+In Figs.
+6 and 7, we assess the effect of the optimized RIS phase shifts on the performance of estimation, compared to the random setup.
 From Fig.
-3 (a), we observe that the inner GD iterations corresponding to each outer iteration converge earlier and earlier.
-This phenomenon occurs because the convergence for outer iteration approaches, the magnitude of the parameter updates decreases, and the variation of the posterior function becomes closer to the predefined stopping threshold.
-From Fig.
-3, it is not difficult to notice that the inner loop typically converges within 2000 iterations, while the outer loop generally converges within 40 iterations.
-VI-B Performance of RIS Phase Shift Optimization
-Figure 4: Apply random RIS phase shift and calculate the PEB of different positions.
-Figure 5: Apply the RIS phase shift optimization algorithm and calculate the PEB of different positions.
-In this subsection, our attention is directed toward evaluating the performance of the optimization algorithm.
-Fig.
-4 and Fig.
-5 illustrate the PEB near the AOI before and after the optimization of RIS phase shift, respectively.
-More precisely, we configure the number of RIS elements, denoted as $N_{R}$, to be $81$, and transmit power $P$ to be $-20{\rm dBm}$, which corresponds to a receiving SNR of approximately $7.5{\rm dB}$.
-Without loss of generality, the selected positions in the figures are located in the $xy$ plane.
-A comparison between the two figures reveals that the optimized PEB has improved by approximately two orders of magnitude, which fully demonstrates the effectiveness of the employed RIS optimization algorithm.
-Additionally, we have observed that the optimized RIS phase shift not only improves the theoretical positioning accuracy within the AOI, but also has a significant improvement in the surrounding areas of it.
-VI-C Localization Accuracy
-Figure 6: The CDF of RMSE and PEB with different $N_{R}$ and transmit powers.
-Figure 7: The RMSE of $\bm{p}_{u}$ estimation with random and optimized RIS phase shift.
+6, the localization performance with random RIS phase shifts does not experience a significant decline as the transmit power increases.
+At the other extreme, the optimized RIS phase shifts provide an improvement of approximately two orders of magnitude in localization accuracy.
+This demonstrates the crucial role of optimizing the RIS phase shifts in localization and underscores the effectiveness of the proposed RIS configuration optimization algorithm.
+On the other hand, Fig.
+7 shows that compared to the optimized RIS phase shifts, using random RIS phase shifts yields even lower joint MSE of CFO and PN, although the curve does not experience any significant decline as the transmit power increases in this case.
+This occurs because we extract the PEB part of HCRLB and set it as the RIS optimization objective.
+Thus, the estimation performance of the CFO and PN parameters is ignored.
 Figure 8: The RMSE of $\bm{p}_{u}$ estimation at $N_{R}=49$, $N_{R}=81$, and $N_{R}=121$.
 Figure 9: ${\rm MSE}_{\phi,\bm{\theta}}$ at $N_{R}=49$, $N_{R}=81$, and $N_{R}=121$.
-Figure 10: ${\rm RMSE}_{u}$ at $\sigma_{\Delta}^{2}=10^{-3}{\rm rad^{2}}$ and $\sigma_{\Delta}^{2}=10^{-4}{\rm rad^{2}}$.
-Figure 11: ${\rm MSE}_{\phi,\bm{\theta}}$ at $\sigma_{\Delta}^{2}=10^{-3}{\rm rad^{2}}$ and $\sigma_{\Delta}^{2}=10^{-4}{\rm rad^{2}}$.
-In this subsection, we first illustrate the overall positioning accuracy.
-Fig.
-6 presents the cumulative distribution functions (CDFs) of RMSE and PEB for 500 random testing positions, under different transmit powers and the number of RIS elements.
-It can be observed that increasing the RIS size and the transmit power both lead to an improvement in the overall positioning accuracy.
-Furthermore, we find that when the RIS size is set to $N_{R}=49$, the CDF of RMSE deviates significantly from the CDF of PEB, and doesn’t achieve the theoretical positioning accuracy.
-However, when the RIS size is increased to $N_{R}=121$, the CDF of the RMSE curve aligns more closely with the CDF of PEB.
-This phenomenon can also be observed in Fig.
-8.
-Next, we select a fixed testing position at coordinate $\rm[1.50m,2.15m,0.45m]^{\top}$ (rounded to two decimal places for convenience) to investigate the impact of transmit power $P$, RIS size $N_{R}$, and PN variance $\sigma_{\Delta}^{2}$ on the estimation accuracy of UE position, CFO, and PN.
-In Fig.
-7, we fix the number of RIS elements at 81 and assess the impact of random RIS phase shift versus optimized RIS configuration on positioning accuracy.
-It is evident that the actual positioning performance with random RIS phase shift does not experience a significant decline as the transmit power increases.
-At the other extreme, the optimized RIS phase shift provides an improvement of approximately two orders of magnitude in positioning accuracy.
-This fully demonstrates the crucial role of optimizing the RIS phase shift in localization problems and the effectiveness of the optimization method we employed.
-Fig.
-8 and Fig.
-9 respectively depict the curves of RMSE for positioning and MSE for estimation of $\underline{\bm{\gamma}}$ under different transmit powers, at the number of RIS sizes of 49, 81, and 121.
-It can be observed that increasing the transmit power and the number of RIS elements results in gains in all parameters to be estimated.
-This demonstrates that increasing the signal transmit power or increasing the number of RIS reflection elements is beneficial for parameter estimation.
-Fig.
-10 and Fig.
-11 respectively illustrate the RMSE for positioning and MSE for estimation of $\underline{\bm{\gamma}}$ when $N_{R}$ is set to 81, with PN variances $\sigma_{\Delta}^{2}=10^{-3}{\rm rad^{2}}$ and $\sigma_{\Delta}^{2}=10^{-4}{\rm rad^{2}}$.
-The results demonstrate that decreasing the PN covariance leads to significant improvements in combined estimation performance.
+Figs.
+8 and 9 depict the curves of the RMSE for UE localization and MSE for the estimation of $\underline{\bm{\gamma}}$, respectively, under different transmit powers, where $N_{R}$ is 49, 81, or 121.
+It is observed that increasing the transmit power and/or the RIS element number results in gains in all parameters to be estimated.
+However, as the transmit SNR increases, the RMSE and MSE do not decrease as sharply as HCRLB because of the approximate PN vector, as well as the Taylor series approximation considered when deriving closed-form expressions for the CFO and PN.
+Figure 10: ${\rm RMSE}_{u}$ at $\sigma_{\Delta}^{2}=10^{-3}{\rm\>rad^{2}}$ and $\sigma_{\Delta}^{2}=10^{-4}{\rm\>rad^{2}}$.
+Figure 11: ${\rm MSE}_{\phi,\bm{\theta}}$ at $\sigma_{\Delta}^{2}=10^{-3}{\rm\>rad^{2}}$ and $\sigma_{\Delta}^{2}=10^{-4}{\rm\>rad^{2}}$.
+Figs.
+10 and 11 plot the RMSE for localization and MSE for estimation of $\underline{\bm{\gamma}}$, respectively, when $N_{R}$ is set to 81, where the PN variance is $\sigma_{\Delta}^{2}=10^{-3}{\rm\>rad^{2}}$ or $\sigma_{\Delta}^{2}=10^{-4}{\rm\>rad^{2}}$.
+The results demonstrate that decreasing the PN covariance leads to significant improvements in the estimation performance of the CFO, PN, and UE position.
+Another important finding is that a smaller PN variance would make the RMSE and MSE curves more closely aligned with the HCRLB curves because of the more accurately approximated parameters when updating the CFO and PN parameters.
+Figure 12: Benchmarking the performance of the proposed algorithms against the state-of-the-art approach in [27].
+Last but not least, we compare the proposed overall algorithm of the CFO, PN, and UE position estimation and the RIS configuration, with the benchmark algorithm developed in [27].
+It is shown in Fig.
+12 that, after optimization, the theoretical localization accuracy of our proposed algorithm outperforms the benchmark, which confirms the conclusion drawn in Fig.
+4(b).
+As for the localization RMSE, the benchmark shows a large gap with its corresponding PEB, and its performance shows no demonstrable correlation with the transmit power because the estimation algorithm totally overlooked the CFO and PN.
+In comparison, the proposed algorithm significantly improves the localization accuracy by about one order of magnitude under relatively low transmit powers of below $-10{\rm\>dBm}$, and by two orders of magnitude under high transmit powers of over $10{\rm\>dBm}$.
 VII Conclusion
-This paper investigates the optimization of RIS phase shift and the joint estimation of CFO, PN, and UE position under imperfect channel conditions for an OFDM system in a near-field static localization scenario.
-An iterative joint estimation algorithm based on the MAP criterion and GD algorithm is proposed.
-Extensive simulation results have demonstrated that by optimizing the RIS phase shift, the theoretical positioning accuracy has been improved by two orders of magnitude.
-Moreover, the maximum positioning accuracy can be lower than $\rm 10^{-2}m$.
-These findings strongly support the efficacy of the proposed estimation and optimization algorithms in the paper.
-In order to further improve the estimation performance and support a larger AOI, the optimization algorithm used in this paper requires averaging a large number of PEBs related to random positions and will result in extremely high computation complexity.
+This paper investigates the optimization of RIS phase shifts and the joint estimation of CFO, PN, and UE position under imperfect hardware conditions for an OFDM system in a near-field static localization scenario.
+An AO-based, iterative joint estimation algorithm based on the MAP criterion and GD algorithm has been proposed.
+Extensive simulations have demonstrated that by optimizing the RIS phase shifts, the theoretical localization accuracy has been improved by two orders of magnitude.
+Moreover, the achievable localization accuracy can be lower than $\rm 10^{-2}\>m$.
+These findings corroborate the effectiveness of the proposed algorithms.
+To further improve the estimation performance and support a larger AOI, the optimization algorithm used in this paper requires averaging a large number of PEBs related to random positions with high computational complexity.
 Hence, it becomes imperative to utilize an optimization algorithm with reduced complexity in this scenario.
-This would be a fruitful area for further work.
+This will be our future work.
 Appendix A Derivation of partial derivatives of $\bm{\mu}[n]$ and $\bm{\mu}[n]^{*}$
-The partial derivatives of $\bm{\mu}[n]$ with respect to all parameters to be estimated are derived as follows
+The partial derivatives of $\bm{\mu}[n]$ with regard to all parameters to be estimated are derived as follows
 $\frac{\partial\bm{\mu}[n]}{\partial\phi}$ $=\sqrt{P}j\frac{2\pi n}{N}e^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}\bm{\bar{w}}^{\rm H}\mathbf{G}[\mathbf{S}]_{:,n},$ (55a)
@@
 $\frac{\partial\bm{\mu}[n]}{\partial\phi_{el}}$ $=\sqrt{P}e^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}\bm{\bar{w}}^{\rm H}\frac{\partial\mathbf{G}}{\partial\phi_{el}}[\mathbf{S}]_{:,n},$ (55f)
-and the partial derivatives of $\bm{\mu}[n]^{*}$ with respect to all parameters to be estimated are derived as follows
+and the partial derivatives of $\bm{\mu}[n]^{*}$ with regard to all parameters to be estimated are derived as follows
 $\frac{\partial\bm{\mu}[n]^{*}}{\partial\phi}$ $=-j\sqrt{P}\frac{2\pi n}{N}e^{-j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}]_{:,n}^{\rm H}\mathbf{G}^{\rm H}\bm{\bar{w}},$ (56a)

### v4

@@
 However, most studies have overlooked the impact of carrier frequency offset (CFO) and phase noise (PN) resulting from hardware impairments on localization.
-This paper presents a novel, alternating optimization (AO)-based algorithm to jointly estimate the CFO, PN, and user position, where, provided the user position, closed-form expressions for the CFO and PN are derived per iteration, significantly reducing the complexity and enhancing the stability of the algorithm.
+This paper presents a novel, alternating optimization (AO)-based algorithm to jointly estimate the CFO, PN, and user equipment (UE) position in orthogonal frequency division multiplexing (OFDM) systems, where, provided the UE position, closed-form expressions for the CFO and PN are derived per iteration, significantly reducing the complexity and enhancing the stability of the algorithm.
 Another important aspect is a new RIS phase shift optimization algorithm developed to minimize the analytical lower bound of localization accuracy, hence benefiting localization.
@@
 I-A Motivation and Challenges
-With the development of industrial Internet of Things, there is an increasing demand for centimeter-level localization accuracy in the sixth-generation (6G) mobile networks [1, 2, 3].
-While global navigation satellite systems (GNSSs) have been widely applied to outdoor scenarios [4], they undergo insufficient localization accuracy and signal instability arising from obstacles [5, 6].
-In recent years, communication systems have been considered to assist in localization, e.g., integrated sensing and communication (ISAC), aiming to achieve more accurate and continuous localization alongside GNSS signals.
-Although communication-assisted localization systems have been extensively studied, few studies have considered the adverse effects of hardware impairments, such as carrier frequency offset (CFO) and phase noise (PN), on localization accuracy.
+With the development of the industrial Internet of Things, there is an increasing demand for centimeter-level localization accuracy in the sixth-generation (6G) mobile networks [1].
+While global navigation satellite systems (GNSSs) have been widely applied to outdoor scenarios, they undergo insufficient localization accuracy and signal instability arising from obstacles [2, 3].
+Targeting at areas with little access to satellite signals, wireless localization [4], also known as communication-assisted localization [5], has been widely studied.
+It can be potentially an integral part of integrated communication and sensing (ISAC) systems [6, 7] by utilizing communication devices and signals to achieve localization functionality.
+While wireless localization systems have been extensively studied, few studies have considered the adverse effects of hardware impairments, such as carrier frequency offset (CFO) and phase noise (PN), on localization accuracy.
 CFO is a constant parameter due to the frequency offset between the transmitter and the receiver.
@@
 Another challenge arises from the increasingly likely blockages of signals, e.g., millimeter wave (mmWave) signals with the quasi-optical property, in many dense urban areas [9].
-Reconfigurable intelligent surface (RIS) [9, 10] has been considered in communication systems to deal with the problem of signal blockage.
+Reconfigurable intelligent surface (RIS) has been considered in communication systems to deal with the problem of signal blockage [10, 11].
 Localization can be implemented with the help of virtual line-of-sight (LoS) paths provided by RISs even when the LoS paths are blocked, and can be improved via passive beamforming.
-In this sense, RIS-assisted user localization is of practical interest, and could be particularly challenging with the further consideration of the CFO and PN.
+In this sense, RIS-assisted user equipment (UE) localization is of practical interest, and could be particularly challenging with the further consideration of the CFO and PN.
 However, this has not been studied in the existing literature.
 I-B Related works
-Communication systems have been widely employed to assist in high-precision user localization, typically under a far-field setting; i.e., the user and the anchors are in the far field of each other.
+I-B1 RIS-free localization with ideal hardware
+Communication systems have been widely employed to assist in high-precision UE localization, typically under a far-field setting; i.e., the UE and the anchors are in the far field of each other.
 For example, Shahmansoori et al.
@@
 Hu et al.
-[12] presented a two-dimensional (2D) localization algorithm based on the celebrated estimating signal parameters via rotational invariance technique (ESPRIT), which estimates user positions by constructing signal subspaces.
+[13] presented a two-dimensional (2D) localization algorithm based on the celebrated estimating signal parameters via rotational invariance technique (ESPRIT), which estimates UE positions by constructing signal subspaces.
+I-B2 RIS-free localization under CFO and PN
 Some studies [14, 15] have explored the impact of realistic hardware impairments, including the CFO and PN, with no RIS considered.
@@
 [14] simulated the effects of CFO and PN on angle-of-arrival (AOA) estimation.
+Nevertheless, CFO and PN were ignored when designing the algorithms and were only included in simulations, showing that the CFO and PN can significantly and adversely affect localization and tracking performance.
 Chen et al.
-[14] considered the impacts of CFO, PN, mutual coupling, and power amplifier nonlinearity on 6G localization, and performed analysis using the derived misspecified Cramér-Rao lower bound (CRLB).
-However, neither of them provided an estimation algorithm to handle the hardware impairments.
-Some other works [15, 16, 17, 18] have delved into localization algorithm design for scenarios with either CFO or PN.
-On the one hand, the CFO was studied in [15, 16].
-Qiu et al.
-[15] designed an maximum likelihood estimation (MLE)-based weighted iterative algorithm for the joint estimation of CFO and AOA, yet the coupling of CFO and position parameters in each iteration incurs high complexity.
+[15] considered the impacts of CFO, PN, mutual coupling, and power amplifier nonlinearity on 6G localization, and derived the misspecified Cramér-Rao lower bound (CRLB).
+It was analytically concluded that the PN affects both angle and delay estimation, whereas the CFO has a more significant effect on angle estimation.
+However, neither of these works, i.e., [14, 15] have provided an estimation algorithm to handle hardware impairments.
+Some other works [16, 17, 18, 19] have delved into localization algorithm design with either CFO or PN.
+On the one hand, Qiu et al.
+[16] designed a maximum likelihood estimation (MLE)-based weighted iterative algorithm for the joint estimation of CFO and AOA.
 Zakariyya et al.
-[16] proposed a high-resolution multiple signal classification (MUSIC) algorithm for localization, but did not utilize the CFO estimates in localization, leading to a loss in localization accuracy.
-On the other hand, the PN was considered in [17, 18].
-Xiao et al.
-[17] combined the expectation maximization (EM) and MUSIC algorithms for joint estimation of AOA and PN, but only considered a single-carrier system under a 2D setting.
+[17] proposed a high-resolution enhancement of the multiple signal classification (MUSIC) algorithm for localization.
+While the algorithm balanced spectrum resolution and accuracy, the CFO was not estimated in localization, leading to a loss in localization accuracy.
+On the other hand, Xiao et al.
+[18] combined the expectation maximization (EM) and MUSIC algorithms for joint estimation of AOA and PN.
+At a typical signal-to-noise ratio (SNR) level, this algorithm can reduce the estimation error by over $55\%$ compared to its benchmarks.
+However, only a simple single-carrier system under a 2D far-field setting was considered.
 Zheng et al.
-[18] formulated a maximum a-posteriori (MAP) problem iteratively solved using the block majorization minimization (MM) algorithm after approximation, where the PN was simplified as a Gaussian variable.
-However, the above methods have not considered the use of RISs nor addressed the signal blockage by obstacles.
+[19] formulated a maximum a-posteriori (MAP) problem solved iteratively using the block majorization minimization (MM) algorithm after approximation.
+Under a medium or small PN, the algorithm improves localization and channel estimation.
+However, none of these methods [16, 17, 18, 19] have considered the use of RISs or addressed the signal blockage by obstacles.
+I-B3 RIS-assisted localization with ideal hardware
 RIS has been integrated into localization systems to address the problem of signal blockage, but little consideration has been given to the above-mentioned hardware impairments, i.e., CFO and PN.
-From the perspective of user mobility, existing studies can be divided between static localization [19, 20, 21] and dynamic localization [22, 23].
-The former encompasses scenarios where the users undergo minimal movement within a coherent time.
-The latter involves scenarios when the users move at a relatively high speed, and the Doppler shift cannot be overlooked.
 From the perspective of localization distance, existing research can be divided between far-field localization [20, 21, 22] and near-field localization [23, 20, 21, 24].
-In the far-field case, the user-RIS distance exceeds Fraunhofer distance [26], and the signal can be approximated as a planar wave.
+In the far-field case, the UE-RIS distance exceeds Fraunhofer distance [25], and the signal can be approximated as a planar wave.
 Conversely, in the near-field case, the signal cannot be treated as parallel propagation and should be considered as a spherical wave.
 For example, Luan et al.
-[27] focused on the near-field static 2D localization scenario, and proposed a robust optimization algorithm to configure the RIS phase shifts for improving localization accuracy.
+[26] focused on the near-field 2D localization scenario, and proposed a robust optimization algorithm to configure the RIS phase shifts for improving localization accuracy.
 The transmit powers were also optimized in the considered localization system.
 Ozturk et al.
-[28] further introduced a more practical RIS model with phase-dependent amplitude variations in the near-field circumstance.
-The misspecified CRLB was derived and an approximated mismatched maximum likelihood-based algorithm was proposed for joint UE localization and RIS parameter estimation.
-The system model in this paper also belongs to the near-field static category.
-Although we adopt the classical ideal RIS phase shift model with unit amplitude for simplicity, we first take the CFO and PN into consideration.
-Elzanaty et al.
-[23] studied RIS-assisted localization for multi-antenna mobile users.
-They proposed a three-dimensional (3D) model that applies to both far-field and near-field scenarios, and analyzed the position and orientation error bounds.
+[27] introduced a more practical RIS model with phase-dependent amplitude variations in the near-field circumstance.
+The mis-specified CRLB was derived and an approximated mismatched maximum likelihood-based algorithm was proposed for joint UE localization and RIS parameter estimation.
 Lin et al.
@@
 They established nonlinear equations exploiting the characteristics of the twin-RIS structure to recover the multipath parameters.
-User localization for static users in the far-field scenario was then performed based on the estimated channel parameters.
-Teng et al.
-[22] introduced a new message-passing algorithm called BULT to locate mobile users in the far-field scenario.
-The Bayesian CRLB was derived for BULT.
-The existing RIS-assisted localization studies have assumed ideal channel conditions without considering realistic hardware impairments.
+Localization for UEs in the far-field scenario was then performed based on the estimated channel parameters.
+I-B4 RIS-assisted localization under CFO and PN
+Most existing RIS-assisted localization studies have assumed ideal channel conditions without considering realistic hardware impairments.
 Only several studies have considered modeling some specific impairments.
@@
 However, the algorithm is only suitable for far-field scenarios.
-To the best of our knowledge, no existing research has specifically analyzed the impact of both CFO and PN on RIS-assisted near-field localization systems.
+Unlike these existing works, we not only analytically quantify the impact of CFO and PN on localization accuracy, but also design efficient localization and RIS configuration algorithms.
+As a result, the localization accuracy can be improved by two orders of magnitude, compared to the existing methods.
+To the best of our knowledge, no existing research has specifically analyzed the impact of both CFO and PN on RIS-assisted near-field localization systems nor developed corresponding localization and RIS configuration algorithms.
 I-C Contributions
-In this paper, we investigate a RIS-assisted OFDM system, where practical CFO and PN convolve with the data symbols, causing signal constellation rotation and ICI and, consequently, hindering user localization in near-field environments.
+In this paper, we investigate a RIS-assisted OFDM system, where practical CFO and PN convolve with the data symbols, causing signal constellation rotation and ICI and, consequently, hindering UE localization in near-field environments.
 The key contributions of this paper are outlined as follows:
-  • We investigate a problem of static near-field localization in a RIS-aided OFDM system, while taking into account the CFO and PN.
+  • We investigate a problem of near-field localization in a RIS-aided OFDM system, while taking into account the CFO and PN.
 We propose an alternating optimization (AO)-based iterative estimation approach, where the MAP criterion and gradient descent (GD) algorithm are employed with closed-form expressions derived for the CFO and PN estimation.
@@
 Compared to the state-of-the-art benchmarks, our algorithms can improve the localization accuracy by about one order of magnitude under relatively low transmit power and even by two orders of magnitude under high transmit power.
-Figure 1: The considered RIS-aided 3D near-field localization system, including an AN, a UE, and an $N_{R}$-element RIS.
+Figure 1: The considered RIS-aided 3D near-field localization system, including an AN, a UE, and an ${\color[rgb]{0,0,0}N_{\text{R}}}$-element RIS.
 The LoS link between the AN and UE is obstructed.
@@
 Notation Definition
-$N_{R}$ Number of RIS elements
+${\color[rgb]{0,0,0}N_{\text{R}}}$ Number of RIS elements
 $N$ Number of subcarriers
@@
 $D$ Aperture of the RIS
-$f_{c}$ Carrier frequency of the system
+${\color[rgb]{0,0,0}f_{\text{c}}}$ Carrier frequency of the system
 $B$ Bandwidth of the system
@@
 $\phi$ Normalized CFO
-$U$ Position sampling amount
+$\color[rgb]{0,0,0}S$ Position sampling amount
 $\sigma^{2}$ Variance of the noise at the UE
 $\sigma_{\Delta}^{2}$ Innovation variance for PN
-$\bm{p}_{a}$ Coordinate of the AN
-$\bm{p}_{r}$ Coordinate of the reference point of the RIS
-$\bm{p}_{u}$ Coordinate of the UE position
+${\color[rgb]{0,0,0}\bm{p}_{\text{a}}}$ Coordinate of the AN
+$\bm{p}_{r}$ Coordinate of the $r$-th element of the RIS
+$\bm{p}_{\color[rgb]{0,0,0}\text{u}}$ Coordinate of the UE position
 $\bm{\theta}$ PN vector
@@
 II System Model
-The considered RIS-assisted OFDM downlink wireless localization system is composed of an anchor node (AN), a user equipment (UE), and an $N_{R}$-element RIS, as illustrated in Fig.
+The considered RIS-assisted OFDM downlink wireless localization system is composed of an anchor node (AN), a UE, and an $N_{\text{R}}$-element RIS, as illustrated in Fig.
 1.
 We make the assumption that the LoS path between the AN and UE is obstructed, and the channel state information (CSI) remains unchanged during localization, as in [20, 26, 29].
-The positions of the AN and each RIS element, denoted as $\bm{p}_{a}$ and $\bm{p}_{r}=[x_{r},y_{r},z_{r}]^{\top}$, respectively, are known, where $r$ is the index of a RIS element.
-The position of UE, denoted as $\bm{p}_{u}=[x_{u},y_{u},z_{u}]^{\top}$, is to be estimated.
+Since both the AN and RIS are part of the network infrastructure, the positions of the AN and each RIS element, denoted as ${\color[rgb]{0,0,0}\bm{p}_{\text{a}}}$ and $\bm{p}_{r}=[x_{r},y_{r},z_{r}]^{\top}$, respectively, are known, as typically considered in RIS-assisted localization studies, e.g., [21, 26, 29, 31].
+Here, $r$ is the index of a RIS element.
+The position of UE, denoted as $\bm{p}_{\color[rgb]{0,0,0}\text{u}}=[x_{\color[rgb]{0,0,0}\text{u}},y_{\color[rgb]{0,0,0}\text{u}},z_{\color[rgb]{0,0,0}\text{u}}]^{\top}$, is to be estimated.
 The RIS elements are organized in a uniform planar array (UPA).
@@
 The RIS is placed on the $yz$-plane, and both edges on the coordinate axes are located in the positive half-axis.
-Hence, the coordinates of the $r$-th RIS element can be represented as $\bm{p}_{r}=[0,d\times((r-1)\mod N_{R}),d\times\lceil\frac{r}{\sqrt{N_{R}}}-1\rceil]^{\top}$, where $d$ stands for the element spacing of the RIS, and is typically half of a wavelength.
-To satisfy the conditions of the near-field setup, both the AN-RIS distance, as well as the RIS-UE distance, should lie within the Fresnel region [26], i.e., the interval $[0.62\sqrt{D^{3}/\lambda},2D^{2}/\lambda]$, where $D=\sqrt{2}d(\sqrt{N_{R}}-1)$ represents the aperture of the RIS and $\lambda$ denotes the wavelength.
-Additionally, we adopt a polar coordinate system for estimating the user’s position, i.e., using $d_{Ru}$, $\phi_{az}$ and $\phi_{el}$, to express the user’s position, indicating the distance, azimuth angle, and elevation angle of UE relative to the coordinate origin, respectively.
-The relationships between these three parameters and $x_{u}$, $y_{u}$ and $z_{u}$ are given by
-$d_{Ru}$ $=\sqrt{x_{u}^{2}+y_{u}^{2}+z_{u}^{2}};$ (1a)
-$\phi_{az}$ $=\arctan\frac{y_{u}}{x_{u}};$ (1b)
-$\phi_{el}$ $=\arccos\frac{z_{u}}{\sqrt{x_{u}^{2}+y_{u}^{2}+z_{u}^{2}}},$ (1c)
-and, in turn, $x_{u}$, $y_{u}$ and $z_{u}$ can be written as
-$x_{u}$ $=d_{Ru}\sin\phi_{el}\cos\phi_{az};$ (2a)
-$y_{u}$ $=d_{Ru}\sin\phi_{el}\sin\phi_{az};$ (2b)
-$z_{u}$ $=d_{Ru}\cos\phi_{el}.$ (2c)
-An OFDM symbol with a center carrier frequency of $f_{c}$, a signal bandwidth of $B$ and $N$ subcarriers is considered in this paper.
-Upon sampling at a rate of $T_{s}=\frac{1}{NB}$ and removing the cyclic prefix (CP) at the UE, the received time-domain signal $\bm{y}$ can be written as
-$\bm{y}=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\bm{h}_{LoS}+\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\bm{h}_{NLoS}+\tilde{\bm{v}},$ (3)
+Hence, the coordinates of the $r$-th RIS element can be represented as $\bm{p}_{r}=[0,d\times(r\mod\sqrt{N_{\text{R}}}),d\times\lceil\frac{r+1}{\sqrt{N_{\text{R}}}}-1\rceil]^{\top}$, where $d$ stands for the element spacing of the RIS.
+We set $d$ to half-wavelength [23], [24], [33], because signals at different RIS elements may couple if $d$ is too small, or the RIS array’s sidelobe can be significant, leading to poor directionality for the RIS if $d$ is too large.
+To satisfy the conditions of the near-field setup, both the AN-RIS distance, as well as the RIS-UE distance, should lie within the Fresnel region [25], i.e., the interval $[0.62\sqrt{D^{3}/\lambda},2D^{2}/\lambda]$, where $D=\sqrt{2}d(\sqrt{{\color[rgb]{0,0,0}N_{\text{R}}}}-1)$ represents the aperture of the RIS and $\lambda$ denotes the wavelength.
+Additionally, we adopt a polar coordinate system for estimating the UE’s position, i.e., using ${\color[rgb]{0,0,0}d_{\text{R}}}$, ${\color[rgb]{0,0,0}\phi_{\text{az}}}$ and ${\color[rgb]{0,0,0}\phi_{\text{el}}}$, to express the UE’s position, indicating the distance, azimuth angle, and elevation angle of UE relative to the coordinate origin, respectively.
+The relationships between these three parameters and $x_{\color[rgb]{0,0,0}\text{u}}$, $y_{\color[rgb]{0,0,0}\text{u}}$ and $z_{\color[rgb]{0,0,0}\text{u}}$ are given by
+${\color[rgb]{0,0,0}d_{\text{R}}}$ $=\sqrt{x_{\color[rgb]{0,0,0}\text{u}}^{2}+y_{\color[rgb]{0,0,0}\text{u}}^{2}+z_{\color[rgb]{0,0,0}\text{u}}^{2}};$ (1a)
+${\color[rgb]{0,0,0}\phi_{\text{az}}}$ $=\arctan\frac{y_{\color[rgb]{0,0,0}\text{u}}}{x_{\color[rgb]{0,0,0}\text{u}}};$ (1b)
+${\color[rgb]{0,0,0}\phi_{\text{el}}}$ $=\arccos\frac{z_{\color[rgb]{0,0,0}\text{u}}}{\sqrt{x_{\color[rgb]{0,0,0}\text{u}}^{2}+y_{\color[rgb]{0,0,0}\text{u}}^{2}+z_{\color[rgb]{0,0,0}\text{u}}^{2}}},$ (1c)
+and, in turn, $x_{\color[rgb]{0,0,0}\text{u}}$, $y_{\color[rgb]{0,0,0}\text{u}}$ and $z_{\color[rgb]{0,0,0}\text{u}}$ can be written as
+$x_{\color[rgb]{0,0,0}\text{u}}$ $={\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\cos{\color[rgb]{0,0,0}\phi_{\text{az}}};$ (2a)
+$y_{\color[rgb]{0,0,0}\text{u}}$ $={\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\sin{\color[rgb]{0,0,0}\phi_{\text{az}}};$ (2b)
+$z_{\color[rgb]{0,0,0}\text{u}}$ $={\color[rgb]{0,0,0}d_{\text{R}}}\cos{\color[rgb]{0,0,0}\phi_{\text{el}}}.$ (2c)
+An OFDM symbol with a center carrier frequency of ${\color[rgb]{0,0,0}f_{\text{c}}}$, a signal bandwidth of $B$ and $N$ subcarriers is considered in this paper.
+Upon sampling at a rate of ${\color[rgb]{0,0,0}T_{\text{s}}}=\frac{1}{NB}$ and removing the cyclic prefix (CP) at the UE, the received time-domain signal $\bm{y}$ can be written as
+$\bm{y}=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}{\color[rgb]{0,0,0}\bm{h}_{\text{LoS}}}+\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}{\color[rgb]{0,0,0}\bm{h}_{\text{NLoS}}}+\tilde{\bm{v}},$ (3)
 where
@@
   • $\mathbf{S}\triangleq\sqrt{N}\mathbf{F}^{\rm H}{\rm Diag}(\bm{s})$, where $\mathbf{F}^{\rm H}$ is the inverse fast Fourier transform (FFT) matrix and $\bm{s}$ is the frequency-domain transmitted symbol sequence at the AN;
-  • $\bm{h}_{LoS}$ and $\bm{h}_{NLoS}$ represent the LoS and non-LoS parts of the frequency-domain channel, respectively;
-  • $\tilde{\bm{v}}$ is the additive Gaussian noise at the UE, and follows the distribution $\tilde{\bm{v}}\sim\mathcal{CN}(\bm{0}_{N\times 1},\sigma_{AWGN}^{2}\mathbf{I}_{N})$.
-Assume that $\bm{h}_{NLoS}$ follows a known Gaussian distribution $\bm{h}_{NLoS}\sim\mathcal{CN}(0,\sigma_{NLoS}^{2}\mathbf{I}_{N})$.
-Thus, $\bm{\mu}_{NLoS}\triangleq\mathbf{X}\bm{h}_{NLoS}$ also follows a known Gaussian distribution $\bm{\mu}_{NLoS}\sim\mathcal{CN}(0,\sigma_{NLoS}^{2}\mathbf{X}\mathbf{X}^{\rm H})$ with $\mathbf{X}\triangleq\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}$.
-Combine $\bm{\mu}_{NLoS}$ with the additive Gaussian noise $\tilde{\bm{v}}$.
-For conciseness, denote $\bm{h}_{LoS}$ with $\bm{h}$.
+  • ${\color[rgb]{0,0,0}\bm{h}_{\text{LoS}}}$ and ${\color[rgb]{0,0,0}\bm{h}_{\text{NLoS}}}$ represent the LoS and non-LoS parts of the frequency-domain channel, respectively [29, 34];
+  • $\tilde{\bm{v}}$ is the additive Gaussian noise at the UE, and follows the distribution $\tilde{\bm{v}}\sim\mathcal{CN}(\bm{0}_{N\times 1},{\color[rgb]{0,0,0}\sigma_{\text{AWGN}}^{2}}\mathbf{I}_{N})$.
+Assume that ${\color[rgb]{0,0,0}\bm{h}_{\text{NLoS}}}$ follows a known Gaussian distribution [35], i.e., ${\color[rgb]{0,0,0}\bm{h}_{\text{NLoS}}}\sim\mathcal{CN}({\color[rgb]{0,0,0}\bm{0}_{N\times 1}},{\color[rgb]{0,0,0}\sigma_{\text{NLoS}}^{2}}\mathbf{I}_{N})$.
+Thus, ${\color[rgb]{0,0,0}\bm{\mu}_{\text{NLoS}}}\triangleq\mathbf{X}{\color[rgb]{0,0,0}\bm{h}_{\text{NLoS}}}$ also follows a known Gaussian distribution ${\color[rgb]{0,0,0}\bm{\mu}_{\text{NLoS}}}\sim\mathcal{CN}({\color[rgb]{0,0,0}\bm{0}_{N\times 1}},{\color[rgb]{0,0,0}\sigma_{\text{NLoS}}^{2}}\mathbf{X}\mathbf{X}^{\rm H})$ with $\mathbf{X}\triangleq\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}$ and $\mathbf{X}\mathbf{X}^{\rm H}=NP\mathbf{I}_{N}$.
+Combine ${\color[rgb]{0,0,0}\bm{\mu}_{\text{NLoS}}}$ with the additive Gaussian noise $\tilde{\bm{v}}$.
+For conciseness, we rewrite ${\color[rgb]{0,0,0}\bm{h}_{\text{LoS}}}$ with $\bm{h}$ in the rest of this paper.
 The received signal is rewritten as
 $\bm{y}=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\bm{h}+\bm{v},$ (6)
-where $\bm{v}\triangleq\bm{\mu}_{NLoS}+\tilde{\bm{v}}$ is Gaussian distributed, i.e, $\bm{v}\sim\mathcal{CN}(0,\sigma^{2}\mathbf{I}_{N})$, with $\sigma^{2}=\sigma_{AWGN}^{2}+\sigma_{NLoS}^{2}$.
-Each element of $\bm{h}$ can be written as $\bm{h}[n]=\sum_{r=1}^{N_{R}}\sqrt{\rho_{aR,n,r}}\sqrt{\rho_{Ru,n,r}}\bm{w}[r]e^{-j2\pi\frac{nB}{N}(\tau_{ar}+\tau_{ru})}$, where $\sqrt{\rho_{aR,n,r}}=\frac{\sqrt{G_{a}}\lambda_{n}}{4\pi d_{ar}}$ and $\sqrt{\rho_{Ru,n,r}}=\frac{\sqrt{G_{t}}\lambda_{n}}{4\pi d_{ru}}$ are the free-space path losses from the AN to the $r$-th RIS element, and from the $r$-th RIS element to the UE, respectively.
-$\bm{w}[r]$ is the phase shift of the $r$-th RIS element, $\tau_{ar}=\frac{d_{ar}}{c}=\frac{\|\bm{p}_{r}-\bm{p}_{a}\|}{c}$ is the delay from the AN to the $r$-th RIS element, while $\tau_{ru}=\frac{d_{ru}}{c}=\frac{\|\bm{p}_{u}-\bm{p}_{r}\|}{c}$ denotes the delay from the $r$-th RIS element to the UE, with the light speed $c$.
+where $\bm{v}\triangleq{\color[rgb]{0,0,0}\bm{\mu}_{\text{NLoS}}}+\tilde{\bm{v}}$ yields a Gaussian distribution, i.e, $\bm{v}\sim\mathcal{CN}({\color[rgb]{0,0,0}\bm{0}_{N\times 1}},\sigma^{2}\mathbf{I}_{N})$, with $\sigma^{2}={\color[rgb]{0,0,0}\sigma_{\text{AWGN}}^{2}}+{\color[rgb]{0,0,0}\sigma_{\text{NLoS}}^{2}NP}$.
+Each element of $\bm{h}$ can be written as [26]
+$\bm{h}[n]=\sum_{r=0}^{N_{\text{R}}-1}\sqrt{\rho^{n,r}_{\text{aR}}}\sqrt{\rho_{{\text{Ru}}}^{n,r}}\bm{w}[r]e^{-j2\pi\frac{nB}{N}(\tau_{\text{a}r}+{\color[rgb]{0,0,0}\tau_{r\text{u}}})},$ (7)
+where ${\color[rgb]{0,0,0}\sqrt{\rho_{\text{aR}}^{n,r}}}=\frac{\sqrt{G_{\color[rgb]{0,0,0}\text{a}}}\lambda_{n}}{4\pi d_{{\color[rgb]{0,0,0}\text{a}}r}}$ and ${\color[rgb]{0,0,0}\sqrt{\rho_{\text{Ru}}^{n,r}}}=\frac{\sqrt{G_{\color[rgb]{0,0,0}\text{t}}}\lambda_{n}}{4\pi{\color[rgb]{0,0,0}d_{r\text{u}}}}$ are the free-space path losses from the AN to the $r$-th RIS element, and from the $r$-th RIS element to the UE, respectively; $\bm{w}[r]$ is the phase shift of the $r$-th RIS element; $\tau_{{\color[rgb]{0,0,0}\text{a}}r}=\frac{d_{{\color[rgb]{0,0,0}\text{a}}r}}{c}=\frac{\|\bm{p}_{r}-{\color[rgb]{0,0,0}\bm{p}_{\text{a}}}\|}{c}$ is the delay from the AN to the $r$-th RIS element, while ${\color[rgb]{0,0,0}\tau_{r\text{u}}}=\frac{{\color[rgb]{0,0,0}d_{r\text{u}}}}{c}=\frac{\|\bm{p}_{\color[rgb]{0,0,0}\text{u}}-\bm{p}_{r}\|}{c}$ is the delay from the $r$-th RIS element to the UE; $c$ is the speed of light.
 III Proposed Joint CFO, PN and UE Position Estimation Algorithm
 In previous works [26, 29], localization was based on frequency-domain signals, and different subcarriers’ signals were often assumed to be independent.
-However, the CFO and PN would result in ICI, which increases the difficulty in the estimation of the CFO, PN and the position of the UE.
+However, the CFO and PN can result in ICI, which increases the difficulty in the estimation of the CFO, PN, and UE position.
 The estimation can only be conducted in the time domain.
 In this section, we propose an AO-based iterative algorithm, which employs the MAP criterion and GD algorithm to jointly estimate the normalized CFO, PN, and the UE position.
-Define the parameter vector $\bm{\lambda}\triangleq\left[\phi,\bm{\theta}^{\top},d_{Ru},\phi_{az},\phi_{el}\right]^{\top}$, wherein $\bm{\theta}$ is a random variable and the others are deterministic.
+To address the new challenge of estimating the CFO and PN, the PN vector is approximated by exploiting the redundancy of its covariance matrix, and the posterior functions concerning the CFO and PN are linearized.
+Closed-form estimates are derived for both CFO and PN.
+The proposed localization algorithm can serve as a prerequisite for tracking.
+Since the Doppler shift can be interpreted as part of CFO, by estimating the CFO, we can provide an estimate of the UE’s speed and improve tracking.
+For example, the UE position and CFO estimates provided by our proposed localization algorithm can be used as initialization inputs for existing tracking algorithms, e.g., [29].
+Define the parameter vector $\bm{\lambda}\triangleq\left[\phi,\bm{\theta}^{\top},{\color[rgb]{0,0,0}d_{\text{R}}},{\color[rgb]{0,0,0}\phi_{\text{az}}},{\color[rgb]{0,0,0}\phi_{\text{el}}}\right]^{\top}$, wherein $\bm{\theta}$ is a random variable and the others are deterministic.
+The $k$-th variable in $\bm{\lambda}$ is denoted as $\lambda_{k}$.
 The posterior distribution of $\bm{\lambda}$, given the received signal $\bm{y}$, can be expressed as
@@
 According to the MAP criterion, $\bm{\lambda}$ can be estimated by optimizing the following unconstrained function:
-$\hat{\bm{\lambda}}\propto\arg\min_{\bm{\lambda}}\mathcal{L}(\bm{\lambda}),$ (8)
-where $\mathcal{L}(\bm{\lambda})=-\log p(\bm{y}\mid\bm{\lambda})-\log p(\bm{\theta})$, which is obtained by calculating the negative logarithmic function of (7) and then suppressing constants.
+$\hat{\bm{\lambda}}\propto\arg\min_{\bm{\lambda}}\>-\log p(\bm{y}\mid\bm{\lambda})-\log p(\bm{\theta}),$ (9)
+which is obtained by calculating the negative logarithmic function of (8) and then suppressing constants.
 III-A Phase Noise Estimation
@@
 To lower the computational complexity of the estimation, we exploit the redundancy of the covariance matrix of the PN vector to approximate it and reduce variables requiring estimation.
+This approximation is based on the commonly used truncated singular value decomposition (SVD) method [36, 37].
 As illustrated in Fig. 2, the eigenvalues of $\bm{\Psi}$ are mostly close to zero, indicating significant redundancy in $\bm{\Psi}$. With this property, we can approximate the PN vector as
@@
 where $\bm{\eta}\sim\mathcal{N}(\bm{0}_{L\times 1},\mathbf{I}_{L})$ contains $L\leq N$ PN variables, and $\bm{\Pi}\in\mathbb{R}^{N\times L}$ is the transition matrix.
-We set $L=\frac{N}{2}$ in this paper.
-To derive $\bm{\Pi}$, we first perform eigenvalue decomposition (EVD) on $\bm{\Psi}$, i.e., $\bm{\Psi}=\mathbf{U}\mathbf{D}\mathbf{U}^{\top}$, where $\mathbf{D}={\rm Diag}(\bm{e})$ and $\bm{e}$ collects the eigenvalues of $\bm{\Psi}$ arranged in descending order.
-Subsequently, $\bm{\Pi}$ can be calculated by $\bm{\Pi}=[\mathbf{U}]_{:,0:L-1}\times\sqrt{\bm{e}}$.
+To derive $\bm{\Pi}$, we first perform eigenvalue decomposition (EVD) on $\bm{\Psi}$.
+For the real symmetric matrix $\mathbf{\Psi}$, the EVD is equivalent to its SVD.
+Based on the EVD, we have $\bm{\Psi}=\mathbf{U}\mathbf{D}\mathbf{U}^{\top}$, where $\mathbf{D}={\rm Diag}(\bm{e})$ and $\bm{e}$ collects the eigenvalues of $\bm{\Psi}$ arranged in descending order.
+Subsequently, $\bm{\Pi}$ can be calculated by $\bm{\Pi}=[\mathbf{U}]_{:,0:L-1}\times{\rm Diag}\left\{\tilde{\bm{e}}\right\}$, with $\tilde{\bm{e}}\triangleq\left[\sqrt{\bm{e}[0]},\sqrt{\bm{e}[1]},\cdots,\sqrt{\bm{e}[L-1]}\right]^{\top}$.
+Define $\tilde{\mathbf{\Psi}}=\mathbf{\Pi}\mathbf{\Pi}^{\top}$. Thus, $\mathbf{\Pi}\bm{\eta}\sim\mathcal{N}(\bm{0}_{N\times 1},\tilde{\mathbf{\Psi}})$. $\tilde{\mathbf{\Psi}}$ can be rewritten as
+$\tilde{\mathbf{\Psi}}$ $=\left[\mathbf{U}\right]_{:,0:L-1}{\rm Diag}\left\{\tilde{\bm{e}}\right\}\left\{\tilde{\bm{e}}\right\}^{\top}\left[\mathbf{U}\right]_{:,0:L-1}^{\top}$
+$=\sum_{l=0}^{L-1}\bm{e}[l]\left[\mathbf{U}\right]_{:,l}\left[\mathbf{U}\right]_{:,l}^{\top},$ (11)
+and is expected to approach $\mathbf{\Psi}$ when $L$ is sufficiently large since most of the eigenvalues of $\mathbf{\Psi}$ are close to zero.
+In this paper, we set $L=\frac{N}{2}$.
 In the following analysis, $\bm{\eta}$ is estimated, instead of $\bm{\theta}$.
@@
 The superscript ``$^{q}"$ indicates the $q$-th iteration.
-Set the partial derivative of (III-A2) regarding $\bm{\eta}$ to zero, i.e.,
+Setting the partial derivative of (III-A2) regarding $\bm{\eta}$ to zero,
 $\frac{\partial\mathcal{L}^{[q+1]}_{\bm{\eta}}}{\partial\bm{\eta}}=-\frac{2}{\sigma^{2}}\Re\left\{\mathbf{Q}^{\rm H}\bar{\bm{y}}\right\}+\frac{2}{\sigma^{2}}\Re\left\{\mathbf{Q}^{\rm H}\mathbf{Q}\right\}\bm{\eta}+\bm{\eta}=\bm{0}_{L\times 1},$ (14)
@@
 $\hat{\phi}\propto\arg\min_{\phi}\mathcal{L}_{\phi}\propto\arg\min_{\phi}\|\bm{y}-\bm{\mu}\|^{2}.$ (16)
-To enhance the tractability of (14) and derive a closed-form solution, the first-order Taylor approximation is employed. The $n$-th term on the diagonal of $\bm{\Lambda}_{\phi}$ can be approximated as
-$\left[\bm{\Lambda}_{\phi}\right]_{n,n}=e^{j2\pi\frac{n}{N}\phi}\approx e^{j2\pi\frac{n}{N}\hat{\phi}}+\left(\phi-\hat{\phi}\right)\frac{j2\pi}{N}e^{j2\pi\frac{n}{N}\hat{\phi}}.$ (15)
+To enhance the tractability of (16) and derive a closed-form solution, the first-order Taylor approximation is employed [38, 39].
+The $n$-th term on the diagonal of $\bm{\Lambda}_{\phi}$ can be approximated as
+$\left[\bm{\Lambda}_{\phi}\right]_{n,n}=e^{j2\pi\frac{n}{N}\phi}\approx e^{j2\pi\frac{n}{N}\hat{\phi}}+\left(\phi-\hat{\phi}\right)\frac{j2\pi{\color[rgb]{0,0,0}n}}{N}e^{j2\pi\frac{n}{N}\hat{\phi}}.$ (17)
 Hence, $\hat{\bm{\Lambda}}_{\phi}^{[q+1]}$ can be approximated as
 $\hat{\bm{\Lambda}}_{\phi}^{[q+1]}\approx\hat{\bm{\Lambda}}_{\phi}^{[q]}+\left(\phi-\hat{\phi}^{[q]}\right)\tilde{\bm{\Lambda}}_{\phi}^{[q]},$ (18)
-where $\left[\tilde{\bm{\Lambda}}_{\phi}^{[q]}\right]_{n,n}=\frac{j2\pi n}{N}e^{\frac{j2\pi n\hat{\phi}^{[q]}}{N}}$.
+where $\left[\tilde{\bm{\Lambda}}_{\phi}^{[q]}\right]_{n,n}=\frac{j2\pi n}{N}e^{\frac{j2\pi n\hat{\phi}^{[q]}}{N}}$. By defining $x=\phi-\hat{\phi}$, the approximation in (17) can be rewritten as
+$f(x)=e^{j2\pi\frac{n}{N}x}\approx 1+\frac{j2\pi{\color[rgb]{0,0,0}n}}{N}x,$ (19)
+which is essentially the first-order Taylor series expansion of $f(x)$ at $x=0$, with $x$ being the estimation accuracy of $\phi$.
+The Lagrange remainder of this Taylor expansion, which reflects the approximation accuracy, can be calculated as
+${\rm R}_{1}(x)$ $=\frac{f^{{}^{\prime\prime}}(\xi)}{2}x^{2}=-2\pi^{2}\frac{n^{2}}{N^{2}}e^{j2\pi\frac{n}{N}\xi}x^{2},$ (20)
+where $\xi\in(0,x)$.
+Since $|e^{j2\pi\frac{n}{N}\xi}|=1$, $|{\rm R}_{1}(x)|\leq 2\pi^{2}\frac{n^{2}}{N^{2}}x^{2}$.
+The remainder of the first-order Taylor expansion of $f(x)$ is upper bounded, and the upper bound depends on the estimation accuracy of $\phi$.
+As the estimation accuracy of $\phi$ improves, the Taylor approximation in (17) becomes increasingly accurate.
 With (18), $\mathcal{L}_{\phi}$ in (16) can be rewritten as
@@
 III-C Position Estimation
-For position estimation, we employ the GD algorithm.
-Due to the same structures in the respective derivatives of the objective function with regard to the three position parameters $d_{Ru}$, $\phi_{az}$, and $\phi_{el}$, we use the symbol $\xi$ to represent any one of them.
+The GD algorithm is a general method for searching maxima or minima in unconstrained optimization problems and has been widely used to solve localization problems in existing studies [40, 41, 42].
+Under the given CFO and PN estimates, the MAP-based problem in (9) is an unconstrained optimization problem.
+We employ the GD algorithm to estimate UE position.
+Due to the same structures in the respective derivatives of the objective function with regard to the three position parameters ${\color[rgb]{0,0,0}d_{\text{R}}}$, ${\color[rgb]{0,0,0}\phi_{\text{az}}}$, and ${\color[rgb]{0,0,0}\phi_{\text{el}}}$, we use the symbol $\xi$ to represent any one of them.
 The partial derivative of the objective function with regard to $\xi$ is given by
@@
 where $\frac{\partial\bm{h}}{\partial\xi}$ is calculated by
-$\frac{\partial\bm{h}}{\partial\xi}[n]$ $=\frac{\partial\bm{h}[n]}{\partial\xi}=-\sum_{r=1}^{N_{R}}\sqrt{\rho_{aR,n,r}}\sqrt{\rho_{Ru,n,r}}\bm{w}[r]$
-×$\times e^{-j2\pi\frac{nB}{N}(\tau_{ar}+\tau_{ru})}\left(\frac{1}{d_{ru}}+j2\pi\frac{nB}{Nc}\right)\frac{\partial d_{ru}}{\partial\xi},$ (20)
-with $\frac{\partial d_{ru}}{\partial\xi}$, $\xi=d_{Ru},\phi_{az},\phi_{el}$, being
-$\frac{\partial d_{ru}}{\partial d_{Ru}}$ $=\frac{1}{d_{Ru}}\left[(d_{Ru}\sin\phi_{el}\cos\phi_{az}-x_{r})\sin\phi_{el}\cos\phi_{az}\right.$
-$\left.\quad+(d_{Ru}\sin\phi_{el}\sin\phi_{az}-y_{r})\sin\phi_{el}\sin\phi_{az}\right.$
-$\left.\quad+(d_{Ru}\cos\phi_{el}-z_{r})\cos\phi_{el}\right],$ (21)
-$\frac{\partial d_{ru}}{\partial\phi_{az}}$ $=\frac{1}{d_{Ru}}\left[(d_{Ru}\sin\phi_{el}\cos\phi_{az}-x_{r})(-\sin\phi_{az}d_{Ru}\sin\phi_{el})\right.$
-$\left.\quad+(d_{Ru}\sin\phi_{el}\sin\phi_{az}-y_{r})d_{Ru}\sin\phi_{el}\cos\phi_{az}\right],$ (22)
+$\frac{\partial\bm{h}}{\partial\xi}[n]$ $=\frac{\partial\bm{h}[n]}{\partial\xi}=-\sum_{r=0}^{{\color[rgb]{0,0,0}N_{\text{R}}-1}}{\color[rgb]{0,0,0}\sqrt{\rho_{{\text{aR}}}^{n,r}}}{\color[rgb]{0,0,0}\sqrt{\rho_{{\text{Ru}}}^{n,r}}}\bm{w}[r]$
+×$\times e^{-j2\pi\frac{nB}{N}({\tau_{{\color[rgb]{0,0,0}\text{a}}r}}+{\color[rgb]{0,0,0}\tau_{r\text{u}}})}\left(\frac{1}{{\color[rgb]{0,0,0}d_{r\text{u}}}}+j2\pi\frac{nB}{Nc}\right)\frac{\partial{\color[rgb]{0,0,0}d_{r\text{u}}}}{\partial\xi},$ (24)
+with $\frac{\partial{\color[rgb]{0,0,0}d_{r\text{u}}}}{\partial\xi}$, $\xi={\color[rgb]{0,0,0}d_{\text{R}}},{\color[rgb]{0,0,0}\phi_{\text{az}}},{\color[rgb]{0,0,0}\phi_{\text{el}}}$, being
+$\frac{\partial{\color[rgb]{0,0,0}d_{r\text{u}}}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}$ $=\frac{1}{{\color[rgb]{0,0,0}{\color[rgb]{0,0,0}d_{r\text{u}}}}}\left[({\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\cos{\color[rgb]{0,0,0}\phi_{\text{az}}}-x_{r})\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\cos{\color[rgb]{0,0,0}\phi_{\text{az}}}\right.$
+$\left.\quad+({\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\sin{\color[rgb]{0,0,0}\phi_{\text{az}}}-y_{r})\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\sin{\color[rgb]{0,0,0}\phi_{\text{az}}}\right.$
+$\left.\quad+({\color[rgb]{0,0,0}d_{\text{R}}}\cos{\color[rgb]{0,0,0}\phi_{\text{el}}}-z_{r})\cos{\color[rgb]{0,0,0}\phi_{\text{el}}}\right],$ (25)
+$\frac{\partial{\color[rgb]{0,0,0}d_{r\text{u}}}}{\partial{\color[rgb]{0,0,0}\phi_{\text{az}}}}$ $=\frac{1}{{\color[rgb]{0,0,0}{\color[rgb]{0,0,0}d_{r\text{u}}}}}\left[({\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\cos{\color[rgb]{0,0,0}\phi_{\text{az}}}-x_{r})(-\sin{\color[rgb]{0,0,0}\phi_{\text{az}}}{\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}})\right.$
+$\left.\quad+({\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\sin{\color[rgb]{0,0,0}\phi_{\text{az}}}-y_{r}){\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\cos{\color[rgb]{0,0,0}\phi_{\text{az}}}\right],$ (26)
 and
-$\frac{\partial d_{ru}}{\partial\phi_{el}}$ $=\frac{1}{d_{Ru}}\left[(d_{Ru}\sin\phi_{el}\cos\phi_{az}-x_{r})d_{Ru}\cos\phi_{az}\cos\phi_{el}\right.$
-$\left.\quad+(d_{Ru}\sin\phi_{el}\sin\phi_{az}-y_{r})d_{Ru}\sin\phi_{az}\cos\phi_{el}\right.$
-$\left.\quad+(d_{Ru}\cos\phi_{el}-z_{r})(-d_{Ru}\sin\phi_{el})\right].$ (23)
+$\frac{\partial{\color[rgb]{0,0,0}d_{r\text{u}}}}{\partial{\color[rgb]{0,0,0}\phi_{\text{el}}}}$ $=\frac{1}{{\color[rgb]{0,0,0}{\color[rgb]{0,0,0}d_{r\text{u}}}}}\left[({\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\cos{\color[rgb]{0,0,0}\phi_{\text{az}}}-x_{r}){\color[rgb]{0,0,0}d_{\text{R}}}\cos{\color[rgb]{0,0,0}\phi_{\text{az}}}\cos{\color[rgb]{0,0,0}\phi_{\text{el}}}\right.$
+$\left.\quad+({\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\sin{\color[rgb]{0,0,0}\phi_{\text{az}}}-y_{r}){\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{az}}}\cos{\color[rgb]{0,0,0}\phi_{\text{el}}}\right.$
+$\left.\quad+({\color[rgb]{0,0,0}d_{\text{R}}}\cos{\color[rgb]{0,0,0}\phi_{\text{el}}}-z_{r})(-{\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}})\right].$ (27)
 III-D Overall Algorithm
 The proposed joint estimation of the CFO, PN and UE position is summarized in Algorithm 1.
-Specifically, we initialize the parameters, including the initial estimation of the normalized CFO $\hat{\phi}$ and UE position $\hat{\bm{p}}_{u}$, the step size, $k_{p}$, the convergence accuracy for outer and inner loops, $\epsilon_{o}$ and $\epsilon_{i}$, and the maximum iteration number for the outer and inner loops, $N_{o}$ and $N_{i}$.
-After transforming the coordinates into the polar coordinates, we estimate the PN, CFO, and UE position alternately.
+Specifically, we initialize the parameters, including the initial estimation of the normalized CFO $\hat{\phi}$ and UE position $\hat{\bm{p}}_{\color[rgb]{0,0,0}\text{u}}$, the step size $k_{\color[rgb]{0,0,0}\text{p}}$, the convergence accuracy requirements for outer and inner loops, $\epsilon_{\color[rgb]{0,0,0}\text{O}}$ and $\epsilon_{\color[rgb]{0,0,0}\text{I}}$, and the maximum numbers of iterations for the outer and inner loops, $N_{\color[rgb]{0,0,0}\text{O}}$ and $N_{\color[rgb]{0,0,0}\text{I}}$.
+Suppose that the range of the normalized CFO $\phi$ is known at the UE, and the initial $\hat{\phi}$ is chosen randomly within this range.
+The initial estimation of the UE position, $\hat{\bm{p}}_{\color[rgb]{0,0,0}\text{u}}$, is obtained using the state of the art proposed in [26].
+The other initial parameters are hyperparameters, and can be fine-tuned to balance both convergence speed and convergence accuracy.
+Then, after transforming the coordinates into the polar coordinates, we estimate the PN, CFO, and UE position alternately.
 Closed-form expressions are utilized to update the PN and CFO estimates in each iteration, while the GD algorithm is employed to estimate the UE position.
 This repeats until either the objective function $\mathcal{L}$ converges or the maximum number of iterations is reached.
-It is noteworthy that this AO-based algorithm takes the RIS phase shifts, i.e., $\bm{w}$, as input, which can affect the algorithm performance significantly.
+Notably, this AO-based algorithm takes the RIS phase shifts, i.e., $\bm{w}$, as input, which can affect the algorithm performance significantly.
 The configuration of the RIS phase shifts, $\bm{w}$, will be discussed in Section V.
@@
 Algorithm 1 converges.
-The computational complexity for updating $\bm{\eta}$, $\bm{\theta}$, $\phi$, and position parameters $d_{Rt}$, $\phi_{az}$, and $\phi_{el}$ in Algorithm 1 are $\mathcal{O}(L^{3}+N^{2}+NL^{2})$, $\mathcal{O}(NL)$, $\mathcal{O}(N^{2})$, and $\mathcal{O}(\max\left\{N^{3},N_{R}\right\})$, respectively.
-Suppose the inner GD algorithm for updating the UE position converges after $K_{i}$ iterations.
-Consequently, the computational complexity of Algorithm 1 is $\mathcal{O}(\max\left\{L^{3}+N^{2}+NL^{2},K_{i}N^{3},K_{i}N_{R}\right\})$ per iteration.
+The computational complexity for updating $\bm{\eta}$, $\bm{\theta}$, $\phi$, and position parameters ${\color[rgb]{0,0,0}d_{\text{R}}}$, ${\color[rgb]{0,0,0}\phi_{\text{az}}}$, and ${\color[rgb]{0,0,0}\phi_{\text{el}}}$ in Algorithm 1 are $\mathcal{O}(L^{3}+N^{2}+NL^{2})$, $\mathcal{O}(NL)$, $\mathcal{O}(N^{2})$, and $\mathcal{O}(\max\left\{N^{3},{\color[rgb]{0,0,0}N_{\text{R}}}\right\})$, respectively.
+Suppose the inner GD algorithm for updating the UE position converges after $K_{\color[rgb]{0,0,0}\text{I}}$ iterations.
+Consequently, the computational complexity of Algorithm 1 is $\mathcal{O}(\max\left\{L^{3}+N^{2}+NL^{2},K_{\color[rgb]{0,0,0}\text{I}}N^{3},K_{\color[rgb]{0,0,0}\text{I}}{\color[rgb]{0,0,0}N_{\text{R}}}\right\})$ per iteration.
+By contrast, the complexity of the UE localization method in [26] is $\mathcal{O}(\max(K_{\text{I}}N^{3},K_{\text{I}}N_{\text{R}}))$ per iteration.
+Nevertheless, by adopting an AO-based approach to jointly estimate the UE position, CFO, and PN, our algorithm can achieve up to a two-order-of-magnitude improvement in localization accuracy under high SNR conditions, compared to the method developed in [26], as demonstrated numerically in Section VI-C.
 Algorithm 1 Proposed iterative joint estimation algorithm.
 1:  Input: $\bm{w}$;
-2:  Initialization: $\hat{\phi}$, $\hat{\bm{p}}_{u}$, $k_{p}$, $\epsilon_{o}$, $\epsilon_{i}$, $N_{o}$, $N_{i}$;
+2:  Initialization: $\hat{\phi}$, $\hat{\bm{p}}_{\color[rgb]{0,0,0}\text{u}}$, $k_{\color[rgb]{0,0,0}\text{p}}$, $\epsilon_{\color[rgb]{0,0,0}\text{O}}$, $\epsilon_{\color[rgb]{0,0,0}\text{I}}$, $N_{\color[rgb]{0,0,0}\text{O}}$, $N_{\color[rgb]{0,0,0}\text{I}}$;
 3:  Calculate polar coordinate of UE using (1);
 4:  repeat
-5:     Given fixed $\hat{\bm{p}}_{u}$, update $\hat{\bm{\eta}}$ with $\hat{\phi}$ using (13);
+5:     Given fixed $\hat{\bm{p}}_{\color[rgb]{0,0,0}\text{u}}$, update $\hat{\bm{\eta}}$ with $\hat{\phi}$ using (15);
 6:     Update $\hat{\bm{\theta}}$ using (10);
-7:     Given fixed $\hat{\bm{p}}_{u}$, update $\hat{\phi}$ with $\hat{\bm{\theta}}$ using (18);
+7:     Given fixed $\hat{\bm{p}}_{\color[rgb]{0,0,0}\text{u}}$, update $\hat{\phi}$ with $\hat{\bm{\theta}}$ using (22);
 8:     repeat
-9:        $\hat{d}_{Ru}\leftarrow\hat{d}_{Ru}-\frac{\partial\mathcal{L}_{\xi}}{\partial d_{Ru}}k_{p}$;
-10:        $\hat{\phi}_{az}\leftarrow\hat{\phi}_{az}-\frac{\partial\mathcal{L}_{\xi}}{\partial\phi_{az}}k_{p}$;
-11:        $\hat{\phi}_{el}\leftarrow\hat{\phi}_{el}-\frac{\partial\mathcal{L}_{\xi}}{\partial\phi_{el}}k_{p}$;
-12:     until $\mathcal{L}^{[p]}-\mathcal{L}^{[p+1]}\leq\epsilon_{i}$ after the $p$-th iteration or $p\textgreater N_{i}$.
-13:  until $\mathcal{L}^{[q]}-\mathcal{L}^{[q+1]}\leq\epsilon_{o}$ after the $q$-th iteration or $q\textgreater N_{o}$.
+9:        $\hat{d}_{\color[rgb]{0,0,0}\text{R}}\leftarrow\hat{d}_{\color[rgb]{0,0,0}\text{R}}-\frac{\partial\mathcal{L}_{\xi}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}k_{\color[rgb]{0,0,0}\text{p}}$;
+10:        $\hat{\phi}_{\color[rgb]{0,0,0}\text{az}}\leftarrow\hat{\phi}_{\color[rgb]{0,0,0}\text{az}}-\frac{\partial\mathcal{L}_{\xi}}{\partial{\color[rgb]{0,0,0}\phi_{\text{az}}}}k_{\color[rgb]{0,0,0}\text{p}}$;
+11:        $\hat{\phi}_{\color[rgb]{0,0,0}\text{el}}\leftarrow\hat{\phi}_{\color[rgb]{0,0,0}\text{el}}-\frac{\partial\mathcal{L}_{\xi}}{\partial{\color[rgb]{0,0,0}\phi_{\text{el}}}}k_{\color[rgb]{0,0,0}\text{p}}$;
+12:     until $\mathcal{L}^{[p]}-\mathcal{L}^{[p+1]}\leq\epsilon_{\color[rgb]{0,0,0}\text{I}}$ after the $p$-th iteration or $p\textgreater N_{\color[rgb]{0,0,0}\text{I}}$.
+13:  until $\mathcal{L}^{[q]}-\mathcal{L}^{[q+1]}\leq\epsilon_{\color[rgb]{0,0,0}\text{O}}$ after the $q$-th iteration or $q\textgreater N_{\color[rgb]{0,0,0}\text{O}}$.
 14:  Calculate Cartesian coordinate of UE using (2);
-15:  Output: $\hat{\phi}$, $\hat{\bm{\theta}}$, $\hat{\bm{p}}_{u}$.
+15:  Output: $\hat{\phi}$, $\hat{\bm{\theta}}$, $\hat{\bm{p}}_{\color[rgb]{0,0,0}\text{u}}$.
 III-E Performance Metrics
@@
 For the position estimation, we take root mean square error (RMSE) as the performance index.
-For any testing UE position $\bm{p}_{u}$, ${\rm RMSE}_{u}$ can be expressed as
-${\rm RMSE}_{u}=\sqrt{\frac{1}{M_{e}}\sum_{m=1}^{M_{e}}\|\hat{\bm{p}}_{u,m}-\bm{p}_{u}\|^{2}},$ (24)
-where $M_{e}$ is the number of Monte Carlo tests, and $\hat{\bm{p}}_{u,m}$ denotes the estimation of the UE position in the $m$-th trial.
+For any testing UE position $\bm{p}_{\color[rgb]{0,0,0}\text{u}}$, ${\rm RMSE}_{\color[rgb]{0,0,0}\text{u}}$ can be expressed as
+${\rm RMSE}_{\color[rgb]{0,0,0}\text{u}}=\sqrt{\frac{1}{M_{\color[rgb]{0,0,0}\text{e}}}{\color[rgb]{0,0,0}\sum_{m=0}^{M_{\text{e}}-1}}\|\hat{\bm{p}}_{{\color[rgb]{0,0,0}\text{u}},m}-\bm{p}_{\color[rgb]{0,0,0}\text{u}}\|^{2}},$ (28)
+where $M_{\color[rgb]{0,0,0}\text{e}}$ is the number of Monte Carlo tests, and $\hat{\bm{p}}_{{\color[rgb]{0,0,0}\text{u}},m}$ denotes the estimation of the UE position in the $m$-th trial.
 For the CFO and PN estimation, it can be straightforwardly observed from (6) that the received signal $\bm{y}$ does not change under a common phase rotation $\epsilon$, i.e.,
@@
 where $\bm{\underline{\gamma}}=\bm{\gamma}-\bm{\gamma}[0]\bm{1}$, $\hat{\bm{\underline{\gamma}}}=\hat{\bm{\gamma}}-\hat{\bm{\gamma}}[0]\bm{1}$, $\bm{\gamma}=[\bm{\gamma}[0],\bm{\gamma}[1],\cdots,\bm{\gamma}[N-1]]^{\top}$, $\bm{\gamma}[n]=\bm{\theta}[n]+\frac{2\pi n\phi}{N}$, $\hat{\bm{\gamma}}=[\hat{\bm{\gamma}}[0],\hat{\bm{\gamma}}[1],\cdots,\hat{\bm{\gamma}}[N-1]]^{\top}$, and $\hat{\bm{\gamma}}[n]=\hat{\bm{\theta}}[n]+\frac{2\pi n\hat{\phi}}{N}$.
-Using this measure, the performance of the joint CFO-PN estimation rather than each of the individual variables alone is considered.
+By focusing on joint CFO-PN estimation rather than individual variables, we ensure a more comprehensive performance evaluation.
+Subtracting the initial elements of $\bm{\gamma}$ and $\hat{\bm{\gamma}}$, specifically the actual PN $\bm{\theta}[0]$ and its estimate $\hat{\bm{\theta}}[0]$ at time 00, helps mitigate the influence of the initial PN $\bm{\theta}[0]$ (and its estimate $\hat{\bm{\theta}}[0]$) on the MSE analysis and resulting in a more precise evaluation of the joint CFO-PN estimation.
 IV Hybrid Cramér-Rao Lower Bound Analysis
-The HCRLB for the joint estimation of the CFO, PN, and UE position is derived in this section.
-This helps quantify the performance lower bound of the proposed AO-based estimation algorithm, and also serves as the objective function for optimizing the RIS phase shifts, $\bm{w}$, in the next section.
-The HCRLB for the estimation of $\bm{\lambda}$ is given by [34, Eq.
-(4.610)]
-$\mathbb{E}_{\bm{y},\bm{\theta}\mid\phi,d_{Ru},\phi_{az},\phi_{el}}\left[(\bm{\lambda}-\hat{\bm{\lambda}})(\bm{\lambda}-\hat{\bm{\lambda}})^{\top}\right]\succeq\mathbf{B}^{\rm-1},$ (27)
-where $\mathbf{B}$ is the Bayesian information matrix (BIM), which can be given by [34, Eq. (4.611)]
+HCRLB for the joint estimation of the CFO, PN, and UE position is derived in this section.
+This helps quantify the performance lower bound of the proposed AO-based estimation algorithm, and serves as the objective function for optimizing the RIS phase shifts, $\bm{w}$, in the next section.
+HCRLB is an extension of the traditional CRLB, suitable for estimation problems with both deterministic and random unknown variables.
+Unlike CRLB, HCRLB exploits the prior distribution of the unknown random variables to improve analysis.
+Since we estimate the deterministic variables, including CFO and UE position, as well as the random variable, i.e., PN, analyzing the HCRLB allows for more accurate estimation error lower bounds for these variables.
+The HCRLB for the estimate of $\bm{\lambda}$ is [43, Eq. (4.610)]
+$\mathbb{E}_{\bm{y},\bm{\theta}\mid\phi,{\color[rgb]{0,0,0}d_{\text{R}}},{\color[rgb]{0,0,0}\phi_{\text{az}}},{\color[rgb]{0,0,0}\phi_{\text{el}}}}\left[(\bm{\lambda}-\hat{\bm{\lambda}})(\bm{\lambda}-\hat{\bm{\lambda}})^{\top}\right]\succeq\mathbf{B}^{\rm-1},$ (31)
+where $\mathbf{B}$ is the Bayesian information matrix (BIM), which can be given by [43, Eq. (4.611)]
 $\mathbf{B}=\mathbb{E}_{\bm{\theta}}\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]+\mathbb{E}_{\bm{\theta}}\left[-\frac{\partial}{\partial\bm{\lambda}}\left(\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}\right)^{\top}\right],$ (32)
@@
 $\frac{\partial\bm{\mu}}{\partial\bm{\theta}[n]}$ $=\sqrt{P}{\rm Diag}(\bm{a}[n])\bm{\Lambda}_{\phi}\mathbf{S}^{\top}\bm{h};$ (34b)
-$\frac{\partial\bm{\mu}}{\partial d_{Ru}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial d_{Ru}};$ (30c)
-$\frac{\partial\bm{\mu}}{\partial\phi_{az}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial\phi_{az}};$ (30d)
-$\frac{\partial\bm{\mu}}{\partial\phi_{el}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial\phi_{el}},$ (30e)
+$\frac{\partial\bm{\mu}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}};$ (34c)
+$\frac{\partial\bm{\mu}}{\partial{\color[rgb]{0,0,0}\phi_{\text{az}}}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial{\color[rgb]{0,0,0}\phi_{\text{az}}}};$ (34d)
+$\frac{\partial\bm{\mu}}{\partial{\color[rgb]{0,0,0}\phi_{\text{el}}}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial{\color[rgb]{0,0,0}\phi_{\text{el}}}},$ (34e)
 where $\bm{\Lambda}$ is a diagnal matrix, $\left[\bm{\Lambda}\right]_{n,n}=\frac{j2\pi n}{N}$, and $\bm{a}[n]=[\bm{0}_{1\times n},je^{j\bm{\theta}[n]},\bm{0}_{1\times(N-n-1)}]^{\top}$.
@@
 $\mathbb{E}_{\bm{\theta}}\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]=\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda}).$ (35)
-Take $\frac{\partial\bm{\mu}^{\rm H}}{\partial d_{Ru}}\frac{\partial\bm{\mu}}{\partial d_{Ru}}$ as an example. We have
-$\frac{\partial\bm{\mu}^{\rm H}}{\partial d_{Ru}}\frac{\partial\bm{\mu}}{\partial d_{Ru}}$ $=\sqrt{P}\frac{\partial\bm{h}^{\rm H}}{\partial d_{Ru}}\mathbf{S}^{*}\bm{\Lambda}_{\bm{\theta}}^{\rm{H}}\bm{\Lambda}_{\phi}^{\rm H}\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial d_{Ru}}$ (32)
-$=P\frac{\partial\bm{h}^{\rm H}}{\partial d_{Ru}}\mathbf{S}^{*}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial d_{Ru}}$
+Take $\frac{\partial\bm{\mu}^{\rm H}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}\frac{\partial\bm{\mu}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}$ as an example. We have
+$\frac{\partial\bm{\mu}^{\rm H}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}\frac{\partial\bm{\mu}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}$ $=\sqrt{P}\frac{\partial\bm{h}^{\rm H}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}\mathbf{S}^{*}\bm{\Lambda}_{\bm{\theta}}^{\rm{H}}\bm{\Lambda}_{\phi}^{\rm H}\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}$ (36)
+$=P\frac{\partial\bm{h}^{\rm H}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}\mathbf{S}^{*}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}},$
 which is unrelated to $\phi$ or $\bm{\theta}$, since $\bm{\Lambda}_{\phi}^{\rm H}\bm{\Lambda}_{\phi}=\bm{\Lambda}_{\bm{\theta}}^{\rm{H}}\bm{\Lambda}_{\bm{\theta}}=\mathbf{I}_{N}$.
@@
 IV-B Derivation of $\mathbb{E}_{\bm{\theta}}\left[-\frac{\partial}{\partial\bm{\lambda}}\left(\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}\right)^{\top}\right]$
-Since $\bm{\theta}\sim\mathcal{N}(\bm{0}_{N\times 1},\bm{\Psi})$, we can calculate $\log p(\bm{\theta})$ as $\log p(\bm{\theta})=-\frac{1}{2}\bm{\theta}^{\top}\bm{\Psi}^{-1}\bm{\theta}-\frac{N}{2}\log(2\pi)-\frac{1}{2}\log\det\bm{\Psi}$, and its partial derivative with respect to $\bm{\lambda}$ can be calculated by
+Since $\bm{\theta}\sim\mathcal{N}(\bm{0}_{N\times 1},\bm{\Psi})$, we can calculate $\log p(\bm{\theta})$ as $\log p(\bm{\theta})=-\frac{1}{2}\bm{\theta}^{\top}\bm{\Psi}^{-1}\bm{\theta}-\frac{N}{2}\log(2\pi)-\frac{1}{2}\log\det\bm{\Psi}$, and its partial derivative with respect to $\bm{\lambda}$ can be calculated as
 $\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}=\left[0,-(\bm{\Psi}^{-1}\bm{\theta})^{\top},0,0,0\right]^{\top}.$ (37)
-Subsequently, $\mathbb{E}_{\bm{\theta}}\left[-\frac{\partial}{\partial\bm{\lambda}}\left(\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}\right)^{\top}\right]$ can be expressed by
+Subsequently, $\mathbb{E}_{\bm{\theta}}\left[-\frac{\partial}{\partial\bm{\lambda}}\left(\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}\right)^{\top}\right]$ can be expressed as
 $\mathbb{E}_{\bm{\theta}}\!\left[-\!\frac{\partial}{\partial\bm{\lambda}}\left(\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}\right)^{\top}\right]$ $=\!-\frac{\partial}{\partial\bm{\lambda}}\left(\!\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}\!\right)^{\top}$
 $=$ $\mathrm{Blkdiag}(0,\bm{\Psi}^{-1},\bm{0}_{3\times 3}).$ (38)
-IV-C Derivation of the Transformed HCRLB
-So far, we have derived the expression for $\mathbf{B}$, which is an HCRLB matrix that gives analytical estimation accuracy lower bounds for the parameters in $\bm{\lambda}$.
+IV-C Derivation of Transformed HCRLB
+So far, we have derived the expression for $\mathbf{B}$, whose inverse is an HCRLB matrix that gives analytical estimation accuracy lower bounds for the parameters in $\bm{\lambda}$.
 However, when assessing the proposed estimation algorithm, we adopt the RMSE and CFO-PN joint MSE metrics introduced in Section III-E.
-Hence, we define a new parameter vector $\bm{\alpha}=\bm{g}(\bm{\lambda})\triangleq\left[\bm{\underline{\gamma}},x_{u},y_{u},z_{u}\right]^{\top}$, corresponding to the two metrics.
-Then, $\mathbf{B}$ can be transformed into the $\bm{\alpha}$-related HCRLB matrix by [35]
-$\mathbf{HCRLB}(\bm{p}_{u},\bm{w})=\bm{\Xi}(\bm{p}_{u})\mathbf{B}^{-1}(\bm{p}_{u},\bm{w})\bm{\Xi}^{\top}({\bm{p}_{u}}),$ (35)
-where, for any UE position $\bm{p}_{u}$, $\mathbf{\Xi}(\bm{p}_{u})$ is the transition matrix defined as
-$\bm{\Xi}(\bm{p}_{u})=\frac{\partial\bm{g}(\bm{\lambda})}{\partial\bm{\lambda}}={\rm Blkdiag}\left(\bm{\Xi}_{1},\bm{\Xi}_{2}(\bm{p}_{u})\right),$ (36)
-with $\bm{\Xi}_{1}$ and $\bm{\Xi}_{2}(\bm{p}_{u})$ given by (37) and (38), respectively.
+Hence, we define a new parameter vector $\bm{\alpha}=\bm{g}(\bm{\lambda})\triangleq\left[\bm{\underline{\gamma}},x_{\color[rgb]{0,0,0}\text{u}},y_{\color[rgb]{0,0,0}\text{u}},z_{\color[rgb]{0,0,0}\text{u}}\right]^{\top}$, corresponding to the two metrics.
+Then, $\mathbf{B}$ can be transformed into the $\bm{\alpha}$-related HCRLB matrix [44]
+$\mathbf{HCRLB}(\bm{p}_{\color[rgb]{0,0,0}\text{u}},\bm{w})=\bm{\Xi}(\bm{p}_{\color[rgb]{0,0,0}\text{u}})\mathbf{B}^{-1}(\bm{p}_{\color[rgb]{0,0,0}\text{u}},\bm{w})\bm{\Xi}^{\top}({\bm{p}_{\color[rgb]{0,0,0}\text{u}}}),$ (39)
+where, for any UE position $\bm{p}_{\color[rgb]{0,0,0}\text{u}}$, $\mathbf{\Xi}(\bm{p}_{\color[rgb]{0,0,0}\text{u}})$ is the transition matrix defined as
+$\bm{\Xi}(\bm{p}_{\color[rgb]{0,0,0}\text{u}})=\frac{\partial\bm{g}(\bm{\lambda})}{\partial\bm{\lambda}}={\rm Blkdiag}\left(\bm{\Xi}_{1},\bm{\Xi}_{2}(\bm{p}_{\color[rgb]{0,0,0}\text{u}})\right),$ (40)
+with $\bm{\Xi}_{1}$ and $\bm{\Xi}_{2}(\bm{p}_{\color[rgb]{0,0,0}\text{u}})$ given by (41) and (42), respectively.
 $\bm{\Xi}_{1}=\begin{bmatrix}\frac{\partial\bm{\underline{\gamma}}[0]}{\partial\phi}&\frac{\partial\bm{\underline{\gamma}}[0]}{\partial\bm{\theta}[0]}&\cdots&\frac{\partial\bm{\underline{\gamma}}[0]}{\partial\bm{\theta}[N-1]}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{\partial\bm{\underline{\gamma}}[N-1]}{\partial\phi}&\frac{\partial\bm{\underline{\gamma}}[N-1]}{\partial\bm{\theta}[0]}&\cdots&\frac{\partial\bm{\underline{\gamma}}[N-1]}{\partial\bm{\theta}[N-1]}\end{bmatrix}=\begin{bmatrix}0&0&0&\cdots&0\\ \frac{2\pi}{N}&-1&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \frac{2\pi(N-1)}{N}&-1&0&\cdots&1\\ \end{bmatrix}\in\mathbb{R}^{N\times(N+1)},$ (41)
-$\bm{\Xi}_{2}(\bm{p}_{u})=\begin{bmatrix}\frac{\partial x_{u}}{\partial d_{Ru}}&\frac{\partial x_{u}}{\partial\phi_{az}}&\frac{\partial x_{u}}{\partial\phi_{el}}\\ \frac{\partial y_{u}}{\partial d_{Ru}}&\frac{\partial y_{u}}{\partial\phi_{az}}&\frac{\partial y_{u}}{\partial\phi_{el}}\\ \frac{\partial z_{u}}{\partial d_{Ru}}&\frac{\partial z_{u}}{\partial\phi_{az}}&\frac{\partial z_{u}}{\partial\phi_{el}}\\ \end{bmatrix}=\begin{bmatrix}\sin\phi_{el}\cos\phi_{az}&-d_{Ru}\sin\phi_{el}\sin\phi_{az}&d_{Ru}\cos\phi_{az}\cos\phi_{el}\\ \sin\phi_{el}\sin\phi_{az}&d_{Ru}\sin\phi_{el}\cos\phi_{az}&d_{Ru}\sin\phi_{az}\cos\phi_{el}\\ \cos\phi_{el}&0&-d_{Ru}\sin\phi_{el}\end{bmatrix}\in\mathbb{R}^{3\times 3}.$ (38)
-After obtaining $\mathbf{HCRLB}(\bm{p}_{u},\bm{w})$, the position error bound (${\rm PEB}$) for evaluating the UE position estimation and $\rm HCRLB_{CFO-PN}$ for evaluating the CFO-PN joint estimation can be given by
-${\rm PEB}=\sqrt{{\rm tr}([\mathbf{HCRLB}(\bm{p}_{u},\bm{w})]_{N:N+2,N:N+2})};$ (39)
-${\rm HCRLB_{CFO-PN}}={\rm tr}([\mathbf{HCRLB}(\bm{p}_{u},\bm{w})]_{0:N-1,0:N-1}).$ (40)
-Both ${\rm PEB}$ and ${\rm HCRLB_{CFO-PN}}$ depend on the UE’s position, $\bm{p}_{u}$, and the RIS phase shifts, $\bm{w}$.
+$\bm{\Xi}_{2}(\bm{p}_{\color[rgb]{0,0,0}\text{u}})=\begin{bmatrix}\frac{\partial x_{\color[rgb]{0,0,0}\text{u}}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}&\frac{\partial x_{\color[rgb]{0,0,0}\text{u}}}{\partial{\color[rgb]{0,0,0}\phi_{\text{az}}}}&\frac{\partial x_{\color[rgb]{0,0,0}\text{u}}}{\partial{\color[rgb]{0,0,0}\phi_{\text{el}}}}\\ \frac{\partial y_{\color[rgb]{0,0,0}\text{u}}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}&\frac{\partial y_{\color[rgb]{0,0,0}\text{u}}}{\partial{\color[rgb]{0,0,0}\phi_{\text{az}}}}&\frac{\partial y_{\color[rgb]{0,0,0}\text{u}}}{\partial{\color[rgb]{0,0,0}\phi_{\text{el}}}}\\ \frac{\partial z_{\color[rgb]{0,0,0}\text{u}}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}&\frac{\partial z_{\color[rgb]{0,0,0}\text{u}}}{\partial{\color[rgb]{0,0,0}\phi_{\text{az}}}}&\frac{\partial z_{\color[rgb]{0,0,0}\text{u}}}{\partial{\color[rgb]{0,0,0}\phi_{\text{el}}}}\\ \end{bmatrix}=\begin{bmatrix}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\cos{\color[rgb]{0,0,0}\phi_{\text{az}}}&-{\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\sin{\color[rgb]{0,0,0}\phi_{\text{az}}}&{\color[rgb]{0,0,0}d_{\text{R}}}\cos{\color[rgb]{0,0,0}\phi_{\text{az}}}\cos{\color[rgb]{0,0,0}\phi_{\text{el}}}\\ \sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\sin{\color[rgb]{0,0,0}\phi_{\text{az}}}&{\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\cos{\color[rgb]{0,0,0}\phi_{\text{az}}}&{\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{az}}}\cos{\color[rgb]{0,0,0}\phi_{\text{el}}}\\ \cos{\color[rgb]{0,0,0}\phi_{\text{el}}}&0&-{\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\end{bmatrix}\in\mathbb{R}^{3\times 3}.$ (42)
+After obtaining $\mathbf{HCRLB}(\bm{p}_{\color[rgb]{0,0,0}\text{u}},\bm{w})$, the position error bound (${\rm PEB}$) for the UE position estimation and $\rm HCRLB_{CFO-PN}$ for the CFO-PN joint estimation can be given by
+${\rm PEB}=\sqrt{{\rm tr}([\mathbf{HCRLB}(\bm{p}_{\color[rgb]{0,0,0}\text{u}},\bm{w})]_{N:N+2,N:N+2})};$ (43)
+${\rm HCRLB_{CFO-PN}}={\rm tr}([\mathbf{HCRLB}(\bm{p}_{\color[rgb]{0,0,0}\text{u}},\bm{w})]_{0:N-1,0:N-1}).$ (44)
+Both ${\rm PEB}$ and ${\rm HCRLB_{CFO-PN}}$ depend on the UE’s position, $\bm{p}_{\color[rgb]{0,0,0}\text{u}}$, and the RIS phase shifts, $\bm{w}$.
 V RIS Configuration
-In this section, we further configure the phase shifts $\bm{w}$ of the RIS to minimize the analytical lower bound of the localization accuracy, i.e., PEB.
+In this section, we configure the phase shifts $\bm{w}$ of the RIS to minimize the analytical lower bound of the localization accuracy, the PEB derived in Section IV.
+The configuration effectively enhances the performance of the algorithm proposed in Section III, and is applied before the localization process.
 According to (43), the PEB depends on the UE position, which indicates that minimizing the PEB at a specific UE position requires knowledge of that position.
 However, this is not possible in practice.
-To address this issue and enhance the algorithm’s robustness, we take the average PEB within the area of interest (AOI) as the objective function.
-Specifically, we randomly sample $U$ positions within the AOI and calculate their average PEB.
+To address this issue and enhance the algorithm’s robustness, we specify the area of interest (AOI), according to the prior knowledge and practical requirements.
+The AOI represents the region in which we anticipate the UE of interest is situated and the high-precision localization of the UE is expected.
+The design objective of the RIS configuration is to achieve high expected accuracy for the localization of UEs located within the AOI, i.e., by minimizing the average PEB within the AOI.
+Specifically, we uniformly and randomly sample $S$ positions within the AOI, denoted as $\bm{p}_{s},s=0,1,\cdots,S-1$, and evaluate their average PEB.
+This problem is transformed into a convex SDP form using the SDR method and the Schur complement.
+Unlike traditional approaches that rely on the classical FIM, the objective function represented by the average PEB is derived from the inverse of the BIM, $\mathbf{B}$.
+To establish the convexity of the transformed problem, we analytically reveal the linear relationship between $\mathbf{B}$ and the optimization variable $\mathbf{W}$.
 Accordingly, the problem of interest is formulated as
-$\min_{\bm{w}}\quad\frac{1}{U}\sum_{u=1}^{U}{\rm PEB}(\bm{p}_{u},\bm{w})$
-$s.t.\quad|\bm{w}[r]|=1,r=1,\cdots,N_{R}.$ (41)
-Problem (V) is not convex because of the non-convexity of the objective function with respect to vector $\bm{w}$, as well as the unit-module constraint.
+$\min_{\bm{w}}\quad\frac{1}{\color[rgb]{0,0,0}S}{\color[rgb]{0,0,0}\sum_{s=0}^{S-1}}{\rm PEB}(\bm{p}_{\color[rgb]{0,0,0}s},\bm{w})$
+${\color[rgb]{0,0,0}\rm s.t.}\quad|\bm{w}[r]|=1,r={\color[rgb]{0,0,0}0,1,\cdots,}{\color[rgb]{0,0,0}N_{\text{R}}-1}.$ (45)
+Problem (V) is not convex because of the non-convexity of the objective function with respect to the vector $\bm{w}$, as well as the unit-module constraint.
 To convexify (V), we first define $\mathbf{W}\triangleq\bm{\bar{w}}\bm{\bar{w}}^{\rm H}$, where $\bm{\bar{w}}\triangleq\bm{w}^{*}$, which satisfies $\mathbf{W}\succcurlyeq 0$ and ${\rm rank}(\mathbf{W})=1$.
@@
 Consequently, the problem in (V) can be transformed into
-$\min_{\mathbf{W}}\quad\frac{1}{U}\sum_{u=1}^{U}{\rm PEB}(\bm{p}_{u},\mathbf{W})$
-$\begin{array}[]{ll}s.t.&\left[\mathbf{W}\right]_{r,r}=1,r=1,\cdots,N_{R},\\ &\mathbf{W}\succcurlyeq 0.\end{array}$ (44)
+$\min_{\mathbf{W}}\quad\frac{1}{\color[rgb]{0,0,0}S}{\color[rgb]{0,0,0}\sum_{s=0}^{S-1}}{\rm PEB}(\bm{p}_{\color[rgb]{0,0,0}s},\mathbf{W})$
+$\begin{array}[]{ll}{\color[rgb]{0,0,0}\rm s.t.}&\left[\mathbf{W}\right]_{r,r}=1,r={\color[rgb]{0,0,0}0,1},\cdots,{\color[rgb]{0,0,0}N_{\text{R}}-1},\\ &\mathbf{W}\succcurlyeq 0.\end{array}$ (48)
 According to (39) and (43), the optimization problem in (V) remains non-convex because of the complex objective function.
-To transform (V) into the convex SDP form, for each calculated PEB at a particular position $\bm{p}_{u}$, we define an auxiliary matrix $\mathbf{Z}_{u}$ that satisfies
-$\mathbf{Z}_{u}\succcurlyeq\tilde{\bm{\Xi}}(\bm{p}_{u})\mathbf{B}^{-1}(\bm{p}_{u},\mathbf{W})\tilde{\bm{\Xi}}^{\top}(\bm{p}_{u}),$ (45)
-where $\tilde{\bm{\Xi}}(\bm{p}_{u})\triangleq\mathbf{T}\mathbf{\Xi}(\bm{p}_{u})$, and $\mathbf{T}\triangleq[\bm{0}_{N\times N}\quad\mathbf{I}_{3}]\in\mathbb{R}^{(N+3)\times(N+3)}$ can extract the PEB from the HCRLB.
-According to the property of the Schur complement and the positive semidefiniteness of $\mathbf{B}(\bm{p}_{u},\mathbf{W})$, (45) can be transformed to
-$\begin{bmatrix}\mathbf{Z}_{u}&\tilde{\bm{\Xi}}(\bm{p}_{u})\\ \tilde{\bm{\Xi}}^{\top}(\bm{p}_{u})&\mathbf{B}(\bm{p}_{u},\mathbf{W})\end{bmatrix}\succcurlyeq 0.$ (46)
+To transform (V) into the convex SDP form, for each calculated PEB at a particular position $\bm{p}_{s}$, we define an auxiliary matrix $\mathbf{Z}_{\color[rgb]{0,0,0}s}$ that satisfies
+$\mathbf{Z}_{\color[rgb]{0,0,0}s}\succcurlyeq\tilde{\bm{\Xi}}(\bm{p}_{\color[rgb]{0,0,0}s})\mathbf{B}^{-1}(\bm{p}_{\color[rgb]{0,0,0}s},\mathbf{W})\tilde{\bm{\Xi}}^{\top}(\bm{p}_{\color[rgb]{0,0,0}s}),$ (49)
+where $\tilde{\bm{\Xi}}(\bm{p}_{\color[rgb]{0,0,0}s})\triangleq\mathbf{T}\mathbf{\Xi}(\bm{p}_{\color[rgb]{0,0,0}s})$, and $\mathbf{T}\triangleq[\bm{0}_{N\times N}\quad\mathbf{I}_{3}]\in\mathbb{R}^{(N+3)\times(N+3)}$ can extract the PEB from the HCRLB.
+According to the property of the Schur complement and the positive semidefiniteness of $\mathbf{B}(\bm{p}_{\color[rgb]{0,0,0}s},\mathbf{W})$, (49) can be transformed to
+$\begin{bmatrix}\mathbf{Z}_{\color[rgb]{0,0,0}s}&\tilde{\bm{\Xi}}(\bm{p}_{\color[rgb]{0,0,0}s})\\ \tilde{\bm{\Xi}}^{\top}(\bm{p}_{\color[rgb]{0,0,0}s})&\mathbf{B}(\bm{p}_{\color[rgb]{0,0,0}s},\mathbf{W})\end{bmatrix}\succcurlyeq 0.$ (50)
 Thus, the problem in (V) can be converted to
-$\min_{\mathbf{W},\mathbf{Z}_{1},\cdots,\mathbf{Z}_{U}}\quad\frac{1}{U}\sum_{u=1}^{U}{\rm tr}(\mathbf{Z}_{u})$
-$\begin{array}[]{ll}s.t.&\left[\mathbf{W}\right]_{r,r}=1,r=1,\cdots,N_{R},\\ &\mathbf{W}\succcurlyeq 0,\\ &\displaystyle\frac{1}{U}\sum_{u=1}^{U}\begin{bmatrix}\mathbf{Z}_{u}&\tilde{\bm{\Xi}}(\bm{p}_{u})\\ \tilde{\bm{\Xi}}^{\top}(\bm{p}_{u})&\mathbf{B}(\bm{p}_{u},\mathbf{W})\end{bmatrix}\succcurlyeq 0,\end{array}$ (50)
+$\min_{\mathbf{W},\mathbf{Z}_{\color[rgb]{0,0,0}0},\cdots,\mathbf{Z}_{\color[rgb]{0,0,0}S-1}}\quad\frac{1}{\color[rgb]{0,0,0}S}{\color[rgb]{0,0,0}\sum_{s=0}^{S-1}}{\rm tr}(\mathbf{Z}_{\color[rgb]{0,0,0}s})$
+$\begin{array}[]{ll}{\color[rgb]{0,0,0}\rm s.t.}&\left[\mathbf{W}\right]_{r,r}=1,r=0,{\color[rgb]{0,0,0}1,}\cdots,{\color[rgb]{0,0,0}N_{\text{R}}-1},\\ &\mathbf{W}\succcurlyeq 0,\\ &\displaystyle\frac{1}{\color[rgb]{0,0,0}S}{\color[rgb]{0,0,0}\sum_{s=0}^{S-1}}\begin{bmatrix}\mathbf{Z}_{\color[rgb]{0,0,0}s}&\tilde{\bm{\Xi}}(\bm{p}_{\color[rgb]{0,0,0}s})\\ \tilde{\bm{\Xi}}^{\top}(\bm{p}_{\color[rgb]{0,0,0}s})&\mathbf{B}(\bm{p}_{\color[rgb]{0,0,0}s},\mathbf{W})\end{bmatrix}\succcurlyeq 0,\end{array}$ (54)
 where, according to the relationship in (49) and the calculation of PEB in (43), minimizing the objective function results in minimizing the average square of PEB, which also minimizes the initial objective function, the average PEB.
-Proposition 1. $\mathbf{B}(\bm{p}_{u},\mathbf{W})$ is linear to $\mathbf{W}$.
+Proposition 1. $\mathbf{B}(\bm{p}_{\color[rgb]{0,0,0}s},\mathbf{W})$ is linear to $\mathbf{W}$.
 Proof.
@@
 $\bm{y}=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\mathbf{G}^{\top}\bm{w}+\bm{v},$ (56)
-with $\mathbf{G}\in\mathbb{C}^{N_{R}\times N}$ and $\left[\mathbf{G}\right]_{r,n}=\sqrt{\rho_{aR,n,r}}\sqrt{\rho_{Ru,n,r}}e^{-j2\pi\frac{nB}{N}(\tau_{ar}+\tau_{ru})}$.
+with $\mathbf{G}\in\mathbb{C}^{{\color[rgb]{0,0,0}N_{\text{R}}}\times N}$ and $\left[\mathbf{G}\right]_{r,n}={\color[rgb]{0,0,0}\sqrt{\rho_{{\text{aR}}}^{n,r}}}{\color[rgb]{0,0,0}\sqrt{\rho_{{\text{Ru}}}^{n,r}}}e^{-j2\pi\frac{nB}{N}({\tau_{{\color[rgb]{0,0,0}\text{a}}r}}+{\color[rgb]{0,0,0}\tau_{r\text{u}}})}$.
 Correspondingly, $\bm{\mu}$ can be rewritten as $\bm{\mu}=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\mathbf{G}^{\top}\bm{\bar{w}}^{*}$.
@@
 As a result, as can be seen from (55), every term of $\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})$ is linear to $\mathbf{W}$ because the operations of taking the real parts, summation, and scalar multiplication preserve linearity.
-Hence, $\mathbf{B}(\bm{p}_{u},\mathbf{W})$ is linear to $\mathbf{W}$.
+Hence, $\mathbf{B}(\bm{p}_{\color[rgb]{0,0,0}s},\mathbf{W})$ is linear to $\mathbf{W}$.
 ∎
@@
 The computational complexity of the proposed RIS configuration is primarily influenced by the interior point method employed to solve (V).
-The worst-case complexity of the interior point method is $\mathcal{O}(\max\left\{N_{con},N_{var}\right\}^{4}\sqrt{N_{v}})$, where $N_{con}$ and $N_{var}$ are the constraint and variable numbers, respectively [39].
-For the problem in (V), ${N_{con}}=N_{R}+2$ and ${N_{var}}=N_{R}^{2}+U(N+3)^{2}$.
+The worst-case complexity of the interior point method is $\mathcal{O}(\max\left\{N_{\color[rgb]{0,0,0}\text{con}},N_{\color[rgb]{0,0,0}\text{var}}\right\}^{4}\sqrt{N_{\color[rgb]{0,0,0}\text{var}}}{\color[rgb]{0,0,0}\log\frac{1}{\epsilon}})$, where $N_{\color[rgb]{0,0,0}\text{con}}$ and $N_{\color[rgb]{0,0,0}\text{var}}$ are the constraint and variable numbers, respectively, and $\epsilon$ is the convergence accuracy [48].
+For the problem in (V), ${N_{\color[rgb]{0,0,0}\text{con}}}={\color[rgb]{0,0,0}N_{\text{R}}}+2$ and ${N_{\color[rgb]{0,0,0}\text{var}}}={\color[rgb]{0,0,0}N_{\text{R}}}^{2}+{\color[rgb]{0,0,0}S}(N+3)^{2}$.
 Suppose that the convergence accuracy is $\epsilon$.
-The complexity of solving (V) is $\mathcal{O}((N_{R}^{2}+UN^{2})^{4.5}\log\frac{1}{\epsilon})$.
+The complexity of solving (V) is $\mathcal{O}(({\color[rgb]{0,0,0}N_{\text{R}}}^{2}+{\color[rgb]{0,0,0}S}N^{2})^{4.5}\log\frac{1}{\epsilon})$.
+In comparison, the complexity of the state of the art [26] is $\mathcal{O}((N_{\text{R}}^{2}+9{\color[rgb]{0,0,0}S})^{4.5}\log\frac{1}{\epsilon})$.
+While incurring a slightly higher complexity, our method improves the PEB by threefold over the benchmark, due to our consideration of CFO and PN, as verified numerically in Section VI-B.
 VI Numerical Evaluation
-This section elaborates on the comprehensive simulation results conducted to assess the proposed CFO, PN and position estimation and the RIS configuration algorithms.
-We adopt a 3D Cartesian coordinate system, where the AN is located at $\bm{p}_{a}=[\rm 2\>m,-2\>m,1\>m]^{\top}$, and the reference point of the RIS is located at $\bm{0}_{3\times 1}{\rm\>m}$.
+This section elaborates on the comprehensive simulation results conducted to assess the proposed CFO, PN and UE position estimation and the RIS configuration algorithms.
+We adopt a 3D Cartesian coordinate system, where the AN is located at ${\color[rgb]{0,0,0}\bm{p}_{\text{a}}}=[\rm 2\>m,-2\>m,1\>m]^{\top}$, and the reference point of the RIS is located at $\bm{0}_{3\times 1}{\rm\>m}$.
 The RIS is placed on the $yz$ plane, without loss of generality.
@@
 The light speed is $c=3\times 10^{8}\>{\rm m/s}$, and the noise power is $\sigma^{2}=-109\;{\rm dBm}$.
-$N=32$ subcarriers are used in an OFDM symbol, the central frequency is $f_{c}=2.8{\rm\>GHz}$, and the bandwidth is $B=100{\rm\>MHz}$.
-The antenna gains $G_{a}$ and $G_{t}$ are both set to $1$.
-Unless otherwise specified, $N_{R}=81$ and $\sigma_{\Delta}^{2}=10^{-3}\;{\rm rad^{2}}$ are used in the simulations.
+$N=32$ subcarriers are used in an OFDM symbol, the central frequency is ${\color[rgb]{0,0,0}f_{\text{c}}}=2.8{\rm\>GHz}$, and the bandwidth is $B=100{\rm\>MHz}$.
+The antenna gains $\color[rgb]{0,0,0}G_{\text{a}}$ and $\color[rgb]{0,0,0}G_{\text{t}}$ are both set to $1$.
+Unless otherwise specified, ${\color[rgb]{0,0,0}N_{\text{R}}}=81$ and $\sigma_{\Delta}^{2}=10^{-3}\;{\rm rad^{2}}$ are used in the simulations.
 The length of the approximate PN vector $\bm{\eta}$ is $L=16$.
@@
 (b) Convergence of the overall algorithm.
-Figure 3: The convergence of the proposed Algorithm 1, where the transmit power is $P=-10{\rm\;dBm}$.
-First, we illustrate the convergence of Algorithm 1.
+Figure 3: The convergence of the proposed Algorithm 1, where the transmit power is $P=10{\rm\;dBm}$.
+First, we demonstrate the convergence of Algorithm 1.
 Fig.
-3 illustrates the case when the transmit power is $-10{\rm\>dBm}$, which corresponds to a receive signal-to-noise ratio (SNR) of approximately $17.5{\rm\>dB}$.
+3 illustrates the case with the transmit power of $-20{\rm\>dBm}$, which corresponds to a receive SNR of approximately $37.5{\rm\>dB}$.
 Specifically, Fig.
-3(a) depicts the change of the objective function when operating the GD algorithm to estimate the UE position, and different curves correspond to the GD algorithm in different rounds of iteration.
+3(a) depicts the change of the objective function when operating the GD algorithm to estimate the UE position.
+The $k$-th curve ($k=0,1,\cdots$) from top in Fig.
+3(a) shows the convergence of the GD algorithm in the $k$-th iteration of Algorithm 1.
+The GD algorithm demonstrates good convergence and effectively reduces the objective function.
 Fig.
 3(b) focuses on the convergence of the overall algorithm, and different points on the curves represent the objective function’s values after different rounds of iteration.
-In general, the GD algorithm typically converges within 2,000 iterations in this setup, which converges increasingly earlier within an iteration as the overall algorithm runs.
-The overall algorithm generally converges within 55 iterations.
-TABLE II: Average number of iterations for Algorithm 1
+The GD algorithm generally converges within 3,000 iterations in this setup, and converges increasingly faster within an iteration.
+The overall algorithm generally converges within 150 iterations.
+Moreover, Table II shows the average number of iterations of Algorithm 1 and the RMSE of UE localization under different transmit powers, $N_{\color[rgb]{0,0,0}\text{O}}$ and $N_{\color[rgb]{0,0,0}\text{I}}$ are set to be sufficiently large, and 100 tests are conducted.
+The results indicate that as the transmit power increases, the number of iterations required decreases while the localization accuracy improves under the proposed algorithm.
+VI-B Performance of RIS Configuration
+TABLE II: Average number of iterations and UE localization RMSE for Algorithm 1, where $\epsilon_{\text{o}}=10^{-7}\times 10^{\rm Round(\log\mathcal{L})}$.
 Transmit Power (dBm) $-20$ $-10$ 00 $10$
-Avg. No. of Iterations 19.45 50.59 84.31 134.95
-Further, Table II shows the average number of iterations under different transmit powers for Algorithm 1.
-In this simulation, 100 tests are conducted, and $N_{o}$ and $N_{i}$ are set to be sufficiently large.
-The results indicate that the required iteration number increases as the transmit power increases.
-VI-B Performance of RIS Configuration
+Avg. No. of Iterations 1238.60 327.90 195.16 137.74
+RMSE (m) $0.1590$ $0.0503$ $0.0197$ $0.0090$
 (a) Random RIS phase shift
@@
 (c) Proposed optimization algorithm
-Figure 4: The PEB within AOI for various RIS configuration strategies.
+Figure 4: The logarithm of PEB within AOI for various RIS configuration strategies.
 Next, our attention is directed towards evaluating the performance of the proposed RIS configuration algorithm.
@@
 Fig.
-4 indicates that the proposed optimization algorithm improves the PEB by approximately two orders of magnitude by comparison with the random RIS phase shifts, and by approximately one order of magnitude compared to the situation where there are no hardware impairments [27].
-This fully demonstrates the significance of considering the CFO and PN in the localization of the UE, as well as the effectiveness, progressiveness, and superiority of the proposed RIS configuration algorithm.
+4 indicates that the proposed optimization algorithm improves the PEB by approximately two orders of magnitude compared to the random RIS phase shifts, and by more than threefold compared to the situation overlooking hardware impairments [26].
+This demonstrates the significance of considering the CFO and PN in the localization of the UE, as well as the effectiveness, progressiveness, and superiority of the proposed RIS configuration algorithm.
 Additionally, observed from Fig.
 4, the optimized RIS phase shifts not only improve the localization accuracy within the AOI but also significantly improve localization performance in its surrounding areas.
+(a) Different RSR volumes
+(b) Different sampling strategies
+Figure 5: The effects of different RSR volumes and random sampling strategies on the CDF of PEB within AOI.
+Furthermore, we explore the effects of the size of the random sampling region (RSR), and the strategy of random sampling on localization.
+Note that the RSR with volume $V$, coincides with the AOI by default in the RIS configuration algorithm.
+Fig.
+5(a) shows the impact of different RSR strategies on the localization performance.
+By setting the center of the RSR to the center of the AOI, it is observed that as $V$ changes, the CDF of PEB does not vary significantly.
+This is because the proposed RIS configuration algorithm can improve the localization accuracy within the encompassed region of the RSR, as already revealed in Fig.
+4.
+Fig.
+5(b) shows the effect of different random sampling strategies on the CDF of PEB within AOI.
+Uniform and Gaussian sampling are considered.
+The strategy of random sampling significantly affects the localization performance, and uniform sampling outperforms Gaussian sampling.
+Moreover, random sampling affects the performance of the RIS configuration algorithm and the localization accuracy more significantly than the RSR strategy, as also shown in Fig.
+5(a).
 VI-C Estimation Accuracy
-Figure 5: The CDF of RMSE and PEB with different $N_{R}$ and transmit powers.
+Figure 6: The CDF of RMSE and PEB with different ${\color[rgb]{0,0,0}N_{\text{R}}}$ and transmit powers.
 Then, we concentrate on the CFO, PN, and UE position estimation accuracy.
@@
 It is observed that increasing the RIS size and the transmit power results in significant enhancement of the overall localization accuracy.
-Furthermore, when the RIS size is $N_{R}=49$, the CDF of the RMSE deviates significantly from the CDF of the PEB, and does not achieve the theoretical localization accuracy.
-However, when the RIS size is increased to $N_{R}=121$, the CDF of the RMSE curve aligns more closely with the CDF of the PEB.
+Furthermore, when the RIS size is ${\color[rgb]{0,0,0}N_{\text{R}}}=49$, the CDF of the RMSE deviates significantly from the CDF of the PEB, and does not achieve the theoretical localization accuracy.
+However, when the RIS size is increased to ${\color[rgb]{0,0,0}N_{\text{R}}}=121$, the CDF of the RMSE curve aligns more closely with the CDF of the PEB.
 This phenomenon can also be observed in Fig.
-8.
-Next, we select a fixed testing position with the coordinates $\rm[1.50\>m,2.15\>m,0.45\>m]^{\top}$ to investigate the impact of the transmit power $P$, the RIS size $N_{R}$, and the PN variance $\sigma_{\Delta}^{2}$ on the CFO, PN, and UE position estimation accuracy.
-Figure 6: The RMSE of $\bm{p}_{u}$ estimation under random and optimized RIS phase shifts.
+9.
+Next, we select a fixed testing position with the coordinates $\rm[1.50\>m,2.15\>m,0.45\>m]^{\top}$ to investigate the impact of the transmit power $P$, the RIS size ${\color[rgb]{0,0,0}N_{\text{R}}}$, and the PN variance $\sigma_{\Delta}^{2}$ on the CFO, PN, and UE position estimation accuracy.
+Figure 7: The RMSE of $\bm{p}_{\color[rgb]{0,0,0}\text{u}}$ estimation under random and optimized RIS phase shifts.
 Figure 8: ${\rm MSE}_{\phi,\bm{\theta}}$ under random and optimized RIS phase shifts.
 In Figs.
-6 and 7, we assess the effect of the optimized RIS phase shifts on the performance of estimation, compared to the random setup.
+7 and 8, we assess the effect of the optimized RIS phase shifts on the performance of estimation, compared to the random setup.
 From Fig.
-6, the localization performance with random RIS phase shifts does not experience a significant decline as the transmit power increases.
+7, the localization performance with random RIS phase shifts does not experience a significant decline as the transmit power increases.
 At the other extreme, the optimized RIS phase shifts provide an improvement of approximately two orders of magnitude in localization accuracy.
-This demonstrates the crucial role of optimizing the RIS phase shifts in localization and underscores the effectiveness of the proposed RIS configuration optimization algorithm.
+This demonstrates the crucial role of optimizing the RIS phase shifts in localization and underscores the effectiveness of the proposed RIS configuration algorithm.
 On the other hand, Fig.
@@
 Thus, the estimation performance of the CFO and PN parameters is ignored.
-Figure 8: The RMSE of $\bm{p}_{u}$ estimation at $N_{R}=49$, $N_{R}=81$, and $N_{R}=121$.
-Figure 9: ${\rm MSE}_{\phi,\bm{\theta}}$ at $N_{R}=49$, $N_{R}=81$, and $N_{R}=121$.
+Figure 9: The RMSE of $\bm{p}_{\color[rgb]{0,0,0}\text{u}}$ estimation at ${\color[rgb]{0,0,0}N_{\text{R}}}=49$, ${\color[rgb]{0,0,0}N_{\text{R}}}=81$, and ${\color[rgb]{0,0,0}N_{\text{R}}}=121$.
+Figure 10: ${\rm MSE}_{\phi,\bm{\theta}}$ at ${\color[rgb]{0,0,0}N_{\text{R}}}=49$, ${\color[rgb]{0,0,0}N_{\text{R}}}=81$, and ${\color[rgb]{0,0,0}N_{\text{R}}}=121$.
 Figs.
-8 and 9 depict the curves of the RMSE for UE localization and MSE for the estimation of $\underline{\bm{\gamma}}$, respectively, under different transmit powers, where $N_{R}$ is 49, 81, or 121.
+9 and 10 depict the curves of the RMSE for UE localization and MSE for the estimation of $\underline{\bm{\gamma}}$, respectively, under different transmit powers, where ${\color[rgb]{0,0,0}N_{\text{R}}}$ is set to 49, 81, or 121.
 It is observed that increasing the transmit power and/or the RIS element number results in gains in all parameters to be estimated.
-However, as the transmit SNR increases, the RMSE and MSE do not decrease as sharply as HCRLB because of the approximate PN vector, as well as the Taylor series approximation considered when deriving closed-form expressions for the CFO and PN.
-Figure 10: ${\rm RMSE}_{u}$ at $\sigma_{\Delta}^{2}=10^{-3}{\rm\>rad^{2}}$ and $\sigma_{\Delta}^{2}=10^{-4}{\rm\>rad^{2}}$.
+However, as the transmit SNR increases, the RMSE and MSE do not decrease as sharply as HCRLB, because of the approximate PN vector, as well as the Taylor series approximation considered when deriving closed-form expressions for the CFO and PN.
+Figure 11: ${\rm RMSE}_{\text{u}}$ at $\sigma_{\Delta}^{2}=10^{-3}{\rm\>rad^{2}}$ and $\sigma_{\Delta}^{2}=10^{-4}{\rm\>rad^{2}}$.
 Figure 12: ${\rm MSE}_{\phi,\bm{\theta}}$ at $\sigma_{\Delta}^{2}=10^{-3}{\rm\>rad^{2}}$ and $\sigma_{\Delta}^{2}=10^{-4}{\rm\>rad^{2}}$.
 Figs.
-10 and 11 plot the RMSE for localization and MSE for estimation of $\underline{\bm{\gamma}}$, respectively, when $N_{R}$ is set to 81, where the PN variance is $\sigma_{\Delta}^{2}=10^{-3}{\rm\>rad^{2}}$ or $\sigma_{\Delta}^{2}=10^{-4}{\rm\>rad^{2}}$.
+11 and 12 plot the RMSE for localization and MSE for estimation of $\underline{\bm{\gamma}}$, respectively, when ${\color[rgb]{0,0,0}N_{\text{R}}}$ is set to 81, where the PN variance is $\sigma_{\Delta}^{2}=10^{-3}{\rm\>rad^{2}}$ or $\sigma_{\Delta}^{2}=10^{-4}{\rm\>rad^{2}}$ [49, 50, 51, 52].
+$\sigma_{\Delta}^{2}=10^{-3}\>{\rm rad^{2}}$ is a considerably high PN variance [49].
 The results demonstrate that decreasing the PN covariance leads to significant improvements in the estimation performance of the CFO, PN, and UE position.
 Another important finding is that a smaller PN variance would make the RMSE and MSE curves more closely aligned with the HCRLB curves because of the more accurately approximated parameters when updating the CFO and PN parameters.
-Figure 12: Benchmarking the performance of the proposed algorithms against the state-of-the-art approach in [27].
-Last but not least, we compare the proposed overall algorithm of the CFO, PN, and UE position estimation and the RIS configuration, with the benchmark algorithm developed in [27].
+Figure 13: Benchmarking the performance of the proposed algorithms against the state-of-the-art approach in [26], as well as the case with perfectly known CFO and PN.
+Furthermore, we compare the proposed overall algorithm of the CFO, PN, and UE position estimation and the RIS configuration, with the benchmark algorithm developed in [26].
 It is shown in Fig.
@@
 As for the localization RMSE, the benchmark shows a large gap with its corresponding PEB, and its performance shows no demonstrable correlation with the transmit power because the estimation algorithm totally overlooked the CFO and PN.
-In comparison, the proposed algorithm significantly improves the localization accuracy by about one order of magnitude under relatively low transmit powers of below $-10{\rm\>dBm}$, and by two orders of magnitude under high transmit powers of over $10{\rm\>dBm}$.
+In comparison, the proposed algorithm significantly improves the localization accuracy by approximately one order of magnitude under relatively low transmit powers of below $-10{\rm\>dBm}$, and by two orders of magnitude under high transmit powers of over $10{\rm\>dBm}$.
+Fig.
+13 also depicts the localization accuracy when the CFO and PN values are perfectly known.
+Compared to the proposed algorithm and the benchmark, effective mitigation of the CFO and PN effects can play a crucial role in improving localization accuracy.
+Furthermore, the utilization of our algorithm results in localization accuracy that is very close to the optimal value with known CFO and PN.
+Figure 14: The impact of the number of antenna elements at the UE on the PEB under different settings of $P$ and $\sigma_{\Delta}^{2}$.
+Figure 15: The impact of the number of antenna elements at the UE on the HCRLB of CFO and PN under different settings of $\sigma_{\Delta}^{2}$.
+In Figs.
+14 and 15, we investigate the impact of the number of antennas at the UE, $M$, on UE localization accuracy.
+The antennas are arranged in a uniform linear array along the $y$-axis, with spacing of $d$.
+Combining signals received by all antennas, we have $\bm{y}_{\text{com}}=\bm{\mu}_{\text{com}}+\bm{v}_{\text{com}}$, where $\bm{y}_{\text{com}}=\left[\bm{y}_{0}^{\top},\bm{y}_{1}^{\top},\cdots,\bm{y}_{M-1}^{\top}\right]^{\top}$, $\bm{\mu}_{\text{com}}=\left[\bm{\mu}_{0}^{\top},\bm{\mu}_{1}^{\top},\cdots,\bm{\mu}_{M-1}^{\top}\right]^{\top}$, and $\bm{v}_{\text{com}}=\left[\bm{v}_{0}^{\top},\bm{v}_{1}^{\top},\cdots,\bm{v}_{M-1}^{\top}\right]^{\top}$.
+$\bm{y}_{m}$, $\bm{\mu}_{m}$, and $\bm{v}_{m}$, $m=0,1,\cdots,M-1$, represent the received signal, noise-free received signal, and combined noise at the $m$-th antenna of the UE, respectively.
+$\bm{\mu}_{m}$ can be further expressed as $\bm{\mu}_{m}=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}_{m}}\mathbf{S}^{\top}\bm{h}_{m}$, where $\bm{h}_{m}$ can be obtained by substituting the coordinates of the UE’s $m$-th antenna into (7).
+$\bm{\theta}_{m},\,m=0,1,\cdots,M-1$, is the PN vector at the UE’s $m$-th antenna and is independent and identically distributed [53].
+Fig.
+14 plots the PEB with the increase in $M$.
+The impacts of the PN variance $\sigma_{\Delta}^{2}$ and the RIS size $N_{\text{R}}$ are considered.
+It is observed that as $M$ increases, the analytical localization accuracy error bound, i.e., PEB, decreases significantly.
+This indicates that increasing the number of UE antennas, like increasing the transmit power and the number of RIS elements, is effective in improving localization accuracy.
+Under our experimental settings, compared to a single-antenna UE, adding 4 receive antennas yields a greater localization performance gain than adding 40 RIS elements; adding 8 receive antennas yields a localization gain greater than increasing the transmit power by 10 dBm.
+Fig.
+15 depicts the respective HCRLBs of the combined PN vector, $\bm{\theta}_{\rm com}\triangleq\left[\bm{\theta}_{0},\bm{\theta}_{1},\cdots,\bm{\theta}_{M-1}\right]$, the PN vector at the 00-th UE, $\bm{\theta}_{0}$, and the CFO, $\phi$, under different $\sigma_{\Delta}^{2}$ settings.
+It is observed that the HCRLBs of $\bm{\theta}_{0}$ and $\phi$ decrease as $M$ increases, due to the additional information introduced by increasing the antennas at the UE.
+The HCRLB of $\phi$ decreases faster than that of $\bm{\theta}_{0}$, because the rise of $M$ increases observations of $\phi$ but does not increase observations of $\bm{\theta}_{0}$.
+Nevertheless, the improved estimation accuracy of $\phi$ and UE position indirectly improves the estimation accuracy of $\bm{\theta}_{0}$.
+By contrast, the HCRLB of the combined PN vector increases with $M$, as the increase of $M$ results in more PN vectors that require estimation.
+Even though the estimation of the PN on each individual antenna of the UE becomes more accurate, considering all PN vectors, the overall estimation accuracy deteriorates.
+VI-D IEEE 802.11n Setup
+(a) Random RIS phase shift
+(b) Proposed optimization algorithm
+Figure 16: The logarithm of PEB within AOI for various RIS configuration strategies under standard IEEE 802.11n settings.
+Figure 17: ${\rm RMSE}_{\text{u}}$ at $\sigma_{\Delta}^{2}=10^{-3}{\rm\>rad^{2}}$ and $\sigma_{\Delta}^{2}=10^{-4}{\rm\>rad^{2}}$ under standard IEEE 802.11n settings.
+Figure 18: ${\rm MSE}_{\phi,\bm{\theta}}$ at $\sigma_{\Delta}^{2}=10^{-3}{\rm\>rad^{2}}$ and $\sigma_{\Delta}^{2}=10^{-4}{\rm\>rad^{2}}$ under standard IEEE 802.11n settings.
+Following the IEEE 802.11n (Wi-Fi 4) [54] settings, the carrier frequency $f_{\color[rgb]{0,0,0}\text{c}}$ is set to $2.4\>{\rm GHz}$, $N=64$ subcarriers with $8$ null subcarriers are utilized, and the bandwidth $B$ is $20\>{\rm MHz}$.
+Fig.
+16 plots the logarithm of PEB near the AOI under both random and optimized RIS configurations.
+The proposed algorithm improves the PEB by approximately two orders of magnitude, compared to the random RIS phase shifts.
+Figs.
+17 and 18 plot the RMSE and joint MSE of CFO and PN under different $\sigma_{\Delta}^{2}$ settings, respectively.
+It is observed that decreasing the PN covariance improves the estimation performance of the CFO, PN, and UE position significantly.
+Overall, Figs.
+16, 17, and 18 provide consistent conclusions with those drawn in Figs.
+4, 11, and 12, demonstrating the ability of our algorithms to be extended to practical systems.
 VII Conclusion
-This paper investigates the optimization of RIS phase shifts and the joint estimation of CFO, PN, and UE position under imperfect hardware conditions for an OFDM system in a near-field static localization scenario.
+This paper investigates the optimization of RIS phase shifts and the joint estimation of CFO, PN, and UE position under imperfect hardware conditions for an OFDM system in a near-field localization scenario.
 An AO-based, iterative joint estimation algorithm based on the MAP criterion and GD algorithm has been proposed.
@@
 $\frac{\partial\bm{\mu}[n]}{\partial\bm{\theta}[m]}$ $=0,m\neq n,$ (59c)
-$\frac{\partial\bm{\mu}[n]}{\partial d_{Ru}}$ $=\sqrt{P}e^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}\bm{\bar{w}}^{\rm H}\frac{\partial\mathbf{G}}{\partial d_{Ru}}[\mathbf{S}]_{:,n},$ (55d)
-$\frac{\partial\bm{\mu}[n]}{\partial\phi_{az}}$ $=\sqrt{P}e^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}\bm{\bar{w}}^{\rm H}\frac{\partial\mathbf{G}}{\partial\phi_{az}}[\mathbf{S}]_{:,n},$ (55e)
-$\frac{\partial\bm{\mu}[n]}{\partial\phi_{el}}$ $=\sqrt{P}e^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}\bm{\bar{w}}^{\rm H}\frac{\partial\mathbf{G}}{\partial\phi_{el}}[\mathbf{S}]_{:,n},$ (55f)
+$\frac{\partial\bm{\mu}[n]}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}$ $=\sqrt{P}e^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}\bm{\bar{w}}^{\rm H}\frac{\partial\mathbf{G}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}[\mathbf{S}]_{:,n},$ (59d)
+$\frac{\partial\bm{\mu}[n]}{\partial{\color[rgb]{0,0,0}\phi_{\text{az}}}}$ $=\sqrt{P}e^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}\bm{\bar{w}}^{\rm H}\frac{\partial\mathbf{G}}{\partial{\color[rgb]{0,0,0}\phi_{\text{az}}}}[\mathbf{S}]_{:,n},$ (59e)
+$\frac{\partial\bm{\mu}[n]}{\partial{\color[rgb]{0,0,0}\phi_{\text{el}}}}$ $=\sqrt{P}e^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}\bm{\bar{w}}^{\rm H}\frac{\partial\mathbf{G}}{\partial{\color[rgb]{0,0,0}\phi_{\text{el}}}}[\mathbf{S}]_{:,n},$ (59f)
 and the partial derivatives of $\bm{\mu}[n]^{*}$ with regard to all parameters to be estimated are derived as follows
@@
 $\frac{\partial\bm{\mu}[n]^{*}}{\partial\bm{\theta}[m]}$ $=0,m\neq n,$ (60c)
-$\frac{\partial\bm{\mu}[n]^{*}}{\partial d_{Ru}}$ $=\sqrt{P}e^{-j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}]_{:,n}^{\rm H}\left(\frac{\partial\mathbf{G}}{\partial d_{Ru}}\right)^{\rm H}\bm{\bar{w}},$ (56d)
-$\frac{\partial\bm{\mu}[n]^{*}}{\partial\phi_{az}}$ $=\sqrt{P}e^{-j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}]_{:,n}^{\rm H}\left(\frac{\partial\mathbf{G}}{\partial\phi_{az}}\right)^{\rm H}\bm{\bar{w}},$ (56e)
-$\frac{\partial\bm{\mu}[n]^{*}}{\partial\phi_{el}}$ $=\sqrt{P}e^{-j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}]_{:,n}^{\rm H}\left(\frac{\partial\mathbf{G}}{\partial\phi_{el}}\right)^{\rm H}\bm{\bar{w}}.$ (56f)
+$\frac{\partial\bm{\mu}[n]^{*}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}$ $=\sqrt{P}e^{-j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}]_{:,n}^{\rm H}\left(\frac{\partial\mathbf{G}}{\partial{\color[rgb]{0,0,0}d_{\text{R}}}}\right)^{\rm H}\bm{\bar{w}},$ (60d)
+$\frac{\partial\bm{\mu}[n]^{*}}{\partial{\color[rgb]{0,0,0}\phi_{\text{az}}}}$ $=\sqrt{P}e^{-j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}]_{:,n}^{\rm H}\left(\frac{\partial\mathbf{G}}{\partial{\color[rgb]{0,0,0}\phi_{\text{az}}}}\right)^{\rm H}\bm{\bar{w}},$ (60e)
+$\frac{\partial\bm{\mu}[n]^{*}}{\partial{\color[rgb]{0,0,0}\phi_{\text{el}}}}$ $=\sqrt{P}e^{-j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}]_{:,n}^{\rm H}\left(\frac{\partial\mathbf{G}}{\partial{\color[rgb]{0,0,0}\phi_{\text{el}}}}\right)^{\rm H}\bm{\bar{w}}.$ (60f)